### v11

# Kinematical Condition for Spontaneous Chiral- and Gauge-Symmetry Breaking: An interpretation of Brout-Englert-Higgs mechanism

Abstract
A new interpretation of the Brout-Englert-Higgs (BEH) mechanism is proposed. The quantum vacuum before symmetry is broken is not a quiet state, but a relativistic many-body state of massless fermions and antifermions that appear and disappear rapidly. In order for their many-body state to be interpretable, they should move along a common direction of time from past to future. This a priori kinematical condition restricts possible states to exactly CP invariant state, leading to another lowest-energy state instead of the Fock vacuum. In this vacuum, whatever effective interaction acts on massless fermion, it comes back on fermion and antifermion as an inertial mass. In this physical vacuum, massless fermions and antifermions are generated as pairs, and they behave as quasi bosons. Due to Bose statistics, their direction of motion is made parallel to each other, and their transverse excitations are suppressed by an energy gap, which makes the gauge boson massive, and explains the origin of the vacuum condensate in the BEH mechanism. The dynamic part of the above effective interaction induces a Higgs-like excitation. This interpretation sheds a new light on the BEH mechanism.
keywords:
spontaneous symmetry breaking , chiral symmetry , antifermion , time direction , Bose statistics , massive gauge boson , Higgs particle
1 Introduction
Spontaneous breaking of symmetry in vacuum is divided roughly into two categories. The first is the Brout-Englert-Higgs (BEH) mechanism [1][2]. For the electroweak interaction, the Glashaw-Weinberg-Salam (GWS) model [3][4][5], which uses the BEH mechanism and predicted the Higgs particle found in 2012 [6][7], does not contradict almost all experimental results to date. The BEH mechanism is a simple model in which both of symmetry breaking and its consequence are derived by adding, to the original Lagrangian density, the following gauge coupling, Higgs potential and Yukawa interaction
$L_{h}(x)=|(i\partial_{\mu}+gB_{\mu})h|^{2}-\mu^{2}|h|^{2}-\lambda|h|^{4}-\frac{m_{f}}{v_{h}}h\bar{\varphi}\varphi.$ (1)
Switching the sign of $\mu^{2}$ leads to the broken-symmetry vacuum with the vacuum condensate $v_{h}$. This $v_{h}$ gives a mass $m_{B}=gv_{h}$ to the vector-Abelian-gauge field $B_{\mu}$, and also gives a finite mass to fermion, but its value $m_{f}$ is substantially a free parameter. The above $h$ becomes $v_{h}+h_{1}+ih_{2}$ composed of the Higgs field $h_{1}$ and the Goldstone mode $h_{2}$. The Higgs potential explains many phenomena using a small number of parameters. This is because it plays a double role: the role of causing symmetry breaking in vacuum and that of predicting the Higgs particle’s mass. Furthermore, it stabilizes the broken-symmetry vacuum, and it represents the interaction between the Higgs particle. In this sense, the Higgs potential is a simple and economical model, and it also has the flexibility to adopt to complex situations of the eletroweak interaction. For its simplicity, however, we face some difficult problems. The switch of the sign of $\mu^{2}$ is an ad hoc assumption without explanation. The origin of the vacuum condensate $v_{h}$ is not clear. When the Higgs potential is used in the perturbation calculation, we must care a lot about the intricate cancellation of the quadratic divergence, and the fine-tuning problem arises. In order to understand the physics behind this phenomenological potential, we need to look at the problem from a somewhat different angle.
The second is the dynamical symmetry breaking. Starting from the Fock vacuum, it is assumed that interactions between particles through gauge field induces the chiral-symmetry-broken vacuum. A symmetry-broken state actually replaces a symmetry-maintained state, which is a physical process we can really observe in experiments. Phase transitions in the non-relativistic condensed matter belong to this category. To understand the mechanism behind the BEH mechanism, many attempts have been made to consider the vacuum by analogy of them. This is a probable analogy, and everything depends on the nature of interaction, but we do not yet have a clear understanding of how and when such a symmetry breaking occurs in vacuum.
1.1 Hypothetical world
In this paper, we propose a different approach to this problem. For a symmetry that is already broken in our world, a hypothetical world in which such a symmetry is maintained is considered. Although all material particles have their own masses, we consider a hypothetical world consisting of massless material particles in which chiral symmetry is maintained. If we can explain why such a world is not exactly realized as far as we follow the premise of relativistic quantum field theory, we will have a deeper understanding of reality. In this sense, the process from hypothetical to real world is a thought experiment.
Let us examine the vacuum of hypothetical world in which fermions and gauge bosons are still massless, using a simple Lagrangian density,
$L_{0}(x)=-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+\bar{\varphi}(i\partial_{\mu}+gB_{\mu})\gamma^{\mu}\varphi,$ (2)
where $\varphi$ is a massless Dirac fermi field [8],
$\varphi(x)=\frac{1}{\sqrt{V}}\sum_{p,s}\left[a^{s}(\mbox{\boldmath$p$})u^{s}(p)e^{-ipx}+b^{s\dagger}(\mbox{\boldmath$p$})v^{s}(p)e^{ipx}\right].$ (3)
If we consider massless particles to be real entities, they become a subject of the theory of relativity, even if the velocity of the massless particle is not so large as that of light.
(1) In the hypothetical world before the mass generation of fermions, due to pair production, the number of massless fermion and antifermion is not fixed [9].
(2) In the hypothetical world before the mass generation of gauge boson, the gauge interaction is still a long-range one.
The hypothetical world is a many-body state of massless fermions and antifermions in which, being mediated by the long-range force, they interact to each other up to infinite order of perturbation
For massive fermion, the lowest-energy state is the Fock vacuum, but for the massless one, we can relax this rule, and it is not correct that the quantum vacuum is quiet with hardly any excitation. The operators $a^{s}(\mbox{\boldmath$p$})$ and $b^{s\dagger}(\mbox{\boldmath$p$})$ as raising and lowering operator is closely linked to the Fock vacuum. For the above hypothetical world, however, a new raising and lowering operator reflecting properties of the above world is needed.
1.2 Raising and lowering operator
CPT invariance is a perfect invariance of nature. When time reversal is applied to one-body states, it only ensures the possibility of reverse motion. However, if it is literally applied to the many-body state, there will be a particle coming from future to present, and its relation to other particles coming from past cannot be described logically. In order to describe relativistic many-body states logically, we should describe them so that all objects move along a common direction of time from past to future. (If local inversion of temporal order is allowed, it makes a consistent interpretation impossible. This is necessary even in the many-body state governed by the underlying law satisfying time-reversal-symmetry.) When we discuss the relativistic many-body state, instead of all possible wide space of states allowed by CPT invariance, it is appropriate to restrict space of states to a narrower one satisfying only CP invariance, because time reversal plays no part in the construction of a physically-interpretable many-body state that keeps the direction of time unreversed.
Momentum, electric charge and spin, which prescribe all properties of massless fermion, are not positive definite, and therefore the effect of annihilation of massless fermion on the state is equivalent to that of creation of massless antifermion on the state. For this equivalence to be ensured naturally in the hypothetical world, a new raising and lowering operator $\widetilde{a}^{s}(\mbox{\boldmath$p$})$ that explicitly ensures CP invariance (when attention is paid to the transfer of charge and parity) is defined as
$\widetilde{a}^{s}(\mbox{\boldmath$p$})=\cos\theta_{\mbox{\boldmath$p$}}a^{s}(\mbox{\boldmath$p$})+\sin\theta_{\mbox{\boldmath$p$}}b^{s\dagger}(-\mbox{\boldmath$p$}),$ (4)
and the physical vacuum $|\widetilde{0}\rangle$ is determined as a lowest-energy state such as $\widetilde{a}^{s}(\mbox{\boldmath$p$})|\widetilde{0}\rangle=0$. If the fermion is massive, the annihilation of fermion and the creation of antifermion cause different effects on the state because mass is positive definite, and such a raising and lowering operator does not exist. Hence, it is possible only for the massless fermion and antifermion in the hypothetical world.
The necessity of antifermion in Eq.(4) changes according to the relative velocity between the incident particle and the observer, which is reflected in $\sin\theta_{\mbox{\boldmath$p$}}$. When $\mbox{\boldmath$p$}=0$, the difference between $a^{s}(\mbox{\boldmath$p$})$ and $b^{s\dagger}(-\mbox{\boldmath$p$})$ in momentum vanishes. Therefore, the importance of $a^{s}(\mbox{\boldmath$p$})$ and $b^{s\dagger}(-\mbox{\boldmath$p$})$ have the same weight for the observer, resulting in $\cos\theta_{\mbox{\boldmath$p$}}=\sin\theta_{\mbox{\boldmath$p$}}$ at $\mbox{\boldmath$p$}=0$. On the contrary, when $\mbox{\boldmath$p$}^{2}\rightarrow\infty$, such observers cannot be found, and antifermion is not needed. Hence, $\sin\theta_{\mbox{\boldmath$p$}}\rightarrow 0$ is expected.
The situation is same also for the antifermion. Hence, a new operator $\widetilde{b}^{s}(-\mbox{\boldmath$p$})$ is defined as a superposition of the annihilation of massless antifermion and the creation of massless fermion
$\widetilde{b}^{s}(-\mbox{\boldmath$p$})=\cos\theta_{\mbox{\boldmath$p$}}b^{s}(-\mbox{\boldmath$p$})-\sin\theta_{\mbox{\boldmath$p$}}a^{s\dagger}(\mbox{\boldmath$p$}).$ (5)
This $\widetilde{b}^{s}(-\mbox{\boldmath$p$})$ is orthogonal to $\widetilde{a}^{s}(-\mbox{\boldmath$p$})$ [10]. These $\widetilde{a}^{s}(\mbox{\boldmath$p$}^{\prime})$ and $\widetilde{b}^{s}(-\mbox{\boldmath$p$}^{\prime})$ define the lowest-energy one $|\widetilde{0}\rangle$ by imposing $\widetilde{a}^{s}(\mbox{\boldmath$p$})|\widetilde{0}\rangle=\widetilde{b}^{s}(-\mbox{\boldmath$p$})|\widetilde{0}\rangle=0$ on it. This lowest-energy state implies physical CP-invariant vacuum of hypothetical world.
In quantum field theory with infinite degrees of freedom, there are inequivalent representations of the same canonical commutation relation realized in different Hilbert spaces. Even if such a representation entails symmetry breaking, it should be chosen by the human side so as to ensure logical consistency. The representation for the hypothetical world should be chosen under the above priori kinematical condition.
As will be shown later, this condition plays a significant role in the spontaneous symmetry breaking. First, it leads to chiral-symmetry breaking. Whatever effective interaction acts on fermions, regardless of being repulsive or attractive, it induces the motion of others, which then comes back on the fermion and antifermion as an inertial mass. Second, massless fermion and antifermion in such a state are generated as a pair, and these pairs obey Bose statistics for certain spatial and temporal lengths. An energy gap due to Bose statistics makes this state robust against transverse perturbations, and an energy gap due to Bose statistics makes the gauge boson massive. This gives us a new viewpoint on the spontaneous symmetry breaking, and on the microscopic interpretation of the BEH mechanism.
This paper is organized as follows. In Section 2, we confirm that the lowest-energy state satisfying this condition is a broken-chiral-symmetry vacuum $|\widetilde{0}\rangle$. In Section 3, we explain the quasi-boson’s property of fermion-antifermion pairs. To describe such a vacuum, some constant parameters of physical vacuum are needed. The $v_{h}$, $\mu$ and $\lambda$ in Eq.(1) are such constants. The question is how to introduce these constants in a physically natural way. For this purpose, we define three constants of the vacuum $|\widetilde{0}\rangle$, (1) a mean spatial distance $d_{m}$ between these pairs, (2) a coherence length $l_{c}$ within which Bose statistics holds on these pairs, and (3) an upper end $\xi_{c}$ of energy-momentum in the excitation in $|\widetilde{0}\rangle$. In Section.4, with these $d_{m}$ and $l_{c}$, we propose an origin of the vacuum condensate $v_{h}$, and calculate the gauge-boson’s mass. In this interpretation, Higgs particle is not an origin of symmetry breaking but its byproduct. In Section 5, using the above $\xi_{c}$, we calculate Higgs mass $m_{H}$ as an excitation energy in $|\widetilde{0}\rangle$. Parameters in the Higgs potential will be explained in terms of these constants. In Section 6, we discuss implications of this interpretation on the BEH mechanism.
2 Physical vacuum of hypothetical world
[…]
$|\widetilde{0}\rangle=\prod_{p,s}\left[\cos\theta_{\mbox{\boldmath$p$}}+\sin\theta_{\mbox{\boldmath$p$}}e^{i\alpha(x)}b^{s\dagger}(-\mbox{\boldmath$p$})a^{s\dagger}(\mbox{\boldmath$p$})\right]|0\rangle.$ (6)
This $|\widetilde{0}\rangle$ was first introduced to elementary-particle physics by [11] in analogy with superconductivity [12]. However, it is not a specific form affected by this analogy, but a general form led inevitably from the raising and lowering operator reflecting massless relativistic many-body state. In superconductivity, the momentum of electrons in metals represents the relative motion of electrons to the center-of-mass of crystal. The naive analogy such as the center-of-mass of world cannot be carried over into the physical vacuum, because the significant momentum is only that of relative motion between observer and particle. The derivation via Eqs.(4) and (5), which uses the relative momentum between observer and particle, is a natural way to describe physical vacuum.
A phase factor $\exp[\alpha(x)]$ concerning $U(1)$ symmetry appears at each point in space-time [13]. This $|\widetilde{0}\rangle$ contains many moving massless particles, but their direction of motion are averaged out, and the relativistic normalization is not needed [14]. The superposition in Eq.(6) implies that massless pairs incessantly appear and disappear. In this sense, the physical vacuum is not a quiet vacuum.
In the BEH model, for deriving the broken-symmetry vacuum, a phenomenological treatment such as the switch of sign of $\mu^{2}$ in the Higgs potential was needed. In contrast, the argument on Eq.(6) is grounded on the basic premise: the direction of time in the relativistic many-body system. In this way, the broken-symmetry vacuum $|\widetilde{0}\rangle$ has a kinematical origin, but some dynamical consequences follow from it. The common direction of time from past to future guaranteed here enables us to describe the dynamics in $|\widetilde{0}\rangle$.
In the physical vacuum, massless fermions and antifermions will interact to each other via $B_{\mu}$. We assume that the mean-field approximation is effective also in the relativistic many-body problem, and without specifying its origin, assume a constant $U_{0}$ as a mean field,
$\bar{\varphi}(x)(i/{\!\!\!}{\partial}+U_{0})\varphi(x).$ (7)
If we would regard this symmetry breaking as a direct analogue of superconductivity, such as the technicolor model, an attractive interaction is necessary to cause symmetry breaking. However, the physical vacuum in Eq.(6) arises from kinematical reason. The comparison of energy between $|\widetilde{0}\rangle$ and the Fock vacuum is not necessary, and therefore the attractive interaction is not necessary. Whatever effective interaction acts on a massless fermion, regardless of being repulsive or attractive, it induces the motion of other fermions, which turns out to be an inertial mass of the original fermion.
[…]
$\cos^{2}\theta_{\mbox{\boldmath$p$}}=\frac{1}{2}\left(1+\frac{\epsilon_{p}}{\sqrt{\epsilon_{p}^{2}+U_{0}^{2}}}\right),\quad\sin^{2}\theta_{\mbox{\boldmath$p$}}=\frac{1}{2}\left(1-\frac{\epsilon_{p}}{\sqrt{\epsilon_{p}^{2}+U_{0}^{2}}}\right).$ (8)
This condition satisfies $\cos\theta_{\mbox{\boldmath$p$}}=\sin\theta_{\mbox{\boldmath$p$}}$ at $\mbox{\boldmath$p$}=0$, and $\sin\theta_{\mbox{\boldmath$p$}}\rightarrow 0$ at $\mbox{\boldmath$p$}\rightarrow\infty$, which satisfy the kinematical requirement we imposed in Eqs.(4) and (5). The diagonalized form of Eq.(7) includes $\sqrt{\epsilon_{p}^{2}+U_{0}^{2}}[\widetilde{a}^{s\dagger}(\mbox{\boldmath$p$})\widetilde{a}^{s}(\mbox{\boldmath$p$})+\widetilde{b}^{s\dagger}(\mbox{\boldmath$p$})\widetilde{b}^{s}(\mbox{\boldmath$p$})]$, in which $U_{0}^{2}$ is a square of the mass $m_{f}^{2}$ of the fermion and antifermion, and does not depend on the sign of $U_{0}$. After particles become massive. they do not necessarily appear as a pair, and this mass turns out to hide the hypothetical world from our eyes.
3 The behavior as Quasi bosons
In the BEH model, the vacuum condensate $v_{h}$ in $L_{h}(x)$ implies that vacuum is robust against perturbation. This robustness should be physically explained by the massless fermion and antifermion in Eq.(6). It is always as a pair that they appear and disappear in $|\widetilde{0}\rangle$ with opposite momentum and spin. For short periods after production, their fields should overlap in position space, and we represent their behavior by dimensionless composite operators $P_{\mbox{\boldmath$k$}}\equiv b(-\mbox{\boldmath$k$},\downarrow)a(\mbox{\boldmath$k$},\uparrow)$ and $P^{\dagger}_{\mbox{\boldmath$k$}}\equiv a^{\dagger}(\mbox{\boldmath$k$},\uparrow)b^{\dagger}(-\mbox{\boldmath$k$},\downarrow)$. ($\uparrow,\downarrow$ denote spins.) Owing to the anti-commutation relation, the composite operator $P_{\mbox{\boldmath$k$}}$ has a following equal-time commutation relation at $\mbox{\boldmath$k$}\neq\mbox{\boldmath$k$}^{\prime}$
$P_{\mbox{\boldmath$k$}},P^{\dagger}_{\mbox{\boldmath$k$}^{\prime}}]=0\quad\mbox{for}\quad\mbox{\boldmath$k$}\neq\mbox{\boldmath$k$}^{\prime},$ (9)
implying that different fermions-antifermions pairs with different momentum follows Bose statistics. However when $\mbox{\boldmath$k$}=\mbox{\boldmath$k$}^{\prime}$, due to Pauli principle, $P_{\mbox{\boldmath$k$}}$ shows a following commutation relation
$P_{\mbox{\boldmath$k$}},P^{\dagger}_{\mbox{\boldmath$k$}}]=1-(n_{\mbox{\boldmath$k$},\uparrow}+n_{-\mbox{\boldmath$k$},\downarrow}),\quad P_{\mbox{\boldmath$k$}}^{2}=P^{\dagger 2}_{\mbox{\boldmath$k$}}=0,$ (10)
where $n_{\mbox{\boldmath$k$},\uparrow}=a^{\dagger}(\mbox{\boldmath$k$},\uparrow)a(\mbox{\boldmath$k$},\uparrow)$ and $n_{-\mbox{\boldmath$k$},\downarrow}=b^{\dagger}(-\mbox{\boldmath$k$},\downarrow)b(-\mbox{\boldmath$k$},\downarrow)$. This is a hybrid of boson’s one $P_{\mbox{\boldmath$k$}}P^{\dagger}_{\mbox{\boldmath$k$}}=N_{\mbox{\boldmath$k$}}+1$ and fermion’s one $P_{\mbox{\boldmath$k$}}P^{\dagger}_{\mbox{\boldmath$k$}}=1-(n_{\mbox{\boldmath$k$},\uparrow}+n_{-\mbox{\boldmath$k$},\downarrow})$. These commutation relations show that the pair denoted by $P_{\mbox{\boldmath$k$}}$ behaves as a quasi bosons.
3.1 Relativistic quantum field in position space
In relativistic quantum field theory, position is complementary not only to momentum, but also to particle number. If we make a precise measurement of position, it causes a wide spread of momentum, and owing to subsequent pair production, the number of particles becomes boundless. However, position of particle is not a completely meaningless concept. In the measurement of position of moving particle with an energy $\epsilon$, the least possible error of position $\delta x$ is $\hbar c/\epsilon$. Unless we attempt to specify the position of particle more precisely than $\delta x$, it has a physical meaning. Quasi bosons appear and disappear randomly in position space, but we consider for simplicity a mean distance between quasi bosons $d_{m}$, which is larger than $\delta x$. Figure 1 schematically illustrates, using small white circles, the distribution of these quasi bosons in position space at a certain moment, and the gauge field $B_{\mu}$ propagating along the horizontal line. Since quasi bosons made of fermion and antifermion with different momentum prevail in vacuum, they follow Bose statistics as in Eq.(9). However, since these quasi bosons are not stable, their behavior as bosons is a limited one in the temporal and spatial sense. Here we define a coherence length $l_{c}$ of vacuum such as $\delta x<d_{m}<l_{c}$ so that as long as the spatial distance between these quasi bosons is less than $l_{c}$, they obey Bose statistics.
Figure 1: A schematic view of the physical vacuum at a certain moment in position space. Each white circle represents a quasi-boson, which is distributed with a mean distance $d_{m}$. (a) The transverse displacements (drdr) of quasi bosons from the horizontal line ($z$ axis) are coupled to the gauge field $B_{\mu}$ propagating along the $z$ axis, and (b) the longitudinal displacements (dzdz) are illustrated.
3.2 Statistical gap
Consider a composite scalar field $f(x)$ representing quasi bosons made of $\varphi(x)$ in Eq.(3)
$f(x)=\frac{1}{\sqrt[3]{V}}\sum_{k,s}[P_{\mbox{\boldmath$k$}}\bar{v}^{s}(-k)u^{s}(k)+P^{\dagger}_{-\mbox{\boldmath$k$}}\bar{u}^{s}(-k)v^{s}(k)]e^{ikx}.$ (11)
(The normalization volume $V$ in $\varphi(x)$ and $f(x)$ is $d_{m}^{3}$.) This $f(x)$ describes the excitations within the coherence length, which has a following kinetic energy
$H_{ef}=\int\left|\frac{\partial}{\partial x_{i}}f(x)\right|^{2}d^{3}x=-\int f^{\dagger}(x)\Delta f(x)d^{3}x.$ (12)
Let us consider a kinetic energy of quasi bosons excited by the gauge field $B_{\mu}$ from the solid horizontal line to a dotted wavy line in Figure 1(a). Bose statistics requires permutation symmetry for quasi bosons, and therefore displacements pointed by a long white arrow from the horizontal line to black circles in Figure 1(a) is indistinguishable from that by small arrows between a black circle and a neighboring white circle. Because the transverse displacements are not additive, there are many short displacements in the excitation, and the long displacements has been replaced by the short ones. (In contrast, since the longitudinal displacement is additive as shown in Figure 1(b), the long displacements are dominant, and the short displacements are exceptional.) The short distance in the denominator of Laplacian operator $\Delta$ in Eq.(12) creates a high excitation energy, and an energy gap $\epsilon_{0}$ from zero appears in the transverse excitation spectrum. This energy gap comes only from statistical property, and we call it statistical gap [15]. As a result, long displacements are substantially forbidden in the transverse displacement of quasi bosons.
This situation is expressed using a metric tensor $g_{ij}(x)$ in differential geometry. For the transverse and longitudinal distances $r$ and $z$ in Figure 1, it is expressed as $dl^{2}=g_{rr}(r)dr^{2}+dz^{2}$. When the quasi bosons in Figure 1(a) move in the transverse direction within the coherent spatial region with a size $l_{c}$, $dr$ of a small distance causes the same effect to $dl^{2}$ as that of large distance. A simple metric representing the situation is given by
$g_{rr}(r)$ $=$ $\frac{r^{2}}{d_{m}^{2}},\quad(d_{m}\leq r<l_{c}),$
$g_{rr}(r)$ $=$ $1,\qquad(0<r<d_{m},\quad l_{c}\leq r).$ (13)
The gradient in Eq.(12) is rewritten as follows (see Appendix.B)
$H_{ef}=\int g_{\mu\nu}\frac{\partial\widehat{f}^{\dagger}}{\partial x^{\mu}}\frac{\partial\widehat{f}}{\partial x^{\nu}}d^{3}x+\int W(x)\widehat{f}^{\dagger}(x)\widehat{f}(x)d^{3}x,$ (14)
where a normalized field $\widehat{f}(x)\equiv g(x)^{1/4}f(x)$ has, in addition to the kinetic energy, a square of the finite energy as follows ($\mu,\nu=x,y,z$) [16]
$W(x)=\frac{1}{4}\frac{\partial}{\partial x^{\mu}}\left(g_{\mu\nu}\frac{\partial\ln g}{\partial x^{\nu}}\right)+\frac{1}{16}g_{\mu\nu}\left(\frac{\partial\ln g}{\partial x^{\mu}}\right)\left(\frac{\partial\ln g}{\partial x^{\nu}}\right),$ (15)
where $g$ is a determinant of $g_{\mu\nu}(x)$. Using Eq.(13) and $g_{zz}=1$, $W(r)$ of each quasi boson is given by
$W(r)=\frac{3}{4}\frac{1}{d_{m}^{2}}.$ (16)
To be transversely excited, each quasi boson should jump the energy barrier $\epsilon_{0}\equiv\sqrt{W(x)}$. The sum of $\epsilon_{0}$ over quasi bosons in the transverse plane within the coherence length $\pi(l_{c}/d_{m})^{2}$ is given by
$\widehat{\epsilon}_{0}=\frac{\sqrt{3}}{2}\pi\left(\frac{l_{c}}{d_{m}}\right)^{2}\frac{1}{d_{m}}.$ (17)
Owing to this gap $\widehat{\epsilon}_{0}$, the state described by $\widehat{f}(x)$ remains in the ground state under transverse perturbations, exhibiting a kind of rigidity of the physical vacuum, thus leading to $\langle\widetilde{0}|\partial_{\mu}\widehat{f}(x)|\widetilde{0}\rangle=0$ within the coherent spatial region. Hence, the state described by $\varphi(x)$ also remains in the ground state. Owing to a flat spatial distribution, it leads to $\langle\widetilde{0}|\partial_{\mu}[\bar{\varphi}(x)\gamma^{\mu}\varphi(x)]|\widetilde{0}\rangle=0$. This is the origin of the vacuum condensate $v_{h}$ in $|(i\partial_{\mu}+gB_{\mu})(v_{h}+h_{1}+ih_{2})|^{2}$ in Eq.(1).
4 Massive gauge boson
In the BEH model, the mass of gauge boson is derived from the phenomenological coupling $|(i\partial_{\mu}+gB_{\mu})(v_{h}+h_{1}+ih_{2})|^{2}$ as $m_{B}^{2}B^{\mu}B_{\mu}=g^{2}v_{h}^{2}B^{\mu}B_{\mu}$. Instead of this, we begin with a simple Lagrangian density $L_{0}(x)$ and the physical vacuum $|\widetilde{0}\rangle$. The physical vacuum $|\widetilde{0}\rangle$ is not a simple system, and therefore the response of $|\widetilde{0}\rangle$ to $B_{\mu}$ gives rise to a non-linear effect. The minimal coupling to fermions $L_{0}^{min}(x)=\bar{\varphi}(x)(i\partial_{\mu}+gB_{\mu})\gamma^{\mu}\varphi(x)$ changes its form in $|\widetilde{0}\rangle$ by the perturbation of ${\cal H}_{I}(x)=gj^{\mu}(x)B_{\mu}(x)$ [17]. Consider a perturbation expansion of $\int d^{4}xL_{0}^{min}(x)$ in powers of $g$
$\langle\widetilde{0}|\int d^{4}x_{1}L_{0}^{min}(x_{1})exp\left(i\int{\cal H}_{I}(x_{2})d^{4}x_{2}\right)|\widetilde{0}\rangle$
$=\langle\widetilde{0}|\int d^{4}x_{1}\bar{\varphi}(x_{1})\gamma^{\mu}[i\partial_{\mu}+gB_{\mu}(x_{1})]\varphi(x_{1})|\widetilde{0}\rangle$
$+\langle\widetilde{0}|\int d^{4}x_{1}\bar{\varphi}(x_{1})\gamma^{\mu}[i\partial_{\mu}+gB_{\mu}(x_{1})]\varphi(x_{1})ig\int d^{4}x_{2}j^{\nu}(x_{2})B_{\nu}(x_{2})|\widetilde{0}\rangle$$\cdots$,
4.1 Gauge-boson’s mass due to statistical gap
(1) In the last term of Eq.(LABEL:eq:34), $B_{\nu}(x_{2})$ is correlated to $B_{\mu}(x_{1})$, yielding the following two-point-correlation function between $j^{\mu}(x_{1})$ and $j^{\nu}(x_{2})$
$\langle\widetilde{0}|\int d^{4}x_{1}{\cal H}_{I}(x_{1})\int d^{4}x_{2}{\cal H}_{I}(x_{2})|\widetilde{0}\rangle$. (19)
Since these $x_{1}$ and $x_{2}$ are separated microscopically, a distant observer regards it as a local phenomenon at $X=(x_{1}+x_{2})/2$. For such an observer, it is useful to rewrite $d^{4}x_{1}d^{4}x_{2}$ in Eq.(19) as $d^{4}Xd^{4}Y$.
$g^{2}$ $\int\langle\widetilde{0}|\int j_{\mu}(x_{1})d^{2}x_{1}\int j^{\nu}(x_{2})d^{2}x_{2}|\widetilde{0}\rangle B^{\mu}(x_{1})B_{\nu}(x_{2})d^{2}x_{1}d^{2}x_{2}$
$=$ $g^{2}$ $\int\langle\widetilde{0}|\int j_{\mu}(Y)j^{\nu}(0)d^{4}Y|\widetilde{0}\rangle\times B^{\mu}(X)B_{\nu}(X)d^{4}X$ (20)
The relative motion along $Y=x_{2}-x_{1}$ is indirectly observed in the coefficient of $B^{\mu}(X)B_{\nu}(X)$
(2) Because the gauge field is a transverse one, the excitation of fermions induced by this $B_{\mu}(x)$ is a transverse one as well. The correlation of currents in Eq.(20) is strongly influenced by the property of quasi bosons. Bose statistics requires that quasi bosons at different points move parallel to each other, and only autocorrelation of the same direction does not vanish as follows
$\langle\widetilde{0}|j_{\mu}(Y)j^{\nu}(0)|\widetilde{0}\rangle \displaystyle\Rightarrow$ $\langle\widetilde{0}|j_{\mu}(Y)j^{\mu}(0)|\widetilde{0}\rangle\quad for\quad\mu=\nu.$
$\langle\widetilde{0}|j_{\mu}(Y)j^{\nu}(0)|\widetilde{0}\rangle \displaystyle\Rightarrow$ $0\qquad\qquad\qquad\qquad for\quad\mu\neq\nu.$ (21)
Figure 2: The coherent space-time region is the inside of a small light-cone specified by $l_{c}$ (a shaded area). (a) The gauge boson $B_{\mu}$ (a wavy line) induces vacuum polarization only outside this region. (b) The chain of creation and annihilation of fermion-antifermion pairs constitutes the Higgs-like excitation $H(x)$, which is induced both inside and outside this region .
Within the coherence length $l_{c}$, the physical vacuum is robust and remains in the ground state, and $\langle\widetilde{0}|\partial_{\mu}[\bar{\varphi}(Y)\gamma^{\mu}\varphi(Y)]|\widetilde{0}\rangle=0$ holds for the transverse component of $Y$. The correlation of currents in Eq.(20) is reduced to
$\langle\widetilde{0}|j_{\mu}(Y)j^{\mu}(0)|\widetilde{0}\rangle$ $=$ $\langle\widetilde{0}|\left(j_{\mu}(0)+\left[\frac{\partial}{\partial Y_{\mu}}[\bar{\varphi}(Y)\gamma_{\mu}\varphi(Y)]\right]_{Y_{\mu}=0}Y^{\mu}+\cdots\right)j^{\mu}(0)|\widetilde{0}\rangle$ (22)
$\Rightarrow$ $\langle\widetilde{0}|j_{\mu}(0)j^{\mu}(0)|\widetilde{0}\rangle$.
Using $j^{\mu}(0)=(\varphi^{\dagger}\varphi,i\varphi^{\dagger}\gamma^{0}\mbox{\boldmath$\gamma$}\varphi)$, we obtain a common coefficient $\langle\widetilde{0}|j_{0}(0)j^{0}(0)|\widetilde{0}\rangle=\langle\widetilde{0}|j_{1}(0)j^{1}(0)|\widetilde{0}\rangle=\langle\widetilde{0}|j_{2}(0)j^{2}(0)|\widetilde{0}\rangle=\langle\widetilde{0}|j_{3}(0)j^{3}(0)|\widetilde{0}\rangle=\langle\widetilde{0}|[\varphi^{\dagger}(0)\varphi(0)]^{2}|\widetilde{0}\rangle$. Hence, the $B^{\mu}(X)B_{\nu}(X)$ term in Eq.(20) is interpreted as a mass term of gauge boson
$m_{B}^{2}\int B^{\mu}(X)B_{\mu}(X)d^{4}X.$ (23)
where $m_{B}^{2}$ is given by
$m_{B}^{2}=g^{2}\langle\widetilde{0}|\int_{Y\in Z_{c}}[\varphi^{\dagger}(0)\varphi(0)]^{2}d^{4}Y|\widetilde{0}\rangle$. (24)
The coherent space-time region $Z_{c}$ in Eq.(24) is the inside of a small light-cone as illustrated in Figure 2 for one spatial direction. Its spatial length is less than $l_{c}$, and time width is smaller than $l_{c}/c$, so that the causal relation is possible between two spatial ends separated by $l_{c}$. The 4-dimensional volume of coherent space-time region $\int_{Y\in Z_{c}}d^{4}Y=l_{c}^{2}\times\frac{1}{2}[l_{c}\times c(l_{c}/c)]\times 2=l_{c}^{4}$ in Figure 2 is Lorentz invariant. Since transverse excitations are suppressed in $Z_{c}$, the vacuum polarization induced by $B_{\mu}$ occurs only outside this region as shown in Figure 2(a).
In Eq.(24), $\varphi(x)$ has a normalization volume $V=d_{m}^{3}$ as in Eq.(3). Every fermion and antifermion in $|\widetilde{0}\rangle$ make the same contribution to $\langle\widetilde{0}|[\varphi^{\dagger}(0)\varphi(0)]^{2}|\widetilde{0}\rangle$ as
$\langle\widetilde{0}|\frac{1}{d_{m}^{6}}\sum_{p,s}\left([a^{s\dagger}(\mbox{\boldmath$p$})u^{s\dagger}(p)+b^{s}(-\mbox{\boldmath$p$})v^{s\dagger}(-p)][a^{s}(\mbox{\boldmath$p$})u^{s}(p)+b^{s\dagger}(-\mbox{\boldmath$p$})v^{s}(-p)]\right)^{2}|\widetilde{0}\rangle$ (25)
$=$ $\frac{1}{d_{m}^{6}}\langle\widetilde{0}|\sum_{p,s}[b^{s}(-\mbox{\boldmath$p$})b^{s\dagger}(-\mbox{\boldmath$p$})]^{2}+\sum_{p,s}[a^{s\dagger}(\mbox{\boldmath$p$})a^{s}(\mbox{\boldmath$p$})]^{2}$
$+$ $\sum_{p,s}b^{s}(-\mbox{\boldmath$p$})a^{s}(\mbox{\boldmath$p$})a^{s\dagger}(\mbox{\boldmath$p$})b^{s\dagger}(-\mbox{\boldmath$p$})+\sum_{p,s}a^{s\dagger}(\mbox{\boldmath$p$})b^{s\dagger}(-\mbox{\boldmath$p$})b^{s}(-\mbox{\boldmath$p$})a^{s}(\mbox{\boldmath$p$})\,|\widetilde{0}\rangle$
$=$ $2\times\frac{1}{d_{m}^{6}}\prod_{p,s}(\cos^{2}\theta_{\mbox{\boldmath$p$}}+\sin^{2}\theta_{\mbox{\boldmath$p$}})=\frac{2}{d_{m}^{6}}.$
where $u^{s\dagger}(p)u^{s}(p)=v^{s\dagger}(-p)v^{s}(-p)=1$. The gauge boson’s mass is given for $l_{c}^{4}$ of $Z_{c}$ by
$m_{B}^{2}=g^{2}\frac{2}{d_{m}^{6}}\int_{Y\in Z_{c}}d^{4}Y=g^{2}\left(\frac{\sqrt{2}l_{c}^{2}}{d_{m}^{3}}\right)^{2}.$ (26)
This $\sqrt{2}l_{c}^{2}/d_{m}^{3}$ is the origin of vacuum condensate $v_{h}$ in $m_{B}^{2}=g^{2}v_{h}^{2}$ of the BEH model, which has approximately the same value as $\widehat{\epsilon}_{0}$ in Eq.(17).
4.2 Goldstone mode
(1) The Goldstone mode exists in the last term of Eq.(LABEL:eq:34). For the first-order term of $B_{\nu}(x_{2})$ in it, integrate $\bar{\varphi}(x_{1})i\partial^{\mu}\gamma_{\mu}\varphi(x_{1})$ over $x_{1}$ by parts. Since $\varphi(x_{1})$ vanishes at $x_{1}\rightarrow\infty$, we obtain two types of terms, one including $i\partial^{\mu}\bar{\varphi}(x_{1})\gamma_{\mu}\varphi(x_{1})$, and the other including $\partial^{\mu}|\widetilde{0}\rangle$. The latter is given by
$g\langle\widetilde{0}|\int d^{4}x_{1}j_{\mu}(x_{1})\int d^{4}x_{2}j^{\nu}(x_{2})B_{\nu}(x_{2})\partial^{\mu}|\widetilde{0}\rangle$ (27)
$+$ $g\partial^{\mu}\langle\widetilde{0}|\int d^{4}x_{1}j_{\mu}(x_{1})\int d^{4}x_{2}j^{\nu}(x_{2})B_{\nu}(x_{2})|\widetilde{0}\rangle$.
Because the physical vacuum $|\widetilde{0}\rangle$ in Eq.(6) has an explicit $x$-dependence in the phase $\alpha(x)$, $\partial^{\mu}|\widetilde{0}\rangle$ contains $\partial^{\mu}\alpha(x)$. The distant observer regards Eq.(27) as representing a local phenomenon at $X$, and rewrites it using $d^{4}x_{1}d^{4}x_{2}=d^{4}Xd^{4}Y$. Hence, $m_{B}^{2}$ is extracted from Eq.(27) with reference to Eq.(22) as
$\frac{2i}{g}m_{B}^{2}\int B_{\mu}(X)\partial^{\mu}\alpha(X)d^{4}X\equiv m_{B}\int B_{\mu}(X)\partial^{\mu}G(X)d^{4}X.$ (28)
Here the Goldstone mode is defined as $G(X)\equiv 2ig^{-1}m_{B}\alpha(X)$. (While the BEH model derives the coupling of the Goldstone mode $h_{2}$ to $B_{\mu}$ from $|(i\partial_{\mu}+gB_{\mu})(v_{h}+h_{1}+ih_{2})|^{2}$, the above coupling is grounded on the response of the physical vacuum $|\widetilde{0}\rangle$ to $B_{\mu}$.)
(2) In the system without the long-range force, the global phase-rotation of fermion’s field requires no energy, and therefore the propagator of the Goldstone mode is given by
$\int\frac{d^{4}X}{(2\pi)^{4}}\langle\widetilde{0}|T[G(X)G(0)]|\widetilde{0}\rangle e^{iqX}=\frac{i}{q^{2}}.$ (29)
However, the long-range force mediated by the gauge boson prohibits a free rotation of the global phase, then preventing the Goldstone mode. This discrepancy is solved by the generation of the gauge-boson’s mass that converts the long-range force into a short-range one. The Fourier transform of Eqs.(20) and (28) are given by $m_{B}^{2}B^{\mu}(q)B_{\mu}(q)$ and $m_{B}q^{\mu}G(q)B_{\mu}(q)$, respectively. Following the usual way, regard the latter as a perturbation to the former, and the second-order perturbation is obtained as
$B^{\mu}(q)\left[im_{B}^{2}g^{\mu\nu}-m_{B}q^{\mu}\frac{i}{q^{2}}m_{B}q^{\nu}\right]B_{\nu}(q)=im_{B}^{2}\left(g^{\mu\nu}-\frac{q^{\mu}q^{\nu}}{q^{2}}\right)B^{\mu}(q)B^{\nu}(q).$ (30)
Adding Eq.(30) to the Fourier transform of $\frac{1}{4}F^{\mu\nu}F_{\mu\nu}$, and performing an inverse transformation on the resulting matrix, we obtain
$D^{\mu\nu}(q)=\frac{-i}{q^{2}-m_{B}^{2}}\left(g^{\mu\nu}-\frac{q^{\mu}q^{\nu}}{q^{2}}\right)\equiv iD(q^{2})\left(g^{\mu\nu}-\frac{q^{\mu}q^{\nu}}{q^{2}}\right),$ (31)
which is the propagator of the massive gauge boson in the Landau gauge. The response of the physical vacuum gives additional terms to $L_{0}(x)$ as $m_{B}^{2}B^{\mu}(q)B_{\mu}(q)$, $m_{B}B_{\mu}(x)\partial^{\mu}G(x)$ and $(\partial_{\mu}G(x))^{2}$.
(3) In the BEH model, the coupling of the Goldstone mode to fermions comes explicitly from the Yukawa interaction, but such a coupling appears implicitly in the first term in the right-hand side of Eq.(LABEL:eq:34). For the zeroth-order term of $B_{\mu}$, the integration over $x_{1}$ by parts yields two types of terms, one including $i\partial^{\mu}\bar{\varphi}(x_{1})\gamma_{\mu}\varphi(x_{1})$, and the other including $\partial^{\mu}|\widetilde{0}\rangle$. The latter term is given by
$i\langle\widetilde{0}|\int d^{4}x_{1}j^{\mu}(x_{1})\partial_{\mu}|\widetilde{0}\rangle+i\partial_{\mu}\langle\widetilde{0}|\int d^{4}x_{1}j^{\mu}(x_{1})|\widetilde{0}\rangle$. (32)
Since $\partial_{\mu}|\widetilde{0}\rangle=\partial_{\mu}\alpha(x)|\widetilde{0}\rangle$ contains the Goldstone mode $G(x)=2ig^{-1}m_{B}\alpha(x)$, Eq.(32) gives the coupling of the Goldstone mode $G$ to fermion
$\frac{g}{m_{B}}\langle\widetilde{0}|\int d^{4}x_{1}\bar{\varphi}(x_{1})\gamma^{\mu}\varphi(x_{1})\partial_{\mu}G(x_{1})|\widetilde{0}\rangle$. (33)
Hence, $(g/m_{B})\bar{\varphi}(x)\gamma^{\mu}\varphi(x)\partial_{\mu}G(x)$ appear in $L_{0}(x)$, which is different from $g(m_{f}/m_{B})\bar{\varphi}(x)\varphi(x)h_{2}(x)$ in the BEH model.
5 Higgs particle
In the BEH model, $\mu^{2}|h|^{2}$ in $L_{h}(x)$ of Eq.(1) plays double roles. The first is the generation of the broken-symmetry vacuum by switching the sign of $\mu^{2}$. The second is giving the Higgs particle a mass $m_{H}$. However, the former is concerned with the global property of the world, and the latter is concerned with the property of one particle. It seems strange that such different scale of things are described by the same parameter. In the present model, since symmetry breaking of vacuum has the kinematical origin, we have an option to assume a possible interaction involving fermions for the mass of Higgs particle, without caring about symmetry breaking.
The constant effective interaction $U_{0}$ in Eq.(7) is generalized to a dynamic one $U_{0}+U_{1}(x)$ as follows
$\bar{\varphi}(x)\left[i/{\!\!\!}{\partial}+U_{0}+U_{1}(x)\right]\varphi(x).$ (34)
We normalize $U_{1}(x)$ as $U_{1}(x)/U_{0}=U_{1}(x)/m_{f}$. If this dynamic interaction $U_{1}(x)$ induces an excitation $H(x)$ of fermions and antifermions, and relates it as
$H(x)\equiv\frac{m_{B}}{g}\frac{U_{1}(x)}{m_{f}},$ (35)
Eq.(34) is rewritten as
$\bar{\varphi}(x)\left[i/{\!\!\!}{\partial}+U_{0}+\widehat{g}H(x)\right]\varphi(x),$ (36)
where $\widehat{g}$ is given by
$\widehat{g}=\frac{m_{f}}{m_{B}}g.$ (37)
Owing to the relation $m_{B}=gv_{h}$ in the BEH model, this $\widehat{g}$ is $m_{f}/v_{h}$, which agrees with the coupling constant of the Higgs particle to fermions in the BEH model. Hence, $H(x)$ in Eq.(35) can be regarded as the Higgs field.
This excitation is made of chains of creations and annihilation of massless fermion-antifermion pairs propagating in space as in Figure 2(b), in which black circles linking bubbles represent a vertex in Eq.(36). Because this excitation is isotropic in space, it is represented by a scalar field. Since this induced excitation is not a transverse one, there is no energy gap in its excitation spectrum. Here we define a finite upper end $\xi_{c}$ of energy-momentum of the excited massive fermion-antifermion pairs as in the non-relativistic physics. The propagator of the Higgs excitation mode $H(x)$ is given by
$\int\frac{d^{4}x}{(2\pi)^{4}}\langle\widetilde{0}|T[H(x)H(0)]|\widetilde{0}\rangle e^{iqx}=\frac{1}{q^{2}\left[1-\chi(q^{2})\right]}.$ (38)
The self energy of the Higgs field $H(x)$ is given by
$iq^{2}\chi(q^{2})=(-i\widehat{g})^{2}(-1)\int^{\xi_{c}}_{0}\frac{d^{4}p}{(2\pi)^{4}}tr\left[\frac{i}{/{\!\!\!}{p}-m_{f}}\frac{i}{/{\!\!\!}{p}+/{\!\!\!}{q}-m_{f}}\right],$ (39)
in which $\gamma^{\mu}$ matrix is not there [18]. This integral does not represent a loop of the virtual process extending to infinite momentum. The integral over $p$ in Eq.(39) is taken from zero to this $\xi_{c}$. According to the ordinary rule, we use a new variable $l=p+xq$. The upper end in the integral over $l$ is $\sqrt{p^{2}+2xp\cdot q+x^{2}q^{2}}$, which depends on the relative direction of $p$ to $q$. Since the sign of $p\cdot q$ oscillates between positive and negative, we use a mean value $\sqrt{p^{2}+x^{2}q^{2}}$ for simplicity. Hence, using an Euclidian 4-momentum $l_{E}$ as $l^{2}=-l_{E}^{2}$, we obtain
$q^{2}\chi(q^{2})=-4\widehat{g}^{2}\int^{1}_{0}dx\int\frac{d\Omega_{4}}{(2\pi)^{4}}\int_{\sqrt{x^{2}q^{2}}}^{\sqrt{\xi_{c}^{2}+x^{2}q^{2}}}l_{E}^{3}dl_{E}\left[\frac{-l_{E}^{2}}{(l_{E}^{2}+\Delta)^{2}}+\frac{\Delta}{(l_{E}^{2}+\Delta)^{2}}\right].$ (40)
where $\Delta=m_{f}^{2}-x(1-x)q^{2}$. If we define a following integral
$I(m,n)\equiv\int l_{E}^{m}(l_{E}^{2}+\Delta)^{n}dl_{E},$ (41)
the indefinite integrals over $l_{E}$ in Eq.(40) are decomposed as follows
$I(5,-2)-\Delta\times I(3,-2)=I(1,0)+2\Delta^{2}\times I(1,-2)-3\Delta\times I(1,-1),$ (42)
where
$I(1,0)=\frac{1}{2}l_{E}^{2},\quad I(1,-2)=-\frac{1}{2(l_{E}^{2}+\Delta)},\quad I(1,-1)=\frac{1}{2}\ln|l_{E}^{2}+\Delta|.$ (43)
The definite integral over $l_{E}$ in Eq.(40) yields
$q^{2}\chi(q^{2})=\frac{\widehat{g}^{2}}{4\pi^{2}}\xi_{c}^{2}$ - $\frac{\widehat{g}^{2}}{2\pi^{2}}\int_{0}^{1}dx\Delta^{2}\left(\frac{1}{\xi_{c}^{2}+x^{2}q^{2}+\Delta}-\frac{1}{x^{2}q^{2}+\Delta}\right)$ (44)
- $\frac{\widehat{g}^{2}}{2\pi^{2}}\int_{0}^{1}dx\frac{3}{2}\Delta\ln\left|1+\frac{\xi_{c}^{2}}{x^{2}q^{2}+\Delta}\right|.$
With this $\chi(q^{2})$, the mass $m_{H}$ of the Higgs particle is defined as $\chi(q^{2})\simeq m_{H}^{2}/q^{2}$ at $q^{2}\rightarrow 0$. Since $\Delta\rightarrow m_{f}^{2}$ at $q^{2}\rightarrow 0$, the integrals over $x$ in the second and third terms of the right-hand side of Eq.(44) have following limits at $q^{2}\rightarrow 0$
$\int_{0}^{1}dx\frac{1}{\xi_{c}^{2}+x^{2}q^{2}+\Delta}\rightarrow\frac{3}{4(\xi_{c}^{2}+m_{f}^{2})},$ (45)
$\int_{0}^{1}dx\ln\left|1+\frac{\xi_{c}^{2}}{x^{2}q^{2}+\Delta}\right|\rightarrow\ln\left|\frac{\xi_{c}^{2}+m_{f}^{2}}{m_{f}^{2}}\right|.$ (46)
Plugging Eqs.(45) and (46) into Eq.(44), and using it in Eq.(38), the mass $m_{H}$ is given by
$m_{H}^{2}=\frac{\widehat{g}^{2}}{4\pi^{2}}\left[\xi_{c}^{2}+\frac{3}{2}m_{f}^{2}\left(1-\frac{m_{f}^{2}}{\xi_{c}^{2}+m_{f}^{2}}\right)-3m_{f}^{2}\ln\left(\frac{\xi_{c}^{2}+m_{f}^{2}}{m_{f}^{2}}\right)\right].$ (47)
The Higgs mass is determined by $\xi_{c}$, $m_{f}$, and $m_{B}$ in $\widehat{g}$. This $\xi_{c}$ is a finite free parameter. Once a finite value of $\xi_{c}$ is assumed, divergence that we face in the renormalization of $m_{H}$ due to virtual processes is not quadratic, but logarithmic divergence.
The Higgs excitation is described by the following effective Lagrangian density
$(\partial_{\mu}H)^{2}-m_{H}^{2}H^{2}+\frac{m_{f}}{m_{B}}g\bar{\varphi}\varphi H.$ (48)
The reason why the mass of the Higgs particle has been an unknown parameter in the electroweak theory is that it is not a quantity inferred from symmetry, but a result of the many-body phenomenon.
6 Discussion
In summary, the phenomena in the physical vacuum $|\widetilde{0}\rangle$ is described by the following total Lagrangian density $\widetilde{L}(x)$. After rewriting $\varphi$ and $\bar{\varphi}$ with $\psi$ and $\bar{\psi}$, $\widetilde{L}(x)$ is given by
$\widetilde{L}(x)=$ - $\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+m_{B}^{2}B^{\mu}B_{\mu}+\bar{\psi}(i\partial_{\mu}+gB_{\mu})\gamma^{\mu}\psi-m_{f}\bar{\psi}\psi$ (49)
$+$ $(\partial_{\mu}G)^{2}+m_{B}B_{\mu}\partial^{\mu}G+\frac{g}{m_{B}}\bar{\psi}\gamma^{\mu}\psi\partial_{\mu}G$
$+$ $(\partial_{\mu}H)^{2}-m_{H}^{2}H^{2}+\frac{m_{f}}{m_{B}}g\bar{\psi}\psi H.$
6.1 Comparison to the BEH model
The double role of the Higgs potential, mentioned in Section.1, is dissolved. Each role is played by each physical process. The broken-symmetry vacuum is derived from the kinematical condition. The vacuum condensate $v_{h}$ is explained by the statistical gap in the transverse excitation. There is no relation between the fermion’s mass and the vacuum condensate of Higgs field. The Higgs particle’s mass is the result of the many-body phenomenon. As a result, the parameters $\mu$, $\lambda$ and $m_{f}$ in $L_{h}(x)$ of Eq.(1) have following physical interpretations
$\frac{\mu^{2}}{2\lambda}$ $=$ $\left(\frac{2l_{c}^{2}}{d_{m}^{3}}\right)^{2},\qquad(=v_{h}^{2})$
$2\mu^{2}$ $=$ $\frac{\widehat{g}^{2}}{4\pi^{2}}\left[\xi_{c}^{2}+\frac{3}{2}U_{0}^{2}\left(1-\frac{U_{0}^{2}}{\xi_{c}^{2}+U_{0}^{2}}\right)-3U_{0}^{2}\ln\left(1+\frac{\xi_{c}^{2}}{U_{0}^{2}}\right)\right],\qquad(=m_{H}^{2})$
$m_{f}$ $=$ $U_{0}.$ (50)
where $\widehat{g}=(m_{f}/m_{B})g=U_{0}(d_{m}^{3}/\sqrt{2}l_{c}^{2})$.
Compared to $L_{0}(x)+L_{h}(x)$ of the BEH model, this $\widetilde{L}(x)$ has following features.
(1) In the BEH model, $|(i\partial_{\mu}+gB_{\mu})(v_{h}+h_{1}+ih_{2})|^{2}$ predicts direct couplings of the massive gauge boson to the Higgs particle $h_{1}$ or Goldstone mode $h_{2}$ such as
$g^{2}v_{h}^{2}B^{\mu}B_{\mu}\left(1+\frac{h_{1}}{v_{h}}\right)^{2}+g^{2}B^{\mu}B_{\mu}h_{2}^{2}+2gB^{\mu}(h_{1}\partial_{\mu}h_{2}+h_{2}\partial_{\mu}h_{1})+(c.c).$ (51)
However, the Higgs-like excitation $H$ is not an elementary field in contrast to $h_{1}$, so that gauge coupling of $H$ with the same coupling constant $g$ as that of fermion does not exist. Instead of $m_{B}^{2}B^{\mu}B_{\mu}(1+h_{1}/v_{h})^{2}$, the effective coupling of $H$ to $B_{\mu}$ arises from the perturbative process through $\bar{\psi}$ and $\psi$ of $\widehat{g}\bar{\psi}\psi H$ and $g\bar{\psi}\gamma^{\mu}\psi B_{\mu}$, which gives different predictions from the BEH model.
(2) One of the important prediction by the Yukawa interaction $(m_{f}/v_{h})(v_{h}+h_{1}+ih_{2})\bar{\varphi}\varphi$ in the BEH model is that the strength of the Higgs’s coupling to fermions is proportional to the fermion’s mass, which is confirmed by experiments in the electroweak interaction. The role of the Yukawa interaction connecting fermions with $H$ or $G$ is played by $(m_{f}/m_{B})g\bar{\psi}\psi H$ and $(g/m_{B})\bar{\psi}\gamma^{\mu}\psi\partial_{\mu}G$ in the present model. The above prediction also appears in $(m_{f}/m_{B})g\bar{\psi}\psi H$. However, the coupling of the Goldstone mode $G$ to the fermion $(g/m_{B})\bar{\psi}\gamma^{\mu}\psi\partial_{\mu}G$ arises from the structure of the physical vacuum as seen in Eqs.(32) and (33). Yukawa interaction is a simple phenomenology of it.
6.2 Further implications
The present model has implications for some fundamental problems that lie in the BEH model.
(a) The present model proposes a solution to the fine-tuning problem of the quadratic divergence. Since it does not assume the Higgs potential $4\lambda v_{h}h_{1}^{3}-\lambda h_{1}^{4}$, the quadratic divergence does not occur in the perturbation calculation. The divergence we must renormalize is only logarithmic one, and there is no fine-tuning problem.
(b) According to the lattice model, which strictly preserves local gauge invariance at each stage of argument, the vacuum expectation value (VEV) of the gauge-dependent quantity vanishes, if it is calculated without gauge fixing. Hence, if $h(x)$ follows $h(x)\rightarrow h(x)\exp(i\theta(x))$ under $A_{\mu}(x)\rightarrow A_{\mu}(x)-ie^{-1}\partial_{\mu}\theta(x)$, $\langle h(x)\rangle=0$ is unavoidable. (Elitzur-De Angelis-De Falco-Guerra theorem) [19][20]. This is because the local character of gauge symmetry effectively breaks the connection in the degrees of freedom defined at different space-time points. If the Higgs particle is an elementary particle, the dependence of the finite $\langle h(x)\rangle$ on the gauge-fixing procedure does not match its fundamental nature. Instead of $v_{h}=\langle h(x)\rangle$, the vacuum is characterized by $\langle\widetilde{0}|\int[\varphi^{\dagger}(0)\varphi(0)]^{2}d^{4}Y|\widetilde{0}\rangle$ in Eq.(24). Because this condensate is gauge invariant, there is no need to worry about the vanishing of its VEV.
(c) The kinematical requirement does not end with the appearance of massless fermion and antifermion in the physical vacuum. Due to $\bar{\varphi}(i\partial_{\mu}+eB_{\mu})\gamma^{\mu}\varphi$, the massless fermion-antifermion pair annihilate to a gauge boson, and this gauge boson annihilates to other massless fermion-antifermion pair in $|\widetilde{0}\rangle$. Such a process between massless objects possesses no threshold energy, and therefore it ends with an equilibrium state between two kinds of condensate. At each point in space, the $U(1)$ gauge field $B_{\mu}$ condenses in the form of $F^{\mu\nu}(x)F_{\mu\nu}(x)$. The free vacuum $|0\rangle$ in the right-hand side of Eq.(6) should be replaced by a condensed vacuum $|0_{r}\rangle$ satisfying $\langle 0_{r}|F^{\mu\nu}(x)F_{\mu\nu}(x)|0_{r}\rangle\equiv\langle\widehat{t}\rangle\neq 0$. This $\langle\widehat{t}\rangle$ is a kind of material constant of vacuum, and we should redefine $|0\rangle$ by $|0_{r}\rangle$ in the right-hand side of Eq.(6). (The explicit form of $|0_{r}\rangle$ is to be studied in the future.)
(d) The vacuum is often probed using the operator-product expansion in the deep inelastic scattering, such as the pair annihilation of electron and positron to hadrons. For the cross section of this experiment, following vacuum condensates are assumed
$\sigma(e^{+}e^{-}\rightarrow hadrons)$
$=\frac{4\pi\alpha^{2}}{s}[\mbox{Im }c^{1}(q^{2})+\mbox{Im }c^{\bar{\varphi}\varphi}\langle 0|m\bar{\varphi}\varphi|0\rangle+\mbox{ Im }c^{F^{2}}(q^{2})\langle 0(F^{a}_{\alpha\beta})^{2}|0\rangle+\cdots].$
Such $\langle 0|m\bar{\varphi}\varphi|0\rangle$ and $\langle 0(F^{a}_{\alpha\beta})^{2}|0\rangle$ may play some role in symmetry breaking as well. The interpretation in this paper may have some implication for this guess.
(e) In the Higgs potential, $\lambda|v_{h}+h_{1}+ih_{2}|^{4}$ predicts the triple and quartic self-couplings of the Higgs particle $h_{1}$. More complex many-body effects than that in Figure 2 may correspond to such self couplings. So far, the agreement of the GWS model to experiments of electroweak interaction is satisfactory. When more precise measurements are performed, however, there is a possibility of deviation, especially for light quarks and leptons of the first and second generations. The above $\widetilde{L}(x)$ predicts some different results from those by the BEH model. The next subject is to extend it to the electroweak interaction.
Appendix A The vacuum $|\widetilde{0}\rangle$ satisfying $\widetilde{a}^{s}(\mbox{\boldmath$p$})|\widetilde{0}\rangle=\widetilde{b}^{s}(-\mbox{\boldmath$p$})|\widetilde{0}\rangle=0$
The vacuum satisfying $\widetilde{a}^{s}(\mbox{\boldmath$p$})|\widetilde{0}\rangle=\widetilde{b}^{s}(-\mbox{\boldmath$p$})|\widetilde{0}\rangle=0$ is as follows. In the expansion
$e^{-iK}Fe^{iK}=F+[-iK,F]+\frac{1}{2!}\left[-iK,[-iK,F]\right]+\cdots,$ (53)
we regard an operator $a^{s}(\mbox{\boldmath$p$})$ or $b^{s}(-\mbox{\boldmath$p$})$ as $F$, and a following operator as $K$,
$i\sum_{p,s}\theta_{\mbox{\boldmath$p$}}[b^{s\dagger}(-\mbox{\boldmath$p$})a^{s\dagger}(\mbox{\boldmath$p$})-a^{s}(\mbox{\boldmath$p$})b^{s}(-\mbox{\boldmath$p$})].$ (54)
Hence, Eqs.(4) and (5) can be rewritten in a compact form
$\widetilde{a}^{s}(\mbox{\boldmath$p$})=e^{-iK}a^{s}(\mbox{\boldmath$p$})e^{iK},\quad\widetilde{b}^{s}(-\mbox{\boldmath$p$})=e^{-iK}b^{s}(-\mbox{\boldmath$p$})e^{iK}.$ (55)
The vacuum $|\widetilde{0}\rangle$ satisfying $\widetilde{a}^{s}(\mbox{\boldmath$p$})|\widetilde{0}\rangle=\widetilde{b}^{s}(-\mbox{\boldmath$p$})|\widetilde{0}\rangle=0$ is simply expressed as $|\widetilde{0}\rangle=e^{-iK}|0\rangle$. Hence, we obtain
$|\widetilde{0}\rangle$ $=$ $\exp\left(\sum_{p,s}\theta_{\mbox{\boldmath$p$}}[b^{s\dagger}(-\mbox{\boldmath$p$})a^{s\dagger}(\mbox{\boldmath$p$})-a^{s}(\mbox{\boldmath$p$})b^{s}(-\mbox{\boldmath$p$})]\right)|0\rangle.$ (56)
$=$ $\prod_{p,s}\left[\sum_{n}\frac{1}{n!}\theta_{\mbox{\boldmath$p$}}^{n}[b^{s\dagger}(-\mbox{\boldmath$p$})a^{s\dagger}(\mbox{\boldmath$p$})-a^{s}(\mbox{\boldmath$p$})b^{s}(-\mbox{\boldmath$p$})]^{n}\right]|0\rangle.$
[…]
$\sum_{n}\frac{\theta^{n}}{n!}(b^{\dagger}a^{\dagger}-ab)^{n}|0\rangle$ $=$ $|0\rangle+\theta b^{\dagger}a^{\dagger}|0\rangle-\frac{\theta^{2}}{2!}abb^{\dagger}a^{\dagger}|0\rangle-\frac{\theta^{3}}{3!}b^{\dagger}a^{\dagger}abb^{\dagger}a^{\dagger}|0\rangle$ (57)
$+$ $\frac{\theta^{4}}{4!}abb^{\dagger}a^{\dagger}abb^{\dagger}a^{\dagger}|0\rangle+\cdots.$
In this expansion, $\cos\theta_{\mbox{\boldmath$p$}}$ appears in the sum of even-order terms of $\theta$, and $\sin\theta_{\mbox{\boldmath$p$}}$ appears in the sum of odd-order terms, and then Eq.(6) is yielded.
Appendix B Statistical gap
(1) Consider a Bose field $f(x)$ with the following kinetic energy
$H_{ef}=-\int d^{3}xf^{\dagger}(x)\Delta f(x).$ (58)
The square of the infinitesimal line-element is a quadratic function of $dx$ as $dl^{2}=g_{\mu\nu}(x)dx^{\mu}dx^{\nu}$, where $g_{\mu\nu}(x)$ is a metric tensor ($\mu,\nu=x,y,z$). The inner product of the field $f(x)$ is defined as
$\langle f(x)|f(x)\rangle=\int\sqrt{g(x)}d^{3}xf^{\dagger}(x)f(x),$ (59)
where $g(x)$ is a determinant of $g_{\mu\nu}(x)$. Consider a gradient of $f$ like $A^{\mu}=g_{\mu\nu}\partial f/\partial x^{\nu}$. Since the metric depends on the position, a derivative of a given vector $A^{\mu}$ with respect to $x^{\mu}$ is replaced by the covariant derivative as follows
$\frac{DA^{\mu}}{dx^{\mu}}=\frac{dA^{\mu}}{dx^{\mu}}+\Gamma^{\mu}_{\nu\mu}A^{\nu},$ (60)
where $\Gamma^{\mu}_{\nu\mu}$ is a connection coefficient. This $\Gamma^{\mu}_{\nu\mu}$ is expressed by the determinant of the metric tensor $g(x)=|g_{\mu\nu}(x)|$ as follows
$\Gamma^{\mu}_{\nu\mu}=\frac{1}{2g}\frac{\partial g}{\partial x^{\nu}}.$ (61)
With this expression, the covariant derivative of $A^{\mu}$ is given by
$\frac{DA^{\mu}}{dx^{\mu}}=\frac{1}{\sqrt{g}}\frac{\partial(\sqrt{g}A^{\mu})}{\partial x^{\mu}}.$ (62)
(2) One can use Eq.(62) to study the Laplacian $D\partial f/\partial^{2}x$. The matrix element of the Laplacian operator is obtained by inserting $D\partial/\partial^{2}x$ between $f^{\dagger}(x)$ and $f(x)$ in Eq.(59). For $D\partial f/\partial^{2}x$, we use Eq.(62) with $A^{\mu}=g_{\mu\nu}\partial f/\partial x^{\nu}$. After integration by parts, we get
$\langle f(x)|-\Delta|f(x)\rangle$ $=$ $-\int\sqrt{g}d^{3}xf^{\dagger}\frac{1}{\sqrt{g}}\frac{\partial}{\partial x^{\mu}}\left(\sqrt{g}g_{\mu\nu}\frac{\partial f}{\partial x^{\nu}}\right)$ (63)
$=$ $\int\sqrt{g(x)}g_{\mu\nu}d^{3}x\frac{\partial f^{\dagger}}{\partial x^{\mu}}\frac{\partial f}{\partial x^{\nu}}.$
Our interest is how the effect of permutation symmetry looks, because we observe it as an effect of Bose statistics in experiments. To rewrite Eq.(63), we introduce a new field $\widehat{f}(x)=g(x)^{1/4}f(x)$ as follows
$\int\sqrt{g(x)}d^{3}xf^{\dagger}(x)f(x)=\int d^{3}x\widehat{f}^{\dagger}(x)\widehat{f}(x).$ (64)
The gradient in Eq.(58) is rewritten using $\widehat{f}(x)$
$\frac{\partial f}{\partial x}=g^{-1/4}\left(\frac{\partial}{\partial x}-\frac{1}{4}\frac{\partial\ln g}{\partial x}\right)\widehat{f}(x).$ (65)
Hence, the matrix element of Laplacian in Eq.(63)
$\langle f(x)|-\Delta|f(x)\rangle=\int d^{3}xg_{\mu\nu}\left(\frac{\partial}{\partial x^{\mu}}-\frac{1}{4}\frac{\partial\ln g}{\partial x^{\mu}}\right)\widehat{f}^{\dagger}(x)\left(\frac{\partial}{\partial x^{\nu}}-\frac{1}{4}\frac{\partial\ln g}{\partial x^{\nu}}\right)\widehat{f}(x),$ (66)
is rewritten using the integration by parts as follows
$\langle f(x)|-\Delta|f(x)\rangle=\int d^{3}xg_{\mu\nu}\frac{\partial\widehat{f}^{\dagger}}{\partial x^{\mu}}\frac{\partial\widehat{f}}{\partial x^{\nu}}+\int W(x)\widehat{f}^{\dagger}(x)\widehat{f}(x)d^{3}x,$ (67)
where
$W(x)=\frac{1}{4}\frac{\partial}{\partial x^{\mu}}\left(g_{\mu\nu}\frac{\partial\ln g}{\partial x^{\nu}}\right)+\frac{1}{16}g_{\mu\nu}\left(\frac{\partial\ln g}{\partial x^{\mu}}\right)\left(\frac{\partial\ln g}{\partial x^{\nu}}\right).$ (68)
This $W(x)$ is a square of the finite energy gap in the excitation spectrum of $\widehat{f}(x)$ [16].