### v13

# Kinematical Condition for Spontaneous Chiral- and Gauge-Symmetry Breaking: An interpretation of Brout-Englert-Higgs mechanism

Abstract
A microscopic interpretation of the Brout-Englert-Higgs (BEH) mechanism is proposed, in which the Higgs particle is not the origin of symmetry breaking, but its byproduct. The state space of massless fermion and antifermion in the chiral-symmetric state is not a Fock space divided sharply into subspaces with definite particle-number. Rather their ground state has unsharp particle-number. From the vast non-Fock state-space, the kinematical condition reflecting unsharp particle-number selects a simple ground state, which comes from equivalence of the annihilation of massless fermion and the creation of massless antifermion with opposite momentum and spin. However, this state is not exactly realized. (1) Whatever interaction acts on them, it induces the motion of other massless ones, and comes back on them as an inertial mass. (2) They exist as pairs, and behave as quasi bosons in a short space-time. Due to Bose statistics, their direction of motion is made parallel to each other, and their transverse excitations are suppressed by an energy gap, which makes the gauge boson coupled to them massive, and explains the origin of the vacuum condensate.
1 Introduction
Spontaneous breaking of symmetry in vacuum is divided roughly into two categories. The first is the Brout-Englert-Higgs (BEH) mechanism [1][2], which is an extension of the Stueckelberg formalism [3]. The BEH mechanism is a simple model in which both of symmetry breaking and its consequence are derived by adding the following gauge coupling and Higgs potential to $\frac{1}{4}F^{\mu\nu}F_{\mu\nu}$,
$L_{h}(x)=|(i\partial_{\mu}+gB_{\mu})h|^{2}-\mu^{2}|h|^{2}-\lambda|h|^{4}.$ (1)
Switching the sign of $\mu^{2}$ leads to the broken-symmetry vacuum with the vacuum condensate $v_{h}$. This $v_{h}$ gives a mass $m_{B}=gv_{h}$ to the vector-Abelian-gauge field $B_{\mu}$. The above $h$ becomes $v_{h}+h_{1}+ih_{2}$ composed of the Higgs field $h_{1}$ and the Goldstone mode $h_{2}$. (Stueckelberg took only the Goldstone mode, and proposed a simple gauge-invariant model of massive gauge boson. )
For the electroweak interaction in which the weak interaction is effective only in the left-handed fermions, the Glashaw-Weinberg-Salam (GWS) model [4][5][6] used the BEH mechanism for the Lagrangian density $\frac{1}{4}F^{a\mu\nu}F^{a}_{\mu\nu}+\bar{\varphi}(i\partial_{\mu}+igA^{a}_{\mu}\tau_{a}+ig^{\prime}B_{\mu})\gamma^{\mu}\varphi$, and gave a finite mass $m_{f}$ to massless fermion $\varphi$ by adding the Yukawa interaction as follows
$L^{\prime}_{h}(x)=L_{h}(x)-\frac{m_{f}}{v_{h}}h\bar{\varphi}\varphi.$ (2)
This model predicted the W and Z bosons and the Higgs particle found in 1983 [7][8] and in 2012 [9][10] respectively, and does not contradict almost all experimental results to date. The Higgs potential explains many phenomena using a small number of parameters. This is because it plays a double role: the role of causing symmetry breaking in vacuum and that of predicting the Higgs particle’s mass. Furthermore, it stabilizes the broken-symmetry vacuum, and it represents the interaction between the Higgs particle. In this sense, the Higgs potential is a simple and economical model, and it also has the flexibility to adapt to complex situations of the eletroweak interaction. For its simplicity, however, we face some difficult problems. The switch of the sign of $\mu^{2}$ is an ad hoc assumption without explanation. The origin of the vacuum condensate $v_{h}$ is not clear. When the Higgs potential is used in the perturbation calculation, we must care a lot about the intricate cancellation of the quadratic divergence, and the fine-tuning problem arises.
The second is the dynamical symmetry breaking. Starting from the Fock vacuum, it is assumed that interactions between particles through gauge field induces the chiral-symmetry-broken vacuum. This change is assumed to be a physical process which we can really observe in experiments. Phase transitions in the non-relativistic condensed matter belong to this category. To understand the mechanism behind the BEH mechanism, many attempts have been made to consider the vacuum by analogy with them. This is a probable analogy, and everything depends on the nature of interaction, but we do not yet have a clear understanding of how and when such a symmetry breaking occurs in vacuum.
In order to understand the physics behind, we need to look at the problem from a different angle. In Section 2, we reconsider chiral-symmetry breaking in vacuum. Unlike massive particles, assuming the Fock state for the massless particles in the chiral-symmetric state is questionable, because mass no longer ensures the existence of individual particle. As a clue to the true ground state, we consider a kinematical condition, rather than dynamics, as a route to a different vacuum from the Fock vacuum. While the Nambu Jona-Lasinio state is selected as such a vacuum, it is not exactly realized, because the massless state is unstable due to interactions. Massless fermion becomes massive whatever interaction acts on them. The order of argument may be the reverse of the model of dynamical-symmetry-breaking beginning with Fock vacuum. The Higgs particle is not the origin of chiral- and gauge-symmetry breaking but its byproduct. In Section 3, it is explained that massless fermion-antifermion pair in this ground state has a quasi-boson’s aspect. To describe them, some constants of vacuum are needed. The $v_{h}$, $\mu$ and $\lambda$ in Eq.(1) are such constants. The question is how to introduce these constants in a physically natural way. For this purpose, we define microscopic three constants of the vacuum. In Section.4, with these constants, we propose an origin of the vacuum condensate $v_{h}$, and calculate the gauge-boson’s mass. The Higgs particle appears after chiral- and gauge-symmetry are broken. In Section 5, we calculate Higgs mass $m_{H}$ as an excitation energy of a collective motion of massive fermion and antifermion. Parameters in the Higgs potential will be explained in terms of these constants. In Section 6, we discuss some implications in the BEH mechanism.
2 Reconsideration of spontaneous symmetry breaking in vacuum
2.1 Massless particles
For a symmetry that is broken in our world, we can consider a fictitious world in which such a symmetry is maintained. Although all material particles have their own masses, we consider a fictitious world consisting of massless material particles, in which chiral symmetry is maintained. If we can explain why such a world is not exactly realized as far as we follow the premise of relativistic quantum field theory, we will have a deeper understanding of reality. In this sense, the process from fictitious to real world is a thought experiment. Because this fictitious world becomes a starting point to understand the real world, we hereafter call it protoworld.
Let us examine the vacuum in which fermions and U(1) gauge bosons are still massless, using a simple Lagrangian density,
$L_{0}(x)=-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+\bar{\varphi}(i\partial_{\mu}+gB_{\mu})\gamma^{\mu}\varphi,$ (3)
where $\varphi$ is a massless Dirac fermi field [11],
$\varphi(x)=\frac{1}{\sqrt{V}}\sum_{p,s}\left[a^{s}(\mbox{$p$})u^{s}(p)e^{-ipx}+b^{s\dagger}(\mbox{$p$})v^{s}(p)e^{ipx}\right].$ (4)
In the case of massive fermions, a finite mass divides a state into subspaces with definite numbers of particles. The operators $a^{s}(\mbox{$p$})$ and $b^{s\dagger}(\mbox{$p$})$, which strictly specifies the number of particle as a lowering or raising operator, is closely linked to Fock state. Fock state is a linear combination of basic vectors $\sum_{n}c_{n}|n_{1},n_{2},\cdots\rangle$ where $n=(n_{1},n_{2}\cdots)$ is a particle-number distribution. While the degree of freedom $\lambda$ in $n_{\lambda}$ extends to infinity, only finite number of $n_{\lambda}$ is nonzero in $|n_{1},n_{2},\cdots\rangle$ of the Fock state, such as $\sum_{n}|c_{n}|^{2}<\infty$. As a result, the number of basic vectors is countable. For the massless fermion, however, there is no reason for this rule to hold, because mass is no longer an indicator of individual particle, and only momentum remains as an continuous parameter. The distribution extending to infinity in $n$ corresponds to a real number, and the number of basic vectors is uncountable as in the cardinality of real numbers. The zero mass does not divide a state sharply into such subspaces [12]. The ground state of massless fermions is a state with unsharp particle-number.
This situation is further supported by the physical reason. If we consider massless particles to be real entities, they become a subject of the theory of relativity, even if their velocity is not so large as that of light. Due to pair production from $B_{\mu}$ in $L_{0}(x)$, the number of massless fermion and antifermion is not fixed, because there is not an energy threshold due to the mass. The state with unsharp particle-number is preferable.
The space expanded by uncountable basic vectors, however, is a vast state-space. Hence, we should proceed with the help of kinematical condition, because kinematics does not depend on the details of the interaction, and kinematics is more fundamental than dynamics in this case. For massless particle, a new raising and lowering operator reflecting this unsharp particle-number is needed.
2.2 Kinematical condition
Massless particles are nothing but a vehicles of charge for the gauge field. The form of $\bar{\varphi}(x)(i\partial_{\mu}+gB_{\mu}(x))\gamma^{\mu}\varphi(x)$ with $\varphi(x)$ in Eq.(4) implies that the annihilation $a^{s}(\mbox{$p$})$ of negative charge, which arises from the departure of massless fermion, affects the gauge field $B_{\mu}(\mbox{$p$})$ similarly to the creation $b^{s\dagger}(-\mbox{$p$})$ of positive charge, which arises from the arrival of massless antifermion with opposite momentum and spin. Furthermore, momentum, electric charge and spin, which prescribe all properties of massless fermion, are not positive definite, and therefore the effect of $a^{s}(\mbox{$p$})$ and $b^{s\dagger}(-\mbox{$p$})$ on the state is equivalent. Hence, there is no reason to assign the equivalent kinematical operations on the state to two different operators. For this kinematical condition to be ensured, a new raising and lowering operator $\widetilde{a}^{s}(\mbox{$p$})$ is defined as follows
$\widetilde{a}^{s}(\mbox{$p$})=\cos\theta_{\mbox{$p$}}a^{s}(\mbox{$p$})+\sin\theta_{\mbox{$p$}}b^{s\dagger}(-\mbox{$p$}).$ (5)
If the fermion is massive, the annihilation of fermion and the creation of antifermion cause different effects on the state because mass is positive definite, and such a raising and lowering operator does not exist. Hence, it is a characteristic of the massless fermion and antifermion.
The necessity of antifermion in Eq.(5) changes according to the relative velocity between the incident particle and the observer, which is reflected in $\sin\theta_{\mbox{$p$}}$. When $\mbox{$p$}=0$, the difference between $a^{s}(\mbox{$p$})$ and $b^{s\dagger}(-\mbox{$p$})$ in momentum vanishes. Therefore, the importance of $a^{s}(\mbox{$p$})$ and $b^{s\dagger}(-\mbox{$p$})$ have the same weight for the observer, resulting in $\cos\theta_{\mbox{$p$}}=\sin\theta_{\mbox{$p$}}$ at $\mbox{$p$}=0$. On the contrary, when $\mbox{$p$}^{2}\rightarrow\infty$, such observers cannot be found, and antifermion is not needed. Hence, $\sin\theta_{\mbox{$p$}}\rightarrow 0$ is expected. The situation is same also for the antifermion. Hence, a new operator $\widetilde{b}^{s}(-\mbox{$p$})$ is defined as a superposition of the annihilation of massless antifermion and the creation of massless fermion
$\widetilde{b}^{s}(-\mbox{$p$})=\cos\theta_{\mbox{$p$}}b^{s}(-\mbox{$p$})-\sin\theta_{\mbox{$p$}}a^{s\dagger}(\mbox{$p$}).$ (6)
This $\widetilde{b}^{s}(-\mbox{$p$})$ is orthogonal to $\widetilde{a}^{s}(-\mbox{$p$})$ [13].
[…]
$|\widetilde{0}\rangle=\prod_{p,s}\left[\cos\theta_{\mbox{$p$}}+\sin\theta_{\mbox{$p$}}e^{i\alpha(x)}b^{s\dagger}(-\mbox{$p$})a^{s\dagger}(\mbox{$p$})\right]|0\rangle.$ (7)
2.3 Reinterpretation of the Nambu Jona-Lasinio state
This $|\widetilde{0}\rangle$ was first introduced to elementary-particle physics by [14] in analogy with superconductivity [15]. However, it is not a specific form influenced by this analogy, but the first approximation of the general form for the relativistic massless many-body state. This $|\widetilde{0}\rangle$ is an example of the states in which the number of basic vector is uncountable as represented by a continuous weight-factor $\sin\theta_{\mbox{$p$}}$, which is selected by the kinematical condition among various candidates in the vast state-space.
The analogy with superconductivity has the following fault. The momentum of electrons in metals represents the relative motion of electrons to the center-of-mass of crystal. The naive analogy such as the center-of-mass of world cannot be carried over into the physical vacuum, because the significant momentum is only that of relative motion between observer and particle. The derivation via Eqs.(5) and (6), which uses the relative momentum between observer and particle, is a natural way to describe physical vacuum [16].
The kinematical condition, rather than dynamics, leads to the vacuum $|\widetilde{0}\rangle$ different from the Fock vacuum. In the real world, however, the dynamics involved in $L_{0}(x)$ prevents the exact realization of such a protoworld. Spontaneous symmetry breaking is a change of protoworld to real world in the following two ways.
(1) The massless state is an unstable existence under the influence of interaction. Chiral symmetry of the state is necessarily broken unless it is protected against perturbations by some stronger principle. Massless fermions and antifermions will interact to each other via $B_{\mu}$. Whatever interaction exists between massless fermion, regardless of being repulsive or attractive, it induces the motion of other fermions, then it comes back to the original fermion as an inertial mass. A simplest way to demonstrate this mechanism is a mean-field approximation to the interaction through $B_{\mu}$ by assuming a mean field $U_{0}$,
$\bar{\varphi}(x)(i/\penalty 50{\!\!\!}{\partial}+U_{0})\varphi(x).$ (8)
Following the procedure in [15], the condition that Eq.(8) is diagonal with respect to $\widetilde{a}^{s}(\mbox{$p$})$ and $\widetilde{b}^{s}(\mbox{$p$})$ is as follows
$\cos^{2}\theta_{\mbox{$p$}}=\frac{1}{2}\left(1+\frac{\epsilon_{p}}{\sqrt{\epsilon_{p}^{2}+U_{0}^{2}}}\right),\quad\sin^{2}\theta_{\mbox{$p$}}=\frac{1}{2}\left(1-\frac{\epsilon_{p}}{\sqrt{\epsilon_{p}^{2}+U_{0}^{2}}}\right).$ (9)
which satisfies the kinematical requirement $\cos\theta_{\mbox{$p$}}=\sin\theta_{\mbox{$p$}}$ at $\mbox{$p$}=0$, and $\sin\theta_{\mbox{$p$}}\rightarrow 0$ at $\mbox{$p$}\rightarrow\infty$, which we imposed in Eqs.(5) and (6). The diagonalized form of Eq.(8) includes $\sqrt{\epsilon_{p}^{2}+U_{0}^{2}}[\widetilde{a}^{s\dagger}(\mbox{$p$})\widetilde{a}^{s}(\mbox{$p$})+\widetilde{b}^{s\dagger}(\mbox{$p$})\widetilde{b}^{s}(\mbox{$p$})]$, in which $U_{0}^{2}$ is a square of the mass $m_{f}^{2}$ of the fermion and antifermion. Here, the sign of $U_{0}$ does not affect the result, hence it does not depend on being repulsive or attractive. If we would regard this symmetry breaking as a direct analogue of superconductivity, such as the technicolor model, an attractive interaction is necessary to cause symmetry breaking. In view of the fact that we do not yet have a clear understanding of how particles interact to each other in the physical vacuum, the attractive interaction is a somewhat wishful assumption. Here, the vacuum in Eq.(7) arises from kinematical reason, and we are free from such an assumption.
(2) These massless fermion and antifermion exist as a pair in $|\widetilde{0}\rangle$, and these pairs obey Bose statistics within certain spatial and temporal lengths. Due to Bose statistics, an energy gap appears in the transverse excitation of these pairs, which makes this state robust against transverse disturbance, and makes the gauge boson massive when it couples to these pairs, as will be explained in Section 3 and 4. This will give us an origin of the vacuum condensate $v_{h}$ in the Higgs potential.
3 The behavior as quasi bosons
In the BEH model, the vacuum condensate $v_{h}$ in $L_{h}(x)$ implies that vacuum is robust against perturbation. This robustness should be physically explained by the massless fermion and antifermion in Eq.(7). When they appear or disappear in $|\widetilde{0}\rangle$, it is always as a pair with mutually opposite momenta and spins. For short periods after production, their fields should overlap in position space, and their behavior is represented by dimensionless composite operators $P_{\mbox{$k$}}\equiv b(-\mbox{$k$},\downarrow)a(\mbox{$k$},\uparrow)$ and $P^{\dagger}_{\mbox{$k$}}\equiv a^{\dagger}(\mbox{$k$},\uparrow)b^{\dagger}(-\mbox{$k$},\downarrow)$. ($\uparrow,\downarrow$ denote spins.) Owing to the anti-commutation relation, the composite operator $P_{\mbox{$k$}}$ has the following equal-time commutation relation at $\mbox{$k$}\neq\mbox{$k$}^{\prime}$
$P_{\mbox{$k$}},P^{\dagger}_{\mbox{$k$}^{\prime}}]=0\quad\mbox{for}\quad\mbox{$k$}\neq\mbox{$k$}^{\prime},$ (10)
implying that massless fermions-antifermions pairs with different momentum follow Bose statistics. However when $\mbox{$k$}=\mbox{$k$}^{\prime}$, due to Pauli principle, $P_{\mbox{$k$}}$ shows the following commutation relation
$P_{\mbox{$k$}},P^{\dagger}_{\mbox{$k$}}]=1-(n_{\mbox{$k$},\uparrow}+n_{-\mbox{$k$},\downarrow}),\quad P_{\mbox{$k$}}^{2}=P^{\dagger 2}_{\mbox{$k$}}=0,$ (11)
where $n_{\mbox{$k$},\uparrow}=a^{\dagger}(\mbox{$k$},\uparrow)a(\mbox{$k$},\uparrow)$ and $n_{-\mbox{$k$},\downarrow}=b^{\dagger}(-\mbox{$k$},\downarrow)b(-\mbox{$k$},\downarrow)$. This is a hybrid of boson’s one $P_{\mbox{$k$}}P^{\dagger}_{\mbox{$k$}}=N_{\mbox{$k$}}+1$ and fermion’s one $P_{\mbox{$k$}}P^{\dagger}_{\mbox{$k$}}=1-(n_{\mbox{$k$},\uparrow}+n_{-\mbox{$k$},\downarrow})$. These commutation relations show that the center-of-mass motion of the pair denoted by $P_{\mbox{$k$}}$ and its response to $B_{\mu}$ are that of a quasi boson.
3.1 Relativistic quantum field in position space
In relativistic quantum field theory, position is complementary not only to momentum, but also to particle number. If we make a precise measurement of position, it causes a wide spread of momentum, and owing to subsequent pair production, the number of particles becomes indefinite. However, position of particle is not a completely meaningless concept. In the measurement of position of moving particle with an energy $\epsilon$, the least possible error of position $\delta x$ is $\hbar c/\epsilon$. Unless we attempt to specify the position of particle more precisely than $\delta x$, it has a physical meaning. Quasi bosons appear and disappear randomly in position space, but we consider for simplicity a mean distance between quasi bosons $d_{m}$, which is larger than $\delta x$. Figure 1 schematically illustrates, using small white circles, the distribution of these quasi bosons in position space at a certain moment, and the gauge field $B_{\mu}$ propagating along the horizontal line. Since quasi bosons made of fermion and antifermion with different momentum prevail in vacuum, they follow Bose statistics as in Eq.(10). However, since these quasi bosons are not stable, their behavior as bosons is a limited one in the temporal and spatial sense. Here we define a coherence length $l_{c}$ in the vacuum such as $\delta x<d_{m}<l_{c}$ so that as long as the spatial distance between these quasi bosons is less than $l_{c}$, they obey Bose statistics.
Figure 1: A schematic view of the physical vacuum at a certain moment in position space. Each white circle represents a quasi-boson, which is distributed with a mean distance $d_{m}$. (a) The transverse displacements (drdr) of quasi bosons from the horizontal line ($z$ axis) are coupled to the gauge field $B_{\mu}$ propagating along the $z$ axis, and (b) the longitudinal displacements (dzdz) are illustrated.
3.2 Statistical gap
Consider a composite scalar field $f(x)$ representing quasi bosons made of $\varphi(x)$ in Eq.(4)
$f(x)=\frac{1}{\sqrt[3]{V}}\sum_{k,s}[P_{\mbox{$k$}}\bar{v}^{s}(-k)u^{s}(k)+P^{\dagger}_{-\mbox{$k$}}\bar{u}^{s}(-k)v^{s}(k)]e^{ikx}.$ (12)
(The normalization volume $V$ in $\varphi(x)$ and $f(x)$ is $d_{m}^{3}$.) This $f(x)$ describes the excitations as bosons within the coherence length, which has the following kinetic energy
$H_{ef}=\int\left|\frac{\partial}{\partial x_{i}}f(x)\right|^{2}d^{3}x=-\int f^{\dagger}(x)\Delta f(x)d^{3}x.$ (13)
Let us consider a kinetic energy of quasi bosons when they are excited by the gauge field $B_{\mu}$ from the solid horizontal line to a dotted wavy line in Figure 1(a). Bose statistics requires permutation symmetry for quasi bosons, and therefore displacements pointed by a long white arrow from the horizontal line to black circles in Figure 1(a) is indistinguishable from that by small arrows between a black circle and a neighboring white circle. Because the transverse displacements are not additive, there are many short displacements in the excitation, and the long displacements has been replaced by the short ones. (In contrast, since the longitudinal displacement is additive as shown in Figure 1(b), the long displacements are dominant, and the short displacements are exceptional.) The short distance in the denominator of Laplacian operator $\Delta$ in Eq.(13) creates a high excitation energy, and an energy gap $\epsilon_{0}$ from zero appears in the transverse excitation spectrum. This energy gap comes only from statistical property, and we call it statistical gap [17]. As a result, long movements are substantially forbidden in the transverse motion of quasi bosons.
This situation is expressed using a metric tensor $g_{ij}(x)$ in differential geometry. For the transverse and longitudinal distances $r$ and $z$ respectively in Figure 1, the square of the line element is expressed as $dl^{2}=g_{rr}(r)dr^{2}+dz^{2}$. When the quasi bosons in Figure 1(a) move in the transverse direction within the coherent spatial region with a size $l_{c}$, $dr$ of a small distance causes the same effect to $dl^{2}$ as that of large distance. A simple metric representing this situation is given by
$g_{rr}(r)$ $=$ $\frac{r^{2}}{d_{m}^{2}},\quad(d_{m}\leq r<l_{c}),$
$g_{rr}(r)$ $=$ $1,\qquad(0<r<d_{m},\quad l_{c}\leq r).$ (14)
The gradient in Eq.(13) is rewritten as follows (see Appendix.B)
$H_{ef}=\int g_{\mu\nu}\frac{\partial\widehat{f}^{\dagger}}{\partial x^{\mu}}\frac{\partial\widehat{f}}{\partial x^{\nu}}d^{3}x+\int W(x)\widehat{f}^{\dagger}(x)\widehat{f}(x)d^{3}x,$ (15)
where a normalized field $\widehat{f}(x)\equiv g(x)^{1/4}f(x)$ has, in addition to the kinetic energy, a square of the finite energy as follows ($\mu,\nu=x,y,z$) [18]
$W(x)=\frac{1}{4}\frac{\partial}{\partial x^{\mu}}\left(g_{\mu\nu}\frac{\partial\ln g}{\partial x^{\nu}}\right)+\frac{1}{16}g_{\mu\nu}\left(\frac{\partial\ln g}{\partial x^{\mu}}\right)\left(\frac{\partial\ln g}{\partial x^{\nu}}\right),$ (16)
where $g$ is a determinant of $g_{\mu\nu}(x)$. Using Eq.(14) and $g_{zz}=1$, $W(r)$ of each quasi boson is given by
$W(r)=\frac{3}{4}\frac{1}{d_{m}^{2}}.$ (17)
For each quasi boson to be transversely excited, it should jump an energy barrier $\epsilon_{0}\equiv\sqrt{W(x)}$. The number of quasi bosons within the coherence length in the transverse plane is $\pi(l_{c}/d_{m})^{2}$. The sum of $\epsilon_{0}$ over quasi bosons in this area is given by
$\widehat{\epsilon}_{0}=\frac{\sqrt{3}}{2}\pi\left(\frac{l_{c}}{d_{m}}\right)^{2}\frac{1}{d_{m}}.$ (18)
Owing to this gap $\widehat{\epsilon}_{0}$, the state described by $\widehat{f}(x)$ remains in the ground state even if it is transversely perturbed. This exhibits a kind of rigidity of the physical vacuum, and leads to $\langle\widetilde{0}|\partial_{\mu}\widehat{f}(x)|\widetilde{0}\rangle=0$ within the coherent spatial region. Hence, the state described by $\varphi(x)$ also remains in the ground state, leading to $\langle\widetilde{0}|\partial_{\mu}[\bar{\varphi}(x)\gamma^{\mu}\varphi(x)]|\widetilde{0}\rangle=0$ for the transverse direction. This is the origin of the vacuum condensate $v_{h}$ in $|(i\partial_{\mu}+gB_{\mu})(v_{h}+h_{1}+ih_{2})|^{2}$ in Eq.(1).
4 Massive gauge boson
In the BEH model, the mass of gauge boson in $m_{B}^{2}B^{\mu}B_{\mu}=g^{2}v_{h}^{2}B^{\mu}B_{\mu}$ is derived from the phenomenological coupling $|(i\partial_{\mu}+gB_{\mu})(v_{h}+h_{1}+ih_{2})|^{2}$. In this paper, we begin with a simple Lagrangian density $L_{0}(x)$ and the vacuum $|\widetilde{0}\rangle$. The physical vacuum is not a simple system, and therefore the response of $|\widetilde{0}\rangle$ to $B_{\mu}$ gives rise to a non-linear effect. The minimal coupling to fermions $L_{0}^{min}(x)=\bar{\varphi}(x)(i\partial_{\mu}+gB_{\mu})\gamma^{\mu}\varphi(x)$ changes its form by the perturbation of ${\cal H}_{I}(x)=gj^{\mu}(x)B_{\mu}(x)$ in $|\widetilde{0}\rangle$ [19]. Consider a perturbation expansion of $\int d^{4}xL_{0}^{min}(x)$ in powers of $g$
$\langle\widetilde{0}|\int d^{4}x_{1}L_{0}^{min}(x_{1})exp\left(i\int{\cal H}_{I}(x_{2})d^{4}x_{2}\right)|\widetilde{0}\rangle$ (19)
$=\langle\widetilde{0}|\int d^{4}x_{1}\bar{\varphi}(x_{1})\gamma^{\mu}[i\partial_{\mu}+gB_{\mu}(x_{1})]\varphi(x_{1})|\widetilde{0}\rangle$
$+\langle\widetilde{0}|\int d^{4}x_{1}\bar{\varphi}(x_{1})\gamma^{\mu}[i\partial_{\mu}+gB_{\mu}(x_{1})]\varphi(x_{1})ig\int d^{4}x_{2}j^{\nu}(x_{2})B_{\nu}(x_{2})|\widetilde{0}\rangle$$\cdots$,
4.1 Gauge-boson’s mass due to statistical gap
(1) In the last term of Eq.(19), $B_{\nu}(x_{2})$ is correlated to $B_{\mu}(x_{1})$, yielding the following two-point-correlation function between $j^{\mu}(x_{1})$ and $j^{\nu}(x_{2})$
$g^{2}\int\langle\widetilde{0}|\int j_{\mu}(x_{1})d^{2}x_{1}\int j^{\nu}(x_{2})d^{2}x_{2}|\widetilde{0}\rangle B^{\mu}(x_{1})B_{\nu}(x_{2})d^{2}x_{1}d^{2}x_{2}$ (20)
Since these $x_{1}$ and $x_{2}$ are separated microscopically, a distant observer regards it as a local phenomenon at $X=(x_{1}+x_{2})/2$. For such an observer, it is useful to rewrite $d^{4}x_{1}d^{4}x_{2}$ in Eq.(20) as $d^{4}Xd^{4}Y$.
$g^{2}\int\langle\widetilde{0}|\int j_{\mu}(Y)j^{\nu}(0)d^{4}Y|\widetilde{0}\rangle\times B^{\mu}(X)B_{\nu}(X)d^{4}X$ (21)
The relative motion along $Y=x_{2}-x_{1}$ is indirectly observed in the coefficient of $B^{\mu}(X)B_{\nu}(X)$
(2) Because the gauge field is a transverse one, the excitation of fermions induced by this $B_{\mu}(x)$ is a transverse one as well. The correlation of currents in Eq.(21) implies a correlation of the center-of-mass motions of different massless fermion-antifermion pairs, which are strongly influenced by Bose statistics. In the response to $B_{\mu}$, quasi bosons move parallel to each other coherently, and autocorrelation of currents remains only in the same direction as follows
$\langle\widetilde{0}|j_{\mu}(Y)j^{\nu}(0)|\widetilde{0}\rangle \displaystyle\Rightarrow$ $\langle\widetilde{0}|j_{\mu}(Y)j^{\mu}(0)|\widetilde{0}\rangle\quad for\quad\mu=\nu.$
$\langle\widetilde{0}|j_{\mu}(Y)j^{\nu}(0)|\widetilde{0}\rangle \displaystyle\Rightarrow$ $0\qquad\qquad\qquad\qquad for\quad\mu\neq\nu.$ (22)
Figure 2: (a) In protoworld, the gauge boson in $B_{\mu}$ (a wavy line) induces vacuum polarization of massless fermion and antifermion only outside the coherent space-time region specified by $l_{c}$ (a shaded area). (b) In real world, the chain of creation and annihilation of massive fermion and antifermion constitutes the Higgs-like excitation $H(x)$, which is induced both inside and outside this region .
Within the coherence length $l_{c}$, the physical vacuum is robust and remains in the ground state, and $\langle\widetilde{0}|\partial_{\mu}[\bar{\varphi}(Y)\gamma^{\mu}\varphi(Y)]|\widetilde{0}\rangle=0$ holds for the transverse component of $Y$. The correlation of currents in Eq.(21) is reduced to
$\langle\widetilde{0}|j_{\mu}(Y)j^{\mu}(0)|\widetilde{0}\rangle$ $=$ $\langle\widetilde{0}|\left(j_{\mu}(0)+\left[\frac{\partial}{\partial Y_{\mu}}[\bar{\varphi}(Y)\gamma_{\mu}\varphi(Y)]\right]_{Y_{\mu}=0}Y^{\mu}+\cdots\right)j^{\mu}(0)|\widetilde{0}\rangle$ (23)
$\Rightarrow$ $\langle\widetilde{0}|j_{\mu}(0)j^{\mu}(0)|\widetilde{0}\rangle$.
Using $j^{\mu}(0)=(\varphi^{\dagger}\varphi,i\varphi^{\dagger}\gamma^{0}\mbox{$\gamma$}\varphi)$, we obtain $\langle\widetilde{0}|j_{0}(0)j^{0}(0)|\widetilde{0}\rangle=\langle\widetilde{0}|j_{1}(0)j^{1}(0)|\widetilde{0}\rangle=\langle\widetilde{0}|j_{2}(0)j^{2}(0)|\widetilde{0}\rangle=\langle\widetilde{0}|j_{3}(0)j^{3}(0)|\widetilde{0}\rangle=\langle\widetilde{0}|[\varphi^{\dagger}(0)\varphi(0)]^{2}|\widetilde{0}\rangle$. Hence, the term including $B^{\mu}(X)B_{\nu}(X)$ in Eq.(21) is interpreted as the mass term of gauge boson
$\frac{1}{2}m_{B}^{2}\int B^{\mu}(X)B_{\mu}(X)d^{4}X.$ (24)
where $m_{B}^{2}$ is given by
$m_{B}^{2}=2g^{2}\langle\widetilde{0}|\int_{Y\in Z_{c}}[\varphi^{\dagger}(0)\varphi(0)]^{2}d^{4}Y|\widetilde{0}\rangle$. (25)
The coherent space-time region $Z_{c}$ in Eq.(25) is the inside of a small light-cone as illustrated in Figure 2 for one spatial direction. Its spatial length is less than $l_{c}$, and time width is smaller than $l_{c}/c$, so that the causal relation is possible between two spatial ends separated by $l_{c}$. The 4-dimensional volume of coherent space-time region $\int_{Y\in Z_{c}}d^{4}Y=l_{c}^{2}\times\frac{1}{2}[l_{c}\times c(l_{c}/c)]\times 2=l_{c}^{4}$ in Figure 2 is Lorentz invariant. Since transverse excitations are suppressed in $Z_{c}$, the vacuum polarization induced by $B_{\mu}$ occurs only outside this region as shown in Figure 2(a).
The $\varphi(x)$ in Eq.(25) has a normalization volume $V=d_{m}^{3}$ as in Eq.(4). Every fermion and antifermion in $|\widetilde{0}\rangle$ make the same contribution to $\langle\widetilde{0}|[\varphi^{\dagger}(0)\varphi(0)]^{2}|\widetilde{0}\rangle$ as
$\langle\widetilde{0}|\frac{1}{d_{m}^{6}}\sum_{p,s}\left([a^{s\dagger}(\mbox{$p$})u^{s\dagger}(p)+b^{s}(-\mbox{$p$})v^{s\dagger}(-p)][a^{s}(\mbox{$p$})u^{s}(p)+b^{s\dagger}(-\mbox{$p$})v^{s}(-p)]\right)^{2}|\widetilde{0}\rangle$ (26)
$=$ $\frac{1}{d_{m}^{6}}\langle\widetilde{0}|\sum_{p,s}[b^{s}(-\mbox{$p$})b^{s\dagger}(-\mbox{$p$})]^{2}+\sum_{p,s}[a^{s\dagger}(\mbox{$p$})a^{s}(\mbox{$p$})]^{2}$
$+$ $\sum_{p,s}b^{s}(-\mbox{$p$})a^{s}(\mbox{$p$})a^{s\dagger}(\mbox{$p$})b^{s\dagger}(-\mbox{$p$})+\sum_{p,s}a^{s\dagger}(\mbox{$p$})b^{s\dagger}(-\mbox{$p$})b^{s}(-\mbox{$p$})a^{s}(\mbox{$p$})\,|\widetilde{0}\rangle$
$=$ $2\times\frac{1}{d_{m}^{6}}\prod_{p,s}(\cos^{2}\theta_{\mbox{$p$}}+\sin^{2}\theta_{\mbox{$p$}})=\frac{2}{d_{m}^{6}}.$
where $u^{s\dagger}(p)u^{s}(p)=v^{s\dagger}(-p)v^{s}(-p)=1$. The gauge boson’s mass for $l_{c}^{4}$ of $Z_{c}$ in Eq.(25) is given by
$m_{B}^{2}=g^{2}\frac{4}{d_{m}^{6}}\int_{Y\in Z_{c}}d^{4}Y=g^{2}\left(\frac{2l_{c}^{2}}{d_{m}^{3}}\right)^{2}.$ (27)
This $2l_{c}^{2}/d_{m}^{3}$ is the origin of vacuum condensate $v_{h}$ in $m_{B}^{2}=g^{2}v_{h}^{2}$ of the BEH model, which has approximately the same value as $\widehat{\epsilon}_{0}$ in Eq.(18).
4.2 Goldstone mode
(1) The Goldstone mode exists in the last term of Eq.(19). For the first-order term of $B_{\nu}(x_{2})$ coupled to $\bar{\varphi}(x_{1})\gamma_{\mu}i\partial^{\mu}\varphi(x_{1})$ in it, integrate it over $x_{1}$ by parts. Since $\varphi(x_{1})$ vanishes at $x_{1}\rightarrow\infty$, we obtain two types of terms, one including $i\partial^{\mu}\bar{\varphi}(x_{1})\gamma_{\mu}\varphi(x_{1})$, and the other including $\partial^{\mu}|\widetilde{0}\rangle$. The latter is given by
$g\langle\widetilde{0}|\int d^{4}x_{1}j_{\mu}(x_{1})\int d^{4}x_{2}j^{\nu}(x_{2})B_{\nu}(x_{2})\partial^{\mu}|\widetilde{0}\rangle$ (28)
$+$ $g\partial^{\mu}\langle\widetilde{0}|\int d^{4}x_{1}j_{\mu}(x_{1})\int d^{4}x_{2}j^{\nu}(x_{2})B_{\nu}(x_{2})|\widetilde{0}\rangle$.
Because the vacuum $|\widetilde{0}\rangle$ in Eq.(7) has an explicit $x$-dependence in the phase $\alpha(x)$, $\partial^{\mu}|\widetilde{0}\rangle$ contains $\partial^{\mu}\alpha(x)$. The distant observer regards Eq.(28) as representing a local phenomenon at $X$, and rewrites it using $d^{4}x_{1}d^{4}x_{2}=d^{4}Xd^{4}Y$. After extracting $m_{B}^{2}$ from Eq.(28) as in Eq.(21), we obtain
$\frac{i}{2g}m_{B}^{2}\int B_{\mu}(X)\partial^{\mu}\alpha(X)d^{4}X\equiv\frac{m_{B}}{\sqrt{2}}\int B_{\mu}(X)\partial^{\mu}G(X)d^{4}X.$ (29)
Here the Goldstone mode is defined as $G(X)\equiv i(\sqrt{2}g)^{-1}m_{B}\alpha(X)$.
(While the BEH model derives the coupling of the Goldstone mode $h_{2}$ to $B_{\mu}$ from $|(i\partial_{\mu}+gB_{\mu})(v_{h}+h_{1}+ih_{2})|^{2}$, the above coupling comes from the response of the vacuum $|\widetilde{0}\rangle$ to $B_{\mu}$.)
(2) In the system without the long-range force, the global phase-rotation of $\alpha(X)$ requires no energy, and therefore the propagator of the Goldstone mode is given by
$\int\frac{d^{4}X}{(2\pi)^{4}}\langle\widetilde{0}|T[G(X)G(0)]|\widetilde{0}\rangle e^{iqX}=\frac{i}{q^{2}}.$ (30)
However, the long-range force mediated by the gauge boson prohibits a free rotation of the global phase, then preventing the Goldstone mode. This discrepancy is resolved by the generation of the gauge-boson’s mass that converts the long-range force into a short-range one. The Fourier transform of Eqs.(24) and (29) are given by $(m_{B}^{2}/2)B^{\mu}(q)B_{\mu}(q)$ and $(m_{B}/\sqrt{2})q^{\mu}G(q)B_{\mu}(q)$, respectively. Following the usual way, regard the latter as a perturbation to the former, and the second-order perturbation is obtained as
$\frac{1}{2}B^{\mu}(q)\left[im_{B}^{2}g^{\mu\nu}-m_{B}q^{\mu}\frac{i}{q^{2}}m_{B}q^{\nu}\right]B_{\nu}(q)=\frac{i}{2}m_{B}^{2}\left(g^{\mu\nu}-\frac{q^{\mu}q^{\nu}}{q^{2}}\right)B^{\mu}(q)B^{\nu}(q).$ (31)
Adding Eq.(31) to the Fourier transform of $\frac{1}{4}F^{\mu\nu}F_{\mu\nu}$, and obtaining an inverse of the resulting matrix, we obtain
$D^{\mu\nu}(q)=\frac{-i}{q^{2}-m_{B}^{2}}\left(g^{\mu\nu}-\frac{q^{\mu}q^{\nu}}{q^{2}}\right)\equiv iD(q^{2})\left(g^{\mu\nu}-\frac{q^{\mu}q^{\nu}}{q^{2}}\right),$ (32)
which is the propagator of the massive gauge boson in the Landau gauge. The response of the physical vacuum gives additional terms to $L_{0}(x)$ as $(m_{B}^{2}/2)B^{\mu}(q)B_{\mu}(q)$, $(m_{B}/\sqrt{2})B_{\mu}(x)\partial^{\mu}G(x)$ and $(\partial_{\mu}G(x))^{2}$.
(3) In the BEH model, the coupling of the Goldstone mode to fermions comes explicitly from the Yukawa interaction. In the present model, however, such a coupling appears implicitly in the first term of the right-hand side in Eq.(19). For the zeroth-order term of $B_{\mu}$, the integration over $x_{1}$ by parts yields two types of terms, one including $i\partial^{\mu}\bar{\varphi}(x_{1})\gamma_{\mu}\varphi(x_{1})$, and the other including $\partial^{\mu}|\widetilde{0}\rangle$. The latter term is given by
$i\langle\widetilde{0}|\int d^{4}x_{1}j^{\mu}(x_{1})\partial_{\mu}|\widetilde{0}\rangle+i\partial_{\mu}\langle\widetilde{0}|\int d^{4}x_{1}j^{\mu}(x_{1})|\widetilde{0}\rangle$. (33)
Since $\partial_{\mu}|\widetilde{0}\rangle=\partial_{\mu}\alpha(x)|\widetilde{0}\rangle$ contains the Goldstone mode $G(x)=i(\sqrt{2}g)^{-1}m_{B}\alpha(x)$, Eq.(33) gives the coupling of the Goldstone mode $G$ to fermion
$\frac{\sqrt{2}g}{m_{B}}\langle\widetilde{0}|\int d^{4}x_{1}\bar{\varphi}(x_{1})\gamma^{\mu}\varphi(x_{1})\partial_{\mu}G(x_{1})|\widetilde{0}\rangle$. (34)
Here, a new term $(\sqrt{2}g/m_{B})\bar{\varphi}(x)\gamma^{\mu}\varphi(x)\partial_{\mu}G(x)$ is added to $L_{0}(x)$, which is different from the corresponding $g(m_{f}/m_{B})\bar{\varphi}(x)\varphi(x)h_{2}(x)$ in the BEH model.
5 Higgs particle
After the protoworld changes to the real world, chiral symmetry and gauge symmetry are broken. The massive fermion and antifermion interact to each other by transmitting the massive gauge boson. They exist not as a pair but as an independent particle, and their creation or annihilation is defined by a field [20]
$i\psi(x)=\int\frac{d^{3}p}{(2\pi)^{3}}\sum_{s}\frac{1}{\sqrt{2E_{p}}}\left[\widehat{a}^{s}(\mbox{$p$})\widehat{u}^{s}(p)e^{-ipx}+\widehat{b}^{s\dagger}(\mbox{$p$})\widehat{v}^{s}(p)e^{ipx}\right].$ (35)
In the BEH model, the Higgs particle appears in a strange manner, that is, $\mu^{2}|h|^{2}$ in $L_{h}(x)$ of Eq.(1) plays double roles. The first is the generation of the broken-symmetry vacuum by switching the sign of $\mu^{2}$. The second is giving the Higgs particle a mass $m_{H}$. However, the former is concerned with the global property of the world, and the latter is concerned with the property of one particle. It seems strange that such different scale of things are described by the same parameter. In the present model, since the vacuum $|\widetilde{0}\rangle$ has the kinematical origin, we have an option to assume a possible interaction for deriving $m_{H}$ without caring about symmetry breaking.
One of the important predictions by the BEH model incorporated into the theory of the electroweak interaction is that the strength of the Higgs’s coupling to fermions is proportional to the fermion’s mass $m_{f}$ as in the Yukawa interaction $(m_{f}/v_{h})(v_{h}+h_{1}+ih_{2})\bar{\varphi}\varphi$, which is confirmed by experiments. Hence, the next step is to go beyond the mean-field $U_{0}$ that gives a mass $m_{f}$ to fermions in Eq.(8). Normally, the kinetic energy of massive fermions is not enough to generate the on-shell massive gauge boson, and therefore the non-relativistic limit is a good approximation. (This situation continues from whatever reference frame we observe it [16].) The first approximation of the massive fermion-fermion scattering amplitude is given by
$i{\cal M}=(-ig)^{2}\bar{\widehat{u}}^{s^{\prime}}(p_{1}^{\prime})\gamma^{\mu}\widehat{u}^{s}(p_{1})\frac{-ig^{\mu\nu}}{q^{2}-m_{B}^{2}}\bar{\widehat{u}}^{s^{\prime}}(p_{2}^{\prime})\gamma_{\nu}\widehat{u}^{s}(p_{2}).$ (36)
[…]
$\bar{\psi}(x)\left[i/\penalty 50{\!\!\!}{\partial}+\widehat{g}H(x)\right]\psi(x),$ (37)
where the scalar field $H(x)$ denotes this collective motion, and $\widehat{g}$ is given by
$\widehat{g}=\frac{m_{f}}{m_{B}}g.$ (38)
The relation $m_{B}=gv_{h}$ in the BEH model implies that this $\widehat{g}$ agrees with the coupling constant $m_{f}/v_{h}$ of the Higgs particle to fermions in the Yukawa interaction in $L^{\prime}_{h}(x)$ (Eq.(2)). Hence, $H(x)$ in Eq.(37) can be regarded as the Higgs field.
The simplest collective excitation is a chain of creations and annihilation of massive fermion and antifermion propagating in space as in Figure 2(b), where black circles represent a vertex made by Eq.(37). The propagator of the Higgs field is given by
$\int\frac{d^{4}x}{(2\pi)^{4}}\langle\widetilde{0}|T[H(x)H(0)]|\widetilde{0}\rangle e^{iqx}=\frac{1}{q^{2}\left[1-\chi(q^{2})\right]}.$ (39)
Since this excitation is not a transverse one, there is no energy gap in its excitation spectrum. Here we define a finite upper end $\xi_{c}$ of energy-momentum of the excitation as in the non-relativistic physics [16]. The self energy of the Higgs particle is given by
$iq^{2}\chi(q^{2})=(-i\widehat{g})^{2}(-1)\int^{\xi_{c}}_{0}\frac{d^{4}p}{(2\pi)^{4}}tr\left[\frac{i}{/\penalty 50{\!\!\!}{p}-m_{f}}\frac{i}{/\penalty 50{\!\!\!}{p}+/\penalty 50{\!\!\!}{q}-m_{f}}\right],$ (40)
in which $\gamma^{\mu}$ matrix is not there. This integral is not a loop of the relativistic process extending to infinite momentum, but an integral over $p$ from zero to this $\xi_{c}$. According to the ordinary rule, we use a new variable $l=p+xq$. The upper end in the integral over $l$ is $\sqrt{p^{2}+2xp\cdot q+x^{2}q^{2}}$, which depends on the relative direction of $p$ to $q$. Since the sign of $p\cdot q$ oscillates between positive and negative, we use a mean value $\sqrt{p^{2}+x^{2}q^{2}}$ for simplicity. Hence, using an Euclidian 4-momentum $l_{E}$ as $l^{2}=-l_{E}^{2}$, we obtain
$q^{2}\chi(q^{2})=-4\widehat{g}^{2}\int^{1}_{0}dx\int\frac{d\Omega_{4}}{(2\pi)^{4}}\int_{\sqrt{x^{2}q^{2}}}^{\sqrt{\xi_{c}^{2}+x^{2}q^{2}}}l_{E}^{3}dl_{E}\left[\frac{-l_{E}^{2}}{(l_{E}^{2}+\Delta)^{2}}+\frac{\Delta}{(l_{E}^{2}+\Delta)^{2}}\right],$ (41)
where $\Delta=m_{f}^{2}-x(1-x)q^{2}$. If we define the following integral
$I(m,n)\equiv\int l_{E}^{m}(l_{E}^{2}+\Delta)^{n}dl_{E},$ (42)
the indefinite integrals over $l_{E}$ in Eq.(41) are decomposed as follows
$I(5,-2)-\Delta\times I(3,-2)=I(1,0)+2\Delta^{2}\times I(1,-2)-3\Delta\times I(1,-1),$ (43)
where
$I(1,0)=\frac{1}{2}l_{E}^{2},\quad I(1,-2)=-\frac{1}{2(l_{E}^{2}+\Delta)},\quad I(1,-1)=\frac{1}{2}\ln|l_{E}^{2}+\Delta|.$ (44)
The definite integral over $l_{E}$ in Eq.(41) yields
$q^{2}\chi(q^{2})=\frac{\widehat{g}^{2}}{4\pi^{2}}\xi_{c}^{2}$ - $\frac{\widehat{g}^{2}}{2\pi^{2}}\int_{0}^{1}dx\Delta^{2}\left(\frac{1}{\xi_{c}^{2}+x^{2}q^{2}+\Delta}-\frac{1}{x^{2}q^{2}+\Delta}\right)$ (45)
- $\frac{\widehat{g}^{2}}{2\pi^{2}}\int_{0}^{1}dx\frac{3}{2}\Delta\ln\left|1+\frac{\xi_{c}^{2}}{x^{2}q^{2}+\Delta}\right|.$
With this $\chi(q^{2})$, the mass $m_{H}$ of the Higgs particle is defined as $\chi(q^{2})\simeq m_{H}^{2}/q^{2}$ at $q^{2}\rightarrow 0$. Since $\Delta\rightarrow m_{f}^{2}$ at $q^{2}\rightarrow 0$, the integrals over $x$ in the second and third terms of the right-hand side of Eq.(45) have following limits at $q^{2}\rightarrow 0$
$\int_{0}^{1}dx\frac{1}{\xi_{c}^{2}+x^{2}q^{2}+\Delta}\rightarrow\frac{3}{4(\xi_{c}^{2}+m_{f}^{2})},$ (46)
$\int_{0}^{1}dx\ln\left|1+\frac{\xi_{c}^{2}}{x^{2}q^{2}+\Delta}\right|\rightarrow\ln\left|\frac{\xi_{c}^{2}+m_{f}^{2}}{m_{f}^{2}}\right|.$ (47)
Plugging Eqs.(46) and (47) into Eq.(45), and using it in Eq.(39), the mass $m_{H}$ is given by
$m_{H}^{2}=\frac{\widehat{g}^{2}}{4\pi^{2}}\left[\xi_{c}^{2}+\frac{3}{2}m_{f}^{2}\left(1-\frac{m_{f}^{2}}{\xi_{c}^{2}+m_{f}^{2}}\right)-3m_{f}^{2}\ln\left(\frac{\xi_{c}^{2}+m_{f}^{2}}{m_{f}^{2}}\right)\right].$ (48)
The Higgs mass is determined by $\xi_{c}$, $m_{f}$, and $m_{B}$ in $\widehat{g}$. This $\xi_{c}$ is a finite parameter to be determined by experiments. Once a finite value of $\xi_{c}$ is used in Eq.(48), divergence that we face in further renormalization is not a quadratic, but a logarithmic divergence.
The Higgs excitation is described by the following effective Lagrangian density
$(\partial_{\mu}H)^{2}-m_{H}^{2}H^{2}+\frac{m_{f}}{m_{B}}g\bar{\psi}\psi H.$ (49)
While the mass generation of fermion and gauge boson has an origin in the protoworld, the Higgs particle is a many-body phenomenon in the real world.
6 Discussion
6.1 Comparison to the BEH model
The phenomena in real world is described by the total Lagrangian density $\widetilde{L}(x)$, in which the responses of $|\widetilde{0}\rangle$ to $B_{\mu}$ are added to $L_{0}(x)$. After rewriting $\varphi$ and $\bar{\varphi}$ with $\psi$ and $\bar{\psi}$, $\widetilde{L}(x)$ is given by
$\widetilde{L}(x)=$ - $\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+\bar{\psi}(i\partial_{\mu}+gB_{\mu})\gamma^{\mu}\psi-m_{f}\bar{\psi}\psi$ (50)
$+$ $\frac{1}{2}m_{B}^{2}B^{\mu}B_{\mu}+\frac{m_{B}}{\sqrt{2}}B_{\mu}\partial^{\mu}G+(\partial_{\mu}G)^{2}+\frac{\sqrt{2}g}{m_{B}}\bar{\psi}\gamma^{\mu}\psi\partial_{\mu}G$
$+$ $(\partial_{\mu}H)^{2}-m_{H}^{2}H^{2}+\frac{m_{f}}{m_{B}}g\bar{\psi}\psi H.$
In this $\widetilde{L}(x)$, the double role of the Higgs potential is dissolved, and each role is played by each physical process. Symmetry breaking is not induced by the Higgs potential, but occurs in the process from protoworld to real world. The vacuum condensate $v_{h}$ arises from the statistical gap in the transverse excitation. Fermion’s mass and the vacuum condensate are not related. While the Higgs and Goldstone fields in the BEH model are defined as radial and tangential modes of the Higgs potential respectively, the Higgs particle is the many-body phenomenon in the real world, and the Goldstone field arises from the phase factor of the superposition in the vacuum $|\widetilde{0}\rangle$. Instead of parameters $\mu$ and $\lambda$ in the Higgs potential, quantities in Eq.(50) have physical interpretations as follows
$m_{B}^{2}$ $=$ $g^{2}\left(\frac{2l_{c}^{2}}{d_{m}^{3}}\right)^{2},\qquad(=g^{2}\frac{\mu^{2}}{2\lambda})$
$m_{f}$ $=$ $|U_{0}|.$
$m_{H}^{2}$ $=$ $\frac{\widehat{g}^{2}}{4\pi^{2}}\left[\xi_{c}^{2}+\frac{3}{2}U_{0}^{2}\left(1-\frac{U_{0}^{2}}{\xi_{c}^{2}+U_{0}^{2}}\right)-3U_{0}^{2}\ln\left(1+\frac{\xi_{c}^{2}}{U_{0}^{2}}\right)\right],\qquad(=2\mu^{2})$ (51)
where $\widehat{g}=(m_{f}/m_{B})g=|U_{0}|(d_{m}^{3}/\sqrt{2}l_{c}^{2})$.
Compared to $L_{0}(x)+L_{h}(x)$ of the BEH model, this $\widetilde{L}(x)$ has following features.
(1) The coupling of Higgs particle to massive gauge bosons has been playing an important role in the experimental tests of Higgs particle. In the BEH model, the gauge coupling $|(i\partial_{\mu}+gB_{\mu})(v_{h}+h_{1}+ih_{2})|^{2}$ predicts direct couplings of the massive gauge boson to the Higgs particle $h_{1}$ or Goldstone mode $h_{2}$ as follows
$g^{2}v_{h}^{2}B^{\mu}B_{\mu}\left(1+\frac{h_{1}}{v_{h}}\right)^{2}+g^{2}B^{\mu}B_{\mu}h_{2}^{2}+2gB^{\mu}(h_{1}\partial_{\mu}h_{2}+h_{2}\partial_{\mu}h_{1})+(c.c).$ (52)
In the present model, however, the Higgs field $H$ is not an elementary field but represents collective excitation, so that gauge coupling of $H$ with the same $g$ does not exist. Instead of $m_{B}^{2}B^{\mu}B_{\mu}(1+h_{1}/v_{h})^{2}$, the effective coupling of $H$ to $B_{\mu}$ arises in the perturbation process induced by $\widehat{g}\bar{\psi}\psi H$ and $g\bar{\psi}\gamma^{\mu}\psi B_{\mu}$ in Eq.(50), in which such couplings are relayed by $\bar{\psi}$ and $\psi$. The above gauge coupling in the BEH mechanism may be a useful phenomenology if $g$ is adjusted properly.
(2) In the GWS model, Higgs and Goldstone couplings to fermions are proportional to fermion’s mass. In the present model, such couplings are made by $(m_{f}/m_{B})g\bar{\psi}\psi H$ and $(\sqrt{2}g/m_{B})\bar{\psi}\gamma^{\mu}\psi\partial_{\mu}G$ in Eq.(50) respectively. For the latter, however, the Goldstone coupling to the fermion arises from the vacuum $|\widetilde{0}\rangle$ itself as seen in Eqs.(33) and (34). The Yukawa interaction $(m_{f}/v_{h})(v_{h}+h_{1}+ih_{2})\bar{\varphi}\varphi$ is a simple phenomenology for the response of physical vacuum.
6.2 Further implications
The present model has implications for some fundamental problems that lie in the BEH model.
(a) Before the GWS model was applied to the electroweak interaction, the BEH mechanism was one of theoretical possibilities. When it was applied to the electroweak interaction, the introduction of massless fermion became necessary. The Yukawa interaction is a useful phenomenological resolution to it, but the meaning of the vacuum condensate $v_{h}$ remains to be obscure. Only when the meaning of massless particle in the chiral-symmetric state is reconsidered and its ground state is discussed throughly, a concrete image of $v_{h}$ can be proposed. Since the present model does not assume the Higgs potential $4\lambda v_{h}h_{1}^{3}-\lambda h_{1}^{4}$, the quadratic divergence does not occur in the perturbation calculation, provided that $\xi_{c}$ is experimentally determined. The divergence we must renormalize is only logarithmic one. The present model proposes a solution to the fine-tuning problem of the quadratic divergence.
(b) According to the lattice model, which strictly preserves local gauge invariance at each stage of argument, the vacuum expectation value (VEV) of the gauge-dependent quantity vanishes, if it is calculated without gauge fixing. Hence, if $h(x)$ follows $h(x)\rightarrow h(x)\exp(i\theta(x))$ under $A_{\mu}(x)\rightarrow A_{\mu}(x)-ie^{-1}\partial_{\mu}\theta(x)$, $\langle h(x)\rangle=0$ is unavoidable. (Elitzur-De Angelis-De Falco-Guerra theorem) [22][23]. This is because the local character of gauge symmetry effectively breaks the connection in the degrees of freedom defined at different space-time points. If the Higgs particle is an elementary particle, the finite $\langle h(x)\rangle$ which depends on the gauge-fixing procedure does not match its fundamental nature. Instead of $v_{h}=\langle h(x)\rangle$, the vacuum is characterized by $\langle\widetilde{0}|\int[\varphi^{\dagger}(0)\varphi(0)]^{2}d^{4}Y|\widetilde{0}\rangle$ in Eq.(25). Because this condensate is gauge invariant, there is no need to worry about the vanishing of its VEV.
(c) The protoworld does not end with the massless fermion and antifermion in $|\widetilde{0}\rangle$. Due to $\bar{\varphi}(i\partial_{\mu}+eB_{\mu})\gamma^{\mu}\varphi$, the massless fermion-antifermion pair annihilate to a gauge boson, and this gauge boson annihilates to other massless fermion-antifermion pair in $|\widetilde{0}\rangle$. Such a process between massless objects possesses no threshold energy, and therefore it ends with an equilibrium state between two kinds of condensate. At each point in space, the $U(1)$ gauge field $B_{\mu}$ condenses in the form of $F^{\mu\nu}(x)F_{\mu\nu}(x)$. The free vacuum $|0\rangle$ in the right-hand side of Eq.(7) should be replaced by a condensed vacuum $|0_{r}\rangle$ satisfying $\langle 0_{r}|F^{\mu\nu}(x)F_{\mu\nu}(x)|0_{r}\rangle\equiv\langle\widehat{t}\rangle\neq 0$. This $\langle\widehat{t}\rangle$ is a kind of material constant of vacuum, and we should redefine $|0\rangle$ by $|0_{r}\rangle$ in the right-hand side of Eq.(7). (The explicit form of $|0_{r}\rangle$ and how it changes from protoworld to real world is to be studied in the future.)
(d) The vacuum is often probed using the operator-product expansion in the deep inelastic scattering, such as the pair annihilation of electron and positron to hadrons. For the cross section of this experiment, following vacuum condensates are assumed
$\sigma(e^{+}e^{-}\rightarrow hadrons)$
$=\frac{4\pi\alpha^{2}}{s}[\mbox{Im }c^{1}(q^{2})+\mbox{Im }c^{\bar{\varphi}\varphi}\langle 0|m\bar{\varphi}\varphi|0\rangle+\mbox{ Im }c^{F^{2}}(q^{2})\langle 0(F^{a}_{\alpha\beta})^{2}|0\rangle+\cdots].$
Such $\langle 0|m\bar{\varphi}\varphi|0\rangle$ and $\langle 0(F^{a}_{\alpha\beta})^{2}|0\rangle$ may play some role in symmetry breaking as well. The interpretation in this paper may have some implication for this guess.
(e) In the Higgs potential, $\lambda|v_{h}+h_{1}+ih_{2}|^{4}$ predicts the triple and quartic self-couplings of the Higgs particle $h_{1}$. More complex many-body effects than that in Figure 2 may correspond to such self couplings. So far, the agreement of the GWS model to experiments of electroweak interaction is satisfactory. When more precise measurements are performed, however, there is a possibility of deviation, especially for light quarks and leptons of the first and second generations. The above $\widetilde{L}(x)$ predicts some different results from those by the BEH model. The next subject is to extend it to the electroweak interaction.
Appendix A The vacuum $|\widetilde{0}\rangle$ satisfying $\widetilde{a}^{s}(\mbox{$p$})|\widetilde{0}\rangle=\widetilde{b}^{s}(-\mbox{$p$})|\widetilde{0}\rangle=0$
The vacuum satisfying $\widetilde{a}^{s}(\mbox{$p$})|\widetilde{0}\rangle=\widetilde{b}^{s}(-\mbox{$p$})|\widetilde{0}\rangle=0$ is as follows. In the expansion
$e^{-iK}Fe^{iK}=F+[-iK,F]+\frac{1}{2!}\left[-iK,[-iK,F]\right]+\cdots,$ (54)
we regard an operator $a^{s}(\mbox{$p$})$ or $b^{s}(-\mbox{$p$})$ as $F$, and the following operator as $K$,
$i\sum_{p,s}\theta_{\mbox{$p$}}[b^{s\dagger}(-\mbox{$p$})a^{s\dagger}(\mbox{$p$})-a^{s}(\mbox{$p$})b^{s}(-\mbox{$p$})].$ (55)
Hence, Eqs.(5) and (6) can be rewritten in a compact form
$\widetilde{a}^{s}(\mbox{$p$})=e^{-iK}a^{s}(\mbox{$p$})e^{iK},\quad\widetilde{b}^{s}(-\mbox{$p$})=e^{-iK}b^{s}(-\mbox{$p$})e^{iK}.$ (56)
The vacuum $|\widetilde{0}\rangle$ satisfying $\widetilde{a}^{s}(\mbox{$p$})|\widetilde{0}\rangle=\widetilde{b}^{s}(-\mbox{$p$})|\widetilde{0}\rangle=0$ is simply expressed as $|\widetilde{0}\rangle=e^{-iK}|0\rangle$. Hence, we obtain
$|\widetilde{0}\rangle$ $=$ $\exp\left(\sum_{p,s}\theta_{\mbox{$p$}}[b^{s\dagger}(-\mbox{$p$})a^{s\dagger}(\mbox{$p$})-a^{s}(\mbox{$p$})b^{s}(-\mbox{$p$})]\right)|0\rangle.$ (57)
$=$ $\prod_{p,s}\left[\sum_{n}\frac{1}{n!}\theta_{\mbox{$p$}}^{n}[b^{s\dagger}(-\mbox{$p$})a^{s\dagger}(\mbox{$p$})-a^{s}(\mbox{$p$})b^{s}(-\mbox{$p$})]^{n}\right]|0\rangle.$
[…]
$\sum_{n}\frac{\theta^{n}}{n!}(b^{\dagger}a^{\dagger}-ab)^{n}|0\rangle$ $=$ $|0\rangle+\theta b^{\dagger}a^{\dagger}|0\rangle-\frac{\theta^{2}}{2!}abb^{\dagger}a^{\dagger}|0\rangle-\frac{\theta^{3}}{3!}b^{\dagger}a^{\dagger}abb^{\dagger}a^{\dagger}|0\rangle$ (58)
$+$ $\frac{\theta^{4}}{4!}abb^{\dagger}a^{\dagger}abb^{\dagger}a^{\dagger}|0\rangle+\cdots.$
In this expansion, $\cos\theta_{\mbox{$p$}}$ appears in the sum of even-order terms of $\theta$, and $\sin\theta_{\mbox{$p$}}$ appears in the sum of odd-order terms, and then Eq.(7) is yielded.
Appendix B Statistical gap
(1) Consider a Bose field $f(x)$ with the following kinetic energy
$H_{ef}=-\int d^{3}xf^{\dagger}(x)\Delta f(x).$ (59)
The square of the infinitesimal line-element is a quadratic function of $dx$ as $dl^{2}=g_{\mu\nu}(x)dx^{\mu}dx^{\nu}$, where $g_{\mu\nu}(x)$ is a metric tensor ($\mu,\nu=x,y,z$). The inner product of the field $f(x)$ is defined as
$\langle f(x)|f(x)\rangle=\int\sqrt{g(x)}d^{3}xf^{\dagger}(x)f(x),$ (60)
where $g(x)$ is a determinant of $g_{\mu\nu}(x)$. Consider a gradient of $f$ like $A^{\mu}=g_{\mu\nu}\partial f/\partial x^{\nu}$. Since the metric depends on the position, a derivative of a given vector $A^{\mu}$ with respect to $x^{\mu}$ is replaced by the covariant derivative as follows
$\frac{DA^{\mu}}{dx^{\mu}}=\frac{dA^{\mu}}{dx^{\mu}}+\Gamma^{\mu}_{\nu\mu}A^{\nu},$ (61)
where $\Gamma^{\mu}_{\nu\mu}$ is a connection coefficient. This $\Gamma^{\mu}_{\nu\mu}$ is expressed by the determinant of the metric tensor $g(x)=|g_{\mu\nu}(x)|$ as follows
$\Gamma^{\mu}_{\nu\mu}=\frac{1}{2g}\frac{\partial g}{\partial x^{\nu}}.$ (62)
With this expression, the covariant derivative of $A^{\mu}$ is given by
$\frac{DA^{\mu}}{dx^{\mu}}=\frac{1}{\sqrt{g}}\frac{\partial(\sqrt{g}A^{\mu})}{\partial x^{\mu}}.$ (63)
(2) One can use Eq.(63) to study the Laplacian $D\partial f/\partial^{2}x$. The matrix element of the Laplacian operator is obtained by inserting $D\partial/\partial^{2}x$ between $f^{\dagger}(x)$ and $f(x)$ in Eq.(60). For $D\partial f/\partial^{2}x$, we use Eq.(63) with $A^{\mu}=g_{\mu\nu}\partial f/\partial x^{\nu}$. After integration by parts, we get
$\langle f(x)|-\Delta|f(x)\rangle$ $=$ $-\int\sqrt{g}d^{3}xf^{\dagger}\frac{1}{\sqrt{g}}\frac{\partial}{\partial x^{\mu}}\left(\sqrt{g}g_{\mu\nu}\frac{\partial f}{\partial x^{\nu}}\right)$ (64)
$=$ $\int\sqrt{g(x)}g_{\mu\nu}d^{3}x\frac{\partial f^{\dagger}}{\partial x^{\mu}}\frac{\partial f}{\partial x^{\nu}}.$
Our interest is how the effect of permutation symmetry looks in the flat space, because we observe it as an effect of Bose statistics in experiments. To rewrite Eq.(64), we introduce a new field $\widehat{f}(x)=g(x)^{1/4}f(x)$ as follows
$\int\sqrt{g(x)}d^{3}xf^{\dagger}(x)f(x)=\int d^{3}x\widehat{f}^{\dagger}(x)\widehat{f}(x).$ (65)
The gradient in Eq.(59) is rewritten using $\widehat{f}(x)$
$\frac{\partial f}{\partial x}=g^{-1/4}\left(\frac{\partial}{\partial x}-\frac{1}{4}\frac{\partial\ln g}{\partial x}\right)\widehat{f}(x).$ (66)
Hence, the matrix element of Laplacian in Eq.(64)
$\langle f(x)|-\Delta|f(x)\rangle=\int d^{3}xg_{\mu\nu}\left(\frac{\partial}{\partial x^{\mu}}-\frac{1}{4}\frac{\partial\ln g}{\partial x^{\mu}}\right)\widehat{f}^{\dagger}(x)\left(\frac{\partial}{\partial x^{\nu}}-\frac{1}{4}\frac{\partial\ln g}{\partial x^{\nu}}\right)\widehat{f}(x),$ (67)
is rewritten using the integration by parts as follows
$\langle f(x)|-\Delta|f(x)\rangle=\int d^{3}xg_{\mu\nu}\frac{\partial\widehat{f}^{\dagger}}{\partial x^{\mu}}\frac{\partial\widehat{f}}{\partial x^{\nu}}+\int W(x)\widehat{f}^{\dagger}(x)\widehat{f}(x)d^{3}x,$ (68)
where
$W(x)=\frac{1}{4}\frac{\partial}{\partial x^{\mu}}\left(g_{\mu\nu}\frac{\partial\ln g}{\partial x^{\nu}}\right)+\frac{1}{16}g_{\mu\nu}\left(\frac{\partial\ln g}{\partial x^{\mu}}\right)\left(\frac{\partial\ln g}{\partial x^{\nu}}\right).$ (69)
This $W(x)$ is a square of the finite energy gap in the excitation spectrum of $\widehat{f}(x)$ [18].