### v15

# Massive gauge boson and Higgs boson as the vestiges of non-Fock vacuum

Abstract
A microscopic model of the Brout-Englert-Higgs (BEH) mechanism is proposed. Massless fermions and antifermions do not belong to the Fock space with definite particle-number distribution, but belong to a non-Fock space with indefinite one. From this vast space, their ground-state is selected by a kinematical condition. Due to the interaction between them, the Fock state in which fermions and antifermions are massive is restored, but this state has the vestiges of the non-Fock state in the massive gauge boson and the Higgs boson. In the non-Fock state, massless fermions and antifermions exist as pairs, and behave as quasi bosons within a small space-time region. Due to Bose statistics, the direction of their center-of-mass motion is parallel to each other, and their transverse excitations are suppressed by an energy gap, making the gauge boson coupled to them massive. The Higgs boson is not an elementary particle, but a quasiparticle appearing after the Fock vacuum is restored.
1 Introduction
Spontaneous breaking of symmetry in vacuum is divided roughly into two categories. The first is the Brout-Englert-Higgs (BEH) mechanism [1][2], which is an extension of the Stueckelberg formalism [3]. The BEH mechanism is a simple model in which both of symmetry breaking and its consequence are derived by adding to $\frac{1}{4}F^{\mu\nu}F_{\mu\nu}$ the following gauge coupling and Higgs potential of a scalar field $h$,
$L_{h}(x)=|(i\partial_{\mu}+gB_{\mu})h|^{2}-\mu^{2}|h|^{2}-\lambda|h|^{4}.$ (1)
Switching the sign of $\mu^{2}$ leads to the broken-symmetry vacuum with the vacuum condensate $v_{h}$. This $v_{h}$ gives a mass $m_{B}=gv_{h}$ to the vector-Abelian-gauge field $B_{\mu}$. The above $h$ becomes $v_{h}+h_{1}+ih_{2}$ composed of the Higgs field $h_{1}$ and the Goldstone mode $h_{2}$. (Stueckelberg took only the Goldstone mode, and proposed a simple gauge-invariant model of massive gauge boson. )
For the electroweak interaction in which the weak interaction is effective only in the left-handed fermions, the Glashaw-Weinberg-Salam (GWS) model [4][5][6] used the BEH mechanism for the chiral-symmetric Lagrangian density $\frac{1}{4}F^{a\mu\nu}F^{a}_{\mu\nu}+\bar{\varphi}(i\partial_{\mu}+igA^{a}_{\mu}\tau_{a}+ig^{\prime}B_{\mu})\gamma^{\mu}\varphi$, and gave a finite mass $m_{f}$ to massless fermion $\varphi$ by adding the Yukawa interaction as follows
$L^{\prime}_{h}(x)=L_{h}(x)-\frac{m_{f}}{v_{h}}h\bar{\varphi}\varphi.$ (2)
This model predicted the W and Z bosons and the Higgs boson found in 1983 [7][8] and in 2012 [9][10], respectively, and does not contradict almost all experimental results to date. The Higgs potential explains many phenomena using a small number of parameters. This is because it plays a double role: the role of causing symmetry breaking in vacuum and that of predicting the mass of Higgs boson. Furthermore, it stabilizes the broken-symmetry vacuum, and it represents the interaction between the Higgs bosons. In this sense, the Higgs potential is a simple and economical model, and it also has the flexibility to adapt to complex situations of the eletroweak interaction. For its simplicity, however, we face some difficult problems. The switch of the sign of $\mu^{2}$ is an ad hoc assumption without explanation. The origin of the vacuum condensate $v_{h}$ is not clear. When the Higgs potential is used in the perturbation calculation, we must care a lot about the intricate cancellation of the quadratic divergence.
The second is dynamical symmetry breaking [11][12][13]. For a broken symmetry in our world, we can consider a symmetric world. Although all material particles have their own masses, massless material particles are considered. Beginning with the Fock vacuum, the interaction through gauge field makes this vacuum unstable, and forms a new vacuum, which belongs to a non-Fock representation at the infinite-volume limit. This is considered as a physical process which we can really observe in experiments. Phase transitions in the non-relativistic condensed matter belong to this category. To understand the mechanism behind the BEH mechanism, many attempts have been made using an analogy with them. This is a probable analogy, but to our knowledge, we do not yet clearly understand how such a symmetry breaking occurs in vacuum.
In order to understand the physics behind, we look at the problem from a different angle. Unlike massive fermions, massless fermions in the Fock state are questionable, because finite mass, which supports the Fock state from behind, is not there. So long as we begin with a chiral-symmetric Lagrangian, the non-Fock state is needed from the beginning, and it is not necessary to derive it dynamically from the Fock state. The order of this argument may be the reverse of the order of the logic in dynamical symmetry breaking. Kinematics of massless antifermion rather than dynamics determines the non-Fock state. There massless fermions acquire a mass by their interaction, and the Fock state is restored, but this state is different from the conventional Fock state in that the massive gauge boson and the Higgs boson exist as the vestiges of the non-Fock state. In this non-Fock state, massless fermion-antifermion pairs respond to the gauge field as quasi bosons, which causes massive gauge boson. It is experimentally confirmed that the fermion’s coupling to Higgs boson is proportional to fermion’s mass. If we take a step of approximation further in the mass acquisition of fermion, the Higgs-like boson will appear. From this viewpoint, the Higgs boson is not an elementary particle, but a quasiparticle appearing after the Fock vacuum is restored.
This paper is organized as follows. In Section 2, we reconsider massless fermion in the non-Fock state. For its ground state, a kinematical condition is used as a route to the non-Fock vacuum. Whereas the Nambu Jona-Lasinio state is recognized as such a vacuum, the massless fermion is an unstable existence by nature and becomes massive whatever interaction acts on them. In Section 3, the quasi-boson’s aspect of massless fermion-antifermion pair is explained. To describe them, some constants of vacuum are needed, and the $v_{h}$, $\mu$ and $\lambda$ in Eq.(1) are examples of such constants. The question is how to introduce these constants in a physically natural way. We define microscopic three constants of the non-Fock vacuum in position space. In Section.4, with these constants and quantum statistics, we propose an origin of the vacuum condensate $v_{h}$, and calculate the gauge-boson’s mass. In Section 5, we calculate Higgs mass $m_{H}$ as an excitation energy of a collective motion of massive fermion and antifermion mediated by massive gauge boson. Parameters in the Higgs potential are reproduced by these constants. In Section 6, we discuss some implications for the BEH mechanism.
2 Non-Fock state under chiral-symmetric Lagrangian
Let us examine the vacuum under the simple chiral-symmetric Lagrangian density,
$L_{0}(x)=-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+\bar{\varphi}(i\partial_{\mu}+gB_{\mu})\gamma^{\mu}\varphi,$ (3)
where $B_{\mu}$ is a U(1) gauge field, and $\varphi$ is a massless Dirac fermi field [14],
$\varphi(x)=\frac{1}{\sqrt{V}}\sum_{p,s}\left[a^{s}(\mbox{$p$})u^{s}(p)e^{-ipx}+b^{s\dagger}(\mbox{$p$})v^{s}(p)e^{ipx}\right].$ (4)
2.1 Massless fermion in the non-Fock state
In the case of massive fermions, a state is divided into subspaces with definite numbers of particles, the existence of which is supported by a finite mass. Fock state is a linear combination of basic vectors $\sum_{n}c_{n}|n_{1},n_{2},\cdots\rangle$ where $n=(n_{1},n_{2}\cdots)$ is a particle-number distribution. Whereas the degree of freedom $\lambda$ in $n_{\lambda}$ extends to infinity, only finite number of $n_{\lambda}$ is nonzero in $|n_{1},n_{2},\cdots\rangle$ of the Fock state, such as $\sum_{n}|c_{n}|^{2}<\infty$, and the sequence $n_{1}n_{2}\cdots$ represents a decimal. As a result, the number of basic vectors is countable. The operators $a^{s}(\mbox{$p$})$ and $b^{s\dagger}(\mbox{$p$})$, which strictly specifies the number of particle as a lowering or raising operator, is closely linked to Fock state. For the massless fermion, however, there is no reason for this rule to hold, because mass is no longer an indicator of individual particle, and momentum remains as a continuous parameter. To describe this parameter, nonzero sequence $n_{1}n_{2}\cdots$ extending to infinity should be prepared, and the number of basic vectors is uncountable as in the cardinality of real numbers. Such a state is not divided sharply into subspaces as in the Fock state [15]. The vacuum of massless fermion and antifermion is a non-Fock state with indefinite particle-number.
This indefinite particle-number is realized by the dynamics in $L_{0}(x)$ as follows. If we consider massless particles to be real entities, they become a subject of the theory of relativity, even if their velocity is not so large as that of light. Due to pair production from $B_{\mu}$, the number of massless fermion and antifermion is not fixed, because mass no longer gives a threshold energy in the pair production. The state with indefinite particle-number is preferable. The space expanded by uncountable basic vectors, however, is a vast state-space. Hence, we should proceed with the help of kinematical condition, because kinematics does not depend on the details of the interaction, and kinematics is more fundamental than dynamics in this case. For massless particle, a new raising and lowering operator reflecting this indefinite particle-number is needed.
2.2 Kinematical condition
For the gauge field, the moving massless fermions or antifermions are nothing but charge’s and spin’s vehicles. Since vacuum is neutral in charge and isotropic in space, the annihilation of negative charge in $a^{s}(\mbox{$p$})$ at the departure of massless fermion brings about the same effect to the gauge field $B_{\mu}(\mbox{$p$})$ as the creation of positive charge in $b^{s\dagger}(-\mbox{$p$})$ at the arrival of massless antifermion with opposite momentum and spin. Furthermore, momentum, electric charge and spin, which prescribe all properties of massless fermion, are not positive definite, and therefore the effect of $a^{s}(\mbox{$p$})$ and that of $b^{s\dagger}(-\mbox{$p$})$ on the state are equivalent. There is no reason to assign the equivalent operations on the state to two different operators. For this to be ensured, a new raising and lowering operator $\widetilde{a}^{s}(\mbox{$p$})$ is defined as follows
$\widetilde{a}^{s}(\mbox{$p$})=\cos\theta_{\mbox{$p$}}a^{s}(\mbox{$p$})+\sin\theta_{\mbox{$p$}}b^{s\dagger}(-\mbox{$p$}).$ (5)
If the fermion is massive, the annihilation of fermion and the creation of antifermion cause different effects on the state because mass is positive definite, and such a raising and lowering operator does not exist. Hence, it is a characteristic of the massless fermion and antifermion.
The necessity of antifermion in Eq.(5) changes according to the relative momentum between the particle and the observer, which is reflected in $\sin\theta_{\mbox{$p$}}$. When $\mbox{$p$}=0$, the difference between $a^{s}(\mbox{$p$})$ and $b^{s\dagger}(-\mbox{$p$})$ in momentum vanishes. Therefore, the operators $a^{s}(\mbox{$p$})$ and $b^{s\dagger}(-\mbox{$p$})$ have the same importance for the observer, resulting in $\cos\theta_{\mbox{$p$}}=\sin\theta_{\mbox{$p$}}$ at $\mbox{$p$}=0$. On the contrary, when $\mbox{$p$}^{2}\rightarrow\infty$, such observers cannot be found, and antifermion is not needed. Hence, $\sin\theta_{\mbox{$p$}}\rightarrow 0$ is expected. The situation is same also for the antifermion. A new operator $\widetilde{b}^{s}(-\mbox{$p$})$ is defined as a superposition of the annihilation of massless antifermion and the creation of massless fermion
$\widetilde{b}^{s}(-\mbox{$p$})=\cos\theta_{\mbox{$p$}}b^{s}(-\mbox{$p$})-\sin\theta_{\mbox{$p$}}a^{s\dagger}(\mbox{$p$}).$ (6)
This $\widetilde{b}^{s}(-\mbox{$p$})$ is orthogonal to $\widetilde{a}^{s}(-\mbox{$p$})$.
[…]
$|\widetilde{0}\rangle=\prod_{p,s}\left[\cos\theta_{\mbox{$p$}}+\sin\theta_{\mbox{$p$}}e^{i\alpha(x)}b^{s\dagger}(-\mbox{$p$})a^{s\dagger}(\mbox{$p$})\right]|0\rangle.$ (7)
This $|\widetilde{0}\rangle$ is an example of non-Fock states in which the number of basic vector is uncountable as represented by a continuous weight-factor $\sin\theta_{\mbox{$p$}}$, which is selected by the kinematical condition among various candidates in the vast state-space.
This $|\widetilde{0}\rangle$ was first introduced to elementary-particle physics by [11] in analogy with superconductivity [16]. However, it is not a form specific to this analogy, but the first approximation of the general form for the relativistic massless many-body state [17]. The analogy with superconductivity has the following limits. In metals, the momentum of electrons represents the relative motion of electrons to the center-of-mass of crystal. In the physical vacuum, the naive analogy such as the center-of-mass of world cannot be carried over, because the significant momentum is only that of relative motion between observer and particle. The derivation via Eqs.(5) and (6), which uses the relative momentum between them, is a natural way to describe physical vacuum [18].
2.3 Restoration of Fock vacuum
The kinematical condition, rather than dynamics, leads to the non-Fock vacuum $|\widetilde{0}\rangle$. However, particles in interacting with others cannot exist as massless ones unless they are protected by some stronger principle. The dynamics in $L_{0}(x)$ this time prevents the exact realization of such a non-Fock vacuum. The restoration of the Fock vacuum occurs in the following way.
Massless fermions and antifermions interact to each other via $B_{\mu}$. The interaction between them, regardless of being repulsive or attractive, induces the motion of other massless fermions and antifermion. As a result, the motion of original fermions and antifermion is not easily affected by the external force, which comes back to the massless fermions and antifermion as an inertial mass. The simplest way to demonstrate this mechanism is a mean-field approximation. The mean field is an approximation that neglects the long-range fluctuation. As long as the gauge boson is massless, the interaction between fermions is a long-range one, which induces the long-range longitudinal excitation. Such excitations involve a large amount of particles and require a high excitation-energy (like the plasmon in the electron gas). Hence, the long-range fluctuation is suppressed, and the mean field is an useful way. Furthermore, the restoration of the Fock vacuum is related to the stability of the world. For this world to continue to exist forever, such a process should depend, not on the time-dependent dynamics, but on the time-independent one that is well described by the mean field [19]. Assuming a scalar mean-field $U_{0}$, we consider
$\bar{\varphi}(x)(i/\penalty 50{\!\!\!}{\partial}+U_{0})\varphi(x).$ (8)
Following the procedure in [16], the condition that Eq.(8) is diagonal with respect to $\widetilde{a}^{s}(\mbox{$p$})$ and $\widetilde{b}^{s}(\mbox{$p$})$ is as follows
$\cos^{2}\theta_{\mbox{$p$}}=\frac{1}{2}\left(1+\frac{\epsilon_{p}}{\sqrt{\epsilon_{p}^{2}+U_{0}^{2}}}\right),\quad\sin^{2}\theta_{\mbox{$p$}}=\frac{1}{2}\left(1-\frac{\epsilon_{p}}{\sqrt{\epsilon_{p}^{2}+U_{0}^{2}}}\right),$ (9)
which satisfies the kinematical requirement $\cos\theta_{\mbox{$p$}}=\sin\theta_{\mbox{$p$}}$ at $\mbox{$p$}=0$, and $\sin\theta_{\mbox{$p$}}\rightarrow 0$ at $\mbox{$p$}\rightarrow\infty$ that we imposed in Eqs.(5) and (6). The diagonalized form of Eq.(8) includes $\sqrt{\epsilon_{p}^{2}+U_{0}^{2}}[\widetilde{a}^{s\dagger}(\mbox{$p$})\widetilde{a}^{s}(\mbox{$p$})+\widetilde{b}^{s\dagger}(\mbox{$p$})\widetilde{b}^{s}(\mbox{$p$})]$, in which $U_{0}^{2}$ is a square of the mass $m_{f}^{2}$ of the fermion and antifermion. Here, the sign of $U_{0}$ does not affect the result, hence it does not depend on whether $U_{0}$ in Eq.(8) is attractive or repulsive. If we regard this symmetry breaking as a direct analogue of superconductivity, such as the technicolor model, the attractive interaction is a necessary assumption to derive the non-Fock vacuum. However, because the non-Fock vacuum is a logical consequence of the zero mass, such an assumption is not necessary.
The restored Fock vacuum is different from the conventional Fock vacuum of the massive fermion in that the vestige of the non-Fock vacuum remains in the massive gauge boson and the Higgs boson.
(a) The massless fermion and antifermion form a pair in $|\widetilde{0}\rangle$, and these pairs obey Bose statistics within a small space-time region after pair production. Due to Bose statistics, an energy gap appears in the transverse excitation of these pairs, which makes this state robust against transverse disturbance, and makes the gauge boson massive when it couples to these pairs. This will give us an origin of the vacuum condensate $v_{h}$ in the Higgs potential, as will be explained in Section 3 and 4.
(b) When we proceed to the beyond-mean-field approximation in Eq.(8), the Higgs particle appears, as will be explained in Section 5.
3 Vacuum condensate
3.1 Massless fermion-antifermion pairs as quasi bosons
In the BEH model, the vacuum condensate $v_{h}$ in $L_{h}(x)$ implies that vacuum is robust against perturbation. This robustness should be physically explained by the massless fermion and antifermion in Eq.(7). When they appear or disappear in $|\widetilde{0}\rangle$, it is always as a pair with mutually opposite momenta and spins. For short periods after pair production, their fields should overlap in position space, and their behavior is represented by dimensionless composite operators $P_{\mbox{$k$}}\equiv b(-\mbox{$k$},\downarrow)a(\mbox{$k$},\uparrow)$ and $P^{\dagger}_{\mbox{$k$}}\equiv a^{\dagger}(\mbox{$k$},\uparrow)b^{\dagger}(-\mbox{$k$},\downarrow)$. ($\uparrow,\downarrow$ denote spins.) Owing to the anti-commutation relation, the composite operator $P_{\mbox{$k$}}$ has the following equal-time commutation relation at $\mbox{$k$}\neq\mbox{$k$}^{\prime}$
$P_{\mbox{$k$}},P^{\dagger}_{\mbox{$k$}^{\prime}}]=0\quad\mbox{for}\quad\mbox{$k$}\neq\mbox{$k$}^{\prime},$ (10)
implying that massless fermions-antifermions pairs with different momentum follow Bose statistics. However when $\mbox{$k$}=\mbox{$k$}^{\prime}$, due to Pauli principle, $P_{\mbox{$k$}}$ shows the following commutation relation
$P_{\mbox{$k$}},P^{\dagger}_{\mbox{$k$}}]=1-(n_{\mbox{$k$},\uparrow}+n_{-\mbox{$k$},\downarrow}),\quad P_{\mbox{$k$}}^{2}=P^{\dagger 2}_{\mbox{$k$}}=0,$ (11)
where $n_{\mbox{$k$},\uparrow}=a^{\dagger}(\mbox{$k$},\uparrow)a(\mbox{$k$},\uparrow)$ and $n_{-\mbox{$k$},\downarrow}=b^{\dagger}(-\mbox{$k$},\downarrow)b(-\mbox{$k$},\downarrow)$. This is a hybrid of boson’s one $P_{\mbox{$k$}}P^{\dagger}_{\mbox{$k$}}=N_{\mbox{$k$}}+1$ and fermion’s one $P_{\mbox{$k$}}P^{\dagger}_{\mbox{$k$}}=1-(n_{\mbox{$k$},\uparrow}+n_{-\mbox{$k$},\downarrow})$. These commutation relations show that the center-of-mass motion of the pair denoted by $P_{\mbox{$k$}}$ and its response to $B_{\mu}$ are that of a quasi boson.
3.2 Relativistic quantum field in position space
In relativistic quantum field theory, position is complementary not only to momentum, but also to particle number. If we make a precise measurement of position, it causes a wide spread of momentum, and owing to subsequent pair production, the number of particles becomes indefinite. However, position of particle is not a completely meaningless concept. In the measurement of position of moving particle with an energy $\epsilon$, the time duration necessary for measurement using light $\delta t\simeq\Delta x/c$ satisfies $\delta t>\hbar/c\Delta p>\hbar/cp\geq\hbar/\epsilon$. Since the particle may move $c\delta t>\hbar c/\epsilon$ during this interval, the least possible error of position $\delta x$ is $\hbar c/\epsilon$ [20]. Unless we attempt to specify the position of particle more precisely than $\delta x$, it has a physical meaning. Quasi bosons appear and disappear randomly in position space, but we consider for simplicity a mean distance between quasi bosons $d_{m}$, which is larger than $\delta x$. Figure 1 schematically illustrates the distribution of these quasi bosons in position space at a moment, using small white circles being blurred with the uncertainty $\delta x=\hbar c/\epsilon$. The gauge field $B_{\mu}$ propagates along the horizontal line. Since quasi bosons made of fermion and antifermion with different momentum prevail in vacuum, they follow Bose statistics as in Eq.(10). However, since these quasi bosons are not stable, their behavior as bosons is a limited one in the temporal and spatial sense. Here we define a coherence length $l_{c}$ in the vacuum such as $\delta x<d_{m}<l_{c}$ so that as long as the spatial distance between these quasi bosons is less than $l_{c}$, they obey Bose statistics.
Figure 1: A schematic view of the physical vacuum at a moment in position space. Each white circle blurred with $\delta x=\hbar c/\epsilon$ represents a quasi-boson, which is distributed with a mean distance $d_{m}$. (a) The transverse displacements (drdr) of quasi bosons from the horizontal line ($z$ axis) are coupled to the gauge field $B_{\mu}$ propagating along the $z$ axis, and (b) the longitudinal displacements (dzdz) are illustrated.
3.3 Statistical gap
Consider quasi bosons using a composite scalar field $f(x)$ made of $\varphi(x)$ in Eq.(4)
$f(x)=\frac{1}{\sqrt[3]{V}}\sum_{k,s}[P_{\mbox{$k$}}\bar{v}^{s}(-k)u^{s}(k)+P^{\dagger}_{-\mbox{$k$}}\bar{u}^{s}(-k)v^{s}(k)]e^{ikx}.$ (12)
(The normalization volume $V$ in $\varphi(x)$ and $f(x)$ is $d_{m}^{3}$.) This $f(x)$ describes the excitations of bosons within the coherence length, which has the following kinetic energy
$H_{ef}=\int\left|\frac{\partial}{\partial x_{i}}f(x)\right|^{2}d^{3}x=-\int f^{\dagger}(x)\Delta f(x)d^{3}x.$ (13)
In Figure 1(a), quasi bosons are excited by the gauge field $B_{\mu}$ from the solid horizontal line to a dotted wavy line. Let us consider their kinetic energy. Bose statistics requires permutation symmetry for quasi bosons, and therefore displacements pointed by a long white arrow from the horizontal line to black circles in Figure 1(a) is indistinguishable from that by small arrows between a black circle and a neighboring white circle. Because the transverse displacements are perpendicular to the wave vector, they are not additive, and there are many short displacements in the excitation. The long transverse displacements can be replaced by the many short ones. (In contrast, since the longitudinal displacement is additive as shown in Figure 1(b), the long displacements are dominant, and the short displacements are exceptional.) The short distance in the denominator of Laplacian operator $\Delta$ in Eq.(13) creates a high excitation energy, and an energy gap $\epsilon_{0}$ from zero appears in the transverse excitation spectrum. This energy gap comes only from statistical property, and we call it statistical gap [21]. The long movements are substantially forbidden in the transverse motion of quasi bosons.
This situation is expressed using a metric tensor $g_{ij}(x)$ in differential geometry. For the transverse and longitudinal distances $r$ and $z$ respectively in Figure 1, the square of the line element is expressed as $dl^{2}=g_{rr}(r)dr^{2}+dz^{2}$. When the quasi bosons in Figure 1(a) move in the transverse direction within the coherent spatial region with a size $l_{c}$, a small distance $dr$ such as $d_{m}$ causes the same effect to $dl^{2}$ as a large distance $r$. A simple metric representing this situation is given by
$g_{rr}(r)$ $=$ $\frac{r^{2}}{d_{m}^{2}},\quad(d_{m}\leq r<l_{c}),$
$g_{rr}(r)$ $=$ $1,\qquad(0<r<d_{m},\quad l_{c}\leq r).$ (14)
The gradient in Eq.(13) is rewritten as follows (see Appendix.B)
$H_{ef}=\int g_{\mu\nu}\frac{\partial\widehat{f}^{\dagger}}{\partial x^{\mu}}\frac{\partial\widehat{f}}{\partial x^{\nu}}d^{3}x+\int W(x)\widehat{f}^{\dagger}(x)\widehat{f}(x)d^{3}x,$ (15)
where a normalized field $\widehat{f}(x)\equiv|g(x)|^{1/4}f(x)$ has, in addition to the kinetic energy, a square of the finite energy as follows ($\mu,\nu=x,y,z$) [22]
$W(x)=\frac{1}{4}\frac{\partial}{\partial x^{\mu}}\left(g_{\mu\nu}\frac{\partial\ln|g|}{\partial x^{\nu}}\right)+\frac{1}{16}g_{\mu\nu}\left(\frac{\partial\ln|g|}{\partial x^{\mu}}\right)\left(\frac{\partial\ln|g|}{\partial x^{\nu}}\right),$ (16)
where $|g(x)|$ is a determinant of $g_{\mu\nu}(x)$. Using Eq.(14) and $g_{zz}=1$, $W(r)$ of each quasi boson is given by
$W(r)=\frac{3}{4}\frac{1}{d_{m}^{2}}.$ (17)
Each quasi boson should jump an energy barrier $\epsilon_{0}\equiv\sqrt{W(x)}$ when it is transversely excited. The number of quasi bosons in the transverse plane within the coherence length is $\pi(l_{c}/d_{m})^{2}$. The sum of $\epsilon_{0}$ over quasi bosons in this area is given by
$\widehat{\epsilon}_{0}=\frac{\sqrt{3}}{2}\pi\left(\frac{l_{c}}{d_{m}}\right)^{2}\frac{1}{d_{m}}.$ (18)
Owing to this gap $\widehat{\epsilon}_{0}$, the state described by $\widehat{f}(x)$ remains in the ground state even if it is transversely perturbed, and leads to $\langle\widetilde{0}|\partial_{\mu}\widehat{f}(x)|\widetilde{0}\rangle=0$ within the coherent spatial region. Hence, the state described by $\varphi(x)$ also remains in the ground state, leading to $\langle\widetilde{0}|\partial_{\mu}[\bar{\varphi}(x)\gamma^{\mu}\varphi(x)]|\widetilde{0}\rangle=0$ for the transverse direction. This exhibits a kind of rigidity of the physical vacuum, which is the origin of the vacuum condensate $v_{h}$ in $|(i\partial_{\mu}+gB_{\mu})(v_{h}+h_{1}+ih_{2})|^{2}$ in Eq.(1).
4 Massive gauge boson
In the BEH model, the mass term of gauge boson $m_{B}^{2}B^{\mu}B_{\mu}=g^{2}v_{h}^{2}B^{\mu}B_{\mu}$ is derived from the phenomenological coupling $|(i\partial_{\mu}+gB_{\mu})(v_{h}+h_{1}+ih_{2})|^{2}$. In this paper, we derive this term microscopically. The physical vacuum is not a simple system, and therefore the response of $|\widetilde{0}\rangle$ to $B_{\mu}$ includes a non-linear effect. The minimal coupling of fermions $L_{0}^{min}(x)=\bar{\varphi}(x)(i\partial_{\mu}+gB_{\mu})\gamma^{\mu}\varphi(x)$ changes its form in $|\widetilde{0}\rangle$ by the perturbation of ${\cal H}_{I}(x)=gj^{\mu}(x)B_{\mu}(x)$ [23]. Consider a perturbation expansion of $\int d^{4}xL_{0}^{min}(x)$ in powers of $g$
$\langle\widetilde{0}|\int d^{4}x_{1}L_{0}^{min}(x_{1})exp\left(i\int{\cal H}_{I}(x_{2})d^{4}x_{2}\right)|\widetilde{0}\rangle$ (19)
$=\langle\widetilde{0}|\int d^{4}x_{1}\bar{\varphi}(x_{1})\gamma^{\mu}[i\partial_{\mu}+gB_{\mu}(x_{1})]\varphi(x_{1})|\widetilde{0}\rangle$
$+\langle\widetilde{0}|\int d^{4}x_{1}\bar{\varphi}(x_{1})\gamma^{\mu}[i\partial_{\mu}+gB_{\mu}(x_{1})]\varphi(x_{1})ig\int d^{4}x_{2}j^{\nu}(x_{2})B_{\nu}(x_{2})|\widetilde{0}\rangle$$\cdots$,
4.1 Gauge-boson’s mass due to statistical gap
(1) In the last term of Eq.(19), $B_{\nu}(x_{2})$ is correlated to $B_{\mu}(x_{1})$, yielding the following two-point-correlation function between $j^{\mu}(x_{1})$ and $j^{\nu}(x_{2})$
$g^{2}\int\langle\widetilde{0}|\int j_{\mu}(x_{1})d^{2}x_{1}\int j^{\nu}(x_{2})d^{2}x_{2}|\widetilde{0}\rangle B^{\mu}(x_{1})B_{\nu}(x_{2})d^{2}x_{1}d^{2}x_{2}$ (20)
Since these $x_{1}$ and $x_{2}$ are separated microscopically, a distant observer regards it as a local phenomenon at $X=(x_{1}+x_{2})/2$. For such an observer, it is useful to rewrite $d^{4}x_{1}d^{4}x_{2}$ in Eq.(20) as $d^{4}Xd^{4}Y$.
$g^{2}\int\langle\widetilde{0}|\int j_{\mu}(Y)j^{\nu}(0)d^{4}Y|\widetilde{0}\rangle\times B^{\mu}(X)B_{\nu}(X)d^{4}X$ (21)
The relative motion along $Y=x_{2}-x_{1}$ is indirectly observed in the coefficient of $B^{\mu}(X)B_{\nu}(X)$
(2) Because the gauge field is a transverse one, the excitation of fermions induced by this $B_{\mu}(x)$ is a transverse one as well. The correlation of currents in Eq.(21) implies a correlation of the center-of-mass motions of different massless fermion-antifermion pairs. Because they are strongly influenced by Bose statistics, quasi bosons move parallel to each other coherently in the response to $B_{\mu}$, and correlation of different currents is nonzero only in the same direction as follows
$\langle\widetilde{0}|j_{\mu}(Y)j^{\nu}(0)|\widetilde{0}\rangle \displaystyle\Rightarrow$ $\langle\widetilde{0}|j_{\mu}(Y)j^{\mu}(0)|\widetilde{0}\rangle\quad for\quad\mu=\nu.$
$\langle\widetilde{0}|j_{\mu}(Y)j^{\nu}(0)|\widetilde{0}\rangle \displaystyle\Rightarrow$ $0\qquad\qquad\qquad\qquad for\quad\mu\neq\nu.$ (22)
Figure 2: (a) In the non-Fock world, the gauge boson in $B_{\mu}$ (a wavy line) induces vacuum polarization of massless fermion and antifermion only outside the coherent space-time region specified by $l_{c}$ (a shaded area). (b) In the real world, the chain of creation and annihilation of massive fermion and antifermion constitutes the Higgs-like excitation $H(x)$, which is induced both inside and outside this region .
Within the coherence length $l_{c}$, $\varphi(x)$ in the current remains in the ground state, and $\langle\widetilde{0}|\partial_{\mu}[\bar{\varphi}(Y)\gamma^{\mu}\varphi(Y)]|\widetilde{0}\rangle=0$ holds for the transverse component of $Y$. The correlation of currents in Eq.(21) is reduced to
$\langle\widetilde{0}|j_{\mu}(Y)j^{\mu}(0)|\widetilde{0}\rangle$ $=$ $\langle\widetilde{0}|\left(j_{\mu}(0)+\left[\frac{\partial}{\partial Y_{\mu}}[\bar{\varphi}(Y)\gamma_{\mu}\varphi(Y)]\right]_{Y_{\mu}=0}Y^{\mu}+\cdots\right)j^{\mu}(0)|\widetilde{0}\rangle$ (23)
$\Rightarrow$ $\langle\widetilde{0}|j_{\mu}(0)j^{\mu}(0)|\widetilde{0}\rangle$.
Using $j^{\mu}(0)=(\varphi^{\dagger}\varphi,i\varphi^{\dagger}\gamma^{0}\mbox{$\gamma$}\varphi)$, we obtain $\langle\widetilde{0}|j_{0}(0)j^{0}(0)|\widetilde{0}\rangle=\langle\widetilde{0}|j_{1}(0)j^{1}(0)|\widetilde{0}\rangle=\langle\widetilde{0}|j_{2}(0)j^{2}(0)|\widetilde{0}\rangle=\langle\widetilde{0}|j_{3}(0)j^{3}(0)|\widetilde{0}\rangle=\langle\widetilde{0}|[\varphi^{\dagger}(0)\varphi(0)]^{2}|\widetilde{0}\rangle$. Hence, Eq.(21) is interpreted as the mass term of gauge boson even after the Fock vacuum is restored
$\frac{1}{2}m_{B}^{2}\int B^{\mu}(X)B_{\mu}(X)d^{4}X,$ (24)
where $m_{B}^{2}$ is given by
$m_{B}^{2}=2g^{2}\langle\widetilde{0}|\int_{Y\in Z_{c}}[\varphi^{\dagger}(0)\varphi(0)]^{2}d^{4}Y|\widetilde{0}\rangle$. (25)
The coherent space-time region $Z_{c}$ in Eq.(25) is the inside of a small light-cone as illustrated in Figure 2 for one spatial direction. Its spatial length is less than $l_{c}$, and time width is smaller than $l_{c}/c$, so that the causal relation is possible between two spatial ends separated by $l_{c}$. The 4-dimensional volume of coherent space-time region $\int_{Y\in Z_{c}}d^{4}Y=l_{c}^{2}\times\frac{1}{2}[l_{c}\times c(l_{c}/c)]\times 2=l_{c}^{4}$ in Figure 2 is Lorentz invariant. Since transverse excitations are suppressed in $Z_{c}$, the vacuum polarization induced by $B_{\mu}$ occurs only outside this region as shown in Figure 2(a).
The $\varphi(x)$ in Eq.(25) has a normalization volume $V=d_{m}^{3}$ as in Eq.(4). Every fermion and antifermion in $|\widetilde{0}\rangle$ make the same contribution to $\langle\widetilde{0}|[\varphi^{\dagger}(0)\varphi(0)]^{2}|\widetilde{0}\rangle$ as follows
$\langle\widetilde{0}|\frac{1}{d_{m}^{6}}\sum_{p,s}\left([a^{s\dagger}(\mbox{$p$})u^{s\dagger}(p)+b^{s}(-\mbox{$p$})v^{s\dagger}(-p)][a^{s}(\mbox{$p$})u^{s}(p)+b^{s\dagger}(-\mbox{$p$})v^{s}(-p)]\right)^{2}|\widetilde{0}\rangle$ (26)
$=$ $\frac{1}{d_{m}^{6}}\langle\widetilde{0}|\sum_{p,s}[b^{s}(-\mbox{$p$})b^{s\dagger}(-\mbox{$p$})]^{2}+\sum_{p,s}[a^{s\dagger}(\mbox{$p$})a^{s}(\mbox{$p$})]^{2}$
$+$ $\sum_{p,s}b^{s}(-\mbox{$p$})a^{s}(\mbox{$p$})a^{s\dagger}(\mbox{$p$})b^{s\dagger}(-\mbox{$p$})+\sum_{p,s}a^{s\dagger}(\mbox{$p$})b^{s\dagger}(-\mbox{$p$})b^{s}(-\mbox{$p$})a^{s}(\mbox{$p$})\,|\widetilde{0}\rangle$
$=$ $2\times\frac{1}{d_{m}^{6}}\prod_{p,s}(\cos^{2}\theta_{\mbox{$p$}}+\sin^{2}\theta_{\mbox{$p$}})=\frac{2}{d_{m}^{6}}.$
where $u^{s\dagger}(p)u^{s}(p)=v^{s\dagger}(-p)v^{s}(-p)=1$. The gauge boson’s mass for $l_{c}^{4}$ of $Z_{c}$ in Eq.(25) is given by
$m_{B}^{2}=g^{2}\frac{4}{d_{m}^{6}}\int_{Y\in Z_{c}}d^{4}Y=g^{2}\left(\frac{2l_{c}^{2}}{d_{m}^{3}}\right)^{2}.$ (27)
This $2l_{c}^{2}/d_{m}^{3}$ is the origin of vacuum condensate $v_{h}$ in $m_{B}^{2}=g^{2}v_{h}^{2}$ of the BEH model, which is about the same as $\widehat{\epsilon}_{0}$ in Eq.(18).
4.2 Goldstone mode
(1) The Goldstone mode exists in the last term of Eq.(19). For the first-order term of $B_{\nu}(x_{2})$ coupled to $\bar{\varphi}(x_{1})\gamma_{\mu}i\partial^{\mu}\varphi(x_{1})$, integrate it over $x_{1}$ by parts. Since $\varphi(x_{1})$ vanishes at $x_{1}\rightarrow\infty$, we obtain two types of terms, one including $i\partial^{\mu}\bar{\varphi}(x_{1})\gamma_{\mu}\varphi(x_{1})$, and the other including $\partial^{\mu}|\widetilde{0}\rangle$. The latter is given by
$g\langle\widetilde{0}|\int d^{4}x_{1}j_{\mu}(x_{1})\int d^{4}x_{2}j^{\nu}(x_{2})B_{\nu}(x_{2})\partial^{\mu}|\widetilde{0}\rangle$ (28)
$+$ $g\partial^{\mu}\langle\widetilde{0}|\int d^{4}x_{1}j_{\mu}(x_{1})\int d^{4}x_{2}j^{\nu}(x_{2})B_{\nu}(x_{2})|\widetilde{0}\rangle$.
Because the vacuum $|\widetilde{0}\rangle$ in Eq.(7) has an explicit $x$-dependence in the phase $\alpha(x)$, $\partial^{\mu}|\widetilde{0}\rangle$ contains $\partial^{\mu}\alpha(x)$. The distant observer regards Eq.(28) as representing a local phenomenon at $X$, and rewrites it using $d^{4}x_{1}d^{4}x_{2}=d^{4}Xd^{4}Y$. After extracting $m_{B}^{2}$ from Eq.(28) as in Eq.(21), we obtain
$\frac{i}{2g}m_{B}^{2}\int B_{\mu}(X)\partial^{\mu}\alpha(X)d^{4}X\equiv\frac{m_{B}}{\sqrt{2}}\int B_{\mu}(X)\partial^{\mu}G(X)d^{4}X.$ (29)
Here the Goldstone mode is defined as $G(X)\equiv i(\sqrt{2}g)^{-1}m_{B}\alpha(X)$. This definition takes over the bound-state realization of the Goldstone mode from [11]. (While the BEH model derives the coupling of the Goldstone mode $h_{2}$ to $B_{\mu}$ from $|(i\partial_{\mu}+gB_{\mu})(v_{h}+h_{1}+ih_{2})|^{2}$, the above coupling comes from the response of the vacuum $|\widetilde{0}\rangle$ to $B_{\mu}$.)
(2) In the system without the long-range force, the global phase-rotation of $\alpha(X)$ requires no energy, and therefore the propagator of the Goldstone mode is given by
$\int\frac{d^{4}X}{(2\pi)^{4}}\langle\widetilde{0}|T[G(X)G(0)]|\widetilde{0}\rangle e^{iqX}=\frac{i}{q^{2}}.$ (30)
However, the long-range force mediated by the gauge boson prevents the free rotation of the global phase, then preventing the Goldstone mode. This discrepancy is resolved by the generation of the gauge-boson’s mass that converts the long-range force into a short-range one. The Fourier transform of Eqs.(24) and (29) are given by $(m_{B}^{2}/2)B^{\mu}(q)B_{\mu}(q)$ and $(m_{B}/\sqrt{2})q^{\mu}G(q)B_{\mu}(q)$, respectively. Following the usual way, regard the latter as a perturbation to the former, and the second-order perturbation is obtained as
$\frac{1}{2}B^{\mu}(q)\left[im_{B}^{2}g^{\mu\nu}-m_{B}q^{\mu}\frac{i}{q^{2}}m_{B}q^{\nu}\right]B_{\nu}(q)=\frac{i}{2}m_{B}^{2}\left(g^{\mu\nu}-\frac{q^{\mu}q^{\nu}}{q^{2}}\right)B^{\mu}(q)B^{\nu}(q).$ (31)
Adding Eq.(31) to the Fourier transform of $\frac{1}{4}F^{\mu\nu}F_{\mu\nu}$, we obtain an inverse of the resulting matrix,
$D^{\mu\nu}(q)=\frac{-i}{q^{2}-m_{B}^{2}}\left(g^{\mu\nu}-\frac{q^{\mu}q^{\nu}}{q^{2}}\right)\equiv iD(q^{2})\left(g^{\mu\nu}-\frac{q^{\mu}q^{\nu}}{q^{2}}\right),$ (32)
which is the propagator of the massive gauge boson in the Landau gauge. The response of the physical vacuum gives additional terms to $L_{0}(x)$ as $(m_{B}^{2}/2)B^{\mu}(q)B_{\mu}(q)$, $(m_{B}/\sqrt{2})B_{\mu}(x)\partial^{\mu}G(x)$ and $(\partial_{\mu}G(x))^{2}$.
(3) In the BEH model, the coupling of the Goldstone mode to fermions comes explicitly from the Yukawa interaction. In the present model, such a coupling appears implicitly in the first term of the right-hand side in Eq.(19). For the zeroth-order term of $B_{\mu}$, the integration over $x_{1}$ by parts yields two types of terms, one including $i\partial^{\mu}\bar{\varphi}(x_{1})\gamma_{\mu}\varphi(x_{1})$, and the other including $\partial^{\mu}|\widetilde{0}\rangle$. The latter term is given by
$i\langle\widetilde{0}|\int d^{4}x_{1}j^{\mu}(x_{1})\partial_{\mu}|\widetilde{0}\rangle+i\partial_{\mu}\langle\widetilde{0}|\int d^{4}x_{1}j^{\mu}(x_{1})|\widetilde{0}\rangle$. (33)
Since $\partial_{\mu}|\widetilde{0}\rangle=\partial_{\mu}\alpha(x)|\widetilde{0}\rangle$ contains the Goldstone mode $G(x)=i(\sqrt{2}g)^{-1}m_{B}\alpha(x)$, Eq.(33) gives the coupling of the Goldstone mode $G$ to fermion
$\frac{\sqrt{2}g}{m_{B}}\langle\widetilde{0}|\int d^{4}x_{1}\bar{\varphi}(x_{1})\gamma^{\mu}\varphi(x_{1})\partial_{\mu}G(x_{1})|\widetilde{0}\rangle$. (34)
Here, a new term $(\sqrt{2}g/m_{B})\bar{\varphi}(x)\gamma^{\mu}\varphi(x)\partial_{\mu}G(x)$ is added to $L_{0}(x)$, which is different from $g(m_{f}/m_{B})\bar{\varphi}(x)\varphi(x)h_{2}(x)$ led by Yukawa coupling in the BEH model.
5 Higgs boson
In the BEH model, the Higgs boson appears in a strange manner, that is, $\mu^{2}|h|^{2}$ in $L_{h}(x)$ of Eq.(1) plays double roles. The first is the generation of the broken-symmetry vacuum by switching the sign of $\mu^{2}$. The second is giving the Higgs boson a mass $m_{H}$. However, the former is concerned with the global property of the world, and the latter is concerned with one-particle property. It seems strange that such different scale of things are described by the same parameter. In the present model, since the vacuum $|\widetilde{0}\rangle$ has the kinematical origin, we have an option to assume a possible interaction for deriving $m_{H}$ without caring about symmetry breaking.
One of the important predictions of the BEH model for the electroweak interaction is that the strength of the Higgs’s coupling to fermions is proportional to the fermion’s mass $m_{f}$ as in the Yukawa interaction $(m_{f}/v_{h})(v_{h}+h_{1}+ih_{2})\bar{\varphi}\varphi$, which is confirmed by experiments. Hence, the next step is to go beyond the mean-field $U_{0}$ in Eq.(8). In the real Fock world, the massive gauge boson is transmitted from one massive fermions to the other. Fermions and antifermions exist not as a pair but as an independent particle, and their creation or annihilation is defined by a field [24]
$i\psi(x)=\int\frac{d^{3}p}{(2\pi)^{3}}\sum_{s}\frac{1}{\sqrt{2E_{p}}}\left[\widehat{a}^{s}(\mbox{$p$})\widehat{u}^{s}(p)e^{-ipx}+\widehat{b}^{s\dagger}(\mbox{$p$})\widehat{v}^{s}(p)e^{ipx}\right].$ (35)
When massive fermions in scattering do not have enough kinetic energy to generate the on-shell massive gauge boson, the non-relativistic limit is a good approximation in observer’s reference frame. The first approximation of the massive fermion-fermion scattering amplitude [25] is given by
$i{\cal M}=(-ig)^{2}\bar{\widehat{u}}^{s^{\prime}}(p_{1}^{\prime})\gamma^{\mu}\widehat{u}^{s}(p_{1})\frac{-ig^{\mu\nu}}{q^{2}-m_{B}^{2}}\bar{\widehat{u}}^{s^{\prime}}(p_{2}^{\prime})\gamma_{\nu}\widehat{u}^{s}(p_{2}).$ (36)
[…]
$\bar{\psi}(x)\left[i/\penalty 50{\!\!\!}{\partial}+\widehat{g}H(x)\right]\psi(x),$ (37)
where the scalar field $H(x)$ denotes this collective motion, and $\widehat{g}$ is given by
$\widehat{g}=\frac{m_{f}}{m_{B}}g.$ (38)
In $L^{\prime}_{h}(x)$ of Eq.(2), the coupling constant $m_{f}/v_{h}$ of the Higgs boson to fermions is, using the relation $m_{B}=gv_{h}$, equal to $(m_{f}/m_{B})g$, which agrees with this $\widehat{g}$. Hence, there arises a possibility that the scalar field $H(x)$ in Eq.(37) represents the Higgs field. Here we consider the possibility that the Higgs boson is not an elementary particle, but a quasiparticle representing a collective excitation.
The simplest collective excitation is a series of creations and annihilation of massive fermion and antifermion that propagate in space as in Figure 2(b), where black circles represent a vertex made of Eq.(37). The propagator of the Higgs field is given by
$\int\frac{d^{4}x}{(2\pi)^{4}}\langle\widetilde{0}|T[H(x)H(0)]|\widetilde{0}\rangle e^{iqx}=\frac{1}{q^{2}\left[1-\chi(q^{2})\right]}.$ (39)
Since this excitation is not a transverse one, there is no energy gap in its excitation spectrum. Here we define a finite upper end $\xi_{c}$ of energy-momentum of the excitation as in the non-relativistic physics. The self energy of the Higgs boson is given by
$iq^{2}\chi(q^{2})=(-i\widehat{g})^{2}(-1)\int^{\xi_{c}}_{0}\frac{d^{4}p}{(2\pi)^{4}}tr\left[\frac{i}{/\penalty 50{\!\!\!}{p}-m_{f}}\frac{i}{/\penalty 50{\!\!\!}{p}+/\penalty 50{\!\!\!}{q}-m_{f}}\right],$ (40)
in which $\gamma^{\mu}$ matrix is not there. This integral is not a loop of the relativistic process extending to infinite momentum, but an integral over $p$ from zero to this $\xi_{c}$. According to the ordinary rule, we use a new variable $l=p+xq$. The upper end in the integral over $l$ is $\sqrt{p^{2}+2xp\cdot q+x^{2}q^{2}}$, which depends on the relative direction of $p$ to $q$. Since the sign of $p\cdot q$ oscillates between positive and negative, we use a mean value $\sqrt{p^{2}+x^{2}q^{2}}$ for simplicity. Hence, using an Euclidian 4-momentum $l_{E}$ as $l^{2}=-l_{E}^{2}$, we obtain
$q^{2}\chi(q^{2})=-4\widehat{g}^{2}\int^{1}_{0}dx\int\frac{d\Omega_{4}}{(2\pi)^{4}}\int_{\sqrt{x^{2}q^{2}}}^{\sqrt{\xi_{c}^{2}+x^{2}q^{2}}}l_{E}^{3}dl_{E}\left[\frac{-l_{E}^{2}}{(l_{E}^{2}+\Delta)^{2}}+\frac{\Delta}{(l_{E}^{2}+\Delta)^{2}}\right],$ (41)
where $\Delta=m_{f}^{2}-x(1-x)q^{2}$. If we define the following integral
$I(m,n)\equiv\int l_{E}^{m}(l_{E}^{2}+\Delta)^{n}dl_{E},$ (42)
the indefinite integrals over $l_{E}$ in Eq.(41) are decomposed as follows
$I(5,-2)-\Delta\times I(3,-2)=I(1,0)+2\Delta^{2}\times I(1,-2)-3\Delta\times I(1,-1),$ (43)
where
$I(1,0)=\frac{1}{2}l_{E}^{2},\quad I(1,-2)=-\frac{1}{2(l_{E}^{2}+\Delta)},\quad I(1,-1)=\frac{1}{2}\ln|l_{E}^{2}+\Delta|.$ (44)
The definite integral over $l_{E}$ in Eq.(41) yields
$q^{2}\chi(q^{2})=\frac{\widehat{g}^{2}}{4\pi^{2}}\xi_{c}^{2}$ - $\frac{\widehat{g}^{2}}{2\pi^{2}}\int_{0}^{1}dx\Delta^{2}\left(\frac{1}{\xi_{c}^{2}+x^{2}q^{2}+\Delta}-\frac{1}{x^{2}q^{2}+\Delta}\right)$ (45)
- $\frac{\widehat{g}^{2}}{2\pi^{2}}\int_{0}^{1}dx\frac{3}{2}\Delta\ln\left|1+\frac{\xi_{c}^{2}}{x^{2}q^{2}+\Delta}\right|.$
With this $\chi(q^{2})$, the mass $m_{H}$ of the Higgs boson is defined as $\chi(q^{2})\simeq m_{H}^{2}/q^{2}$ at $q^{2}\rightarrow 0$. Since $\Delta\rightarrow m_{f}^{2}$ at $q^{2}\rightarrow 0$, the integrals over $x$ in the second and third terms of the right-hand side of Eq.(45) have following limits at $q^{2}\rightarrow 0$
$\int_{0}^{1}dx\frac{1}{\xi_{c}^{2}+x^{2}q^{2}+\Delta}\rightarrow\frac{3}{4(\xi_{c}^{2}+m_{f}^{2})},$ (46)
$\int_{0}^{1}dx\ln\left|1+\frac{\xi_{c}^{2}}{x^{2}q^{2}+\Delta}\right|\rightarrow\ln\left|\frac{\xi_{c}^{2}+m_{f}^{2}}{m_{f}^{2}}\right|.$ (47)
Plugging Eqs.(46) and (47) into Eq.(45), and using it in Eq.(39), the mass $m_{H}$ is given by
$m_{H}^{2}=\frac{\widehat{g}^{2}}{4\pi^{2}}\left[\xi_{c}^{2}+\frac{3}{2}m_{f}^{2}\left(1-\frac{m_{f}^{2}}{\xi_{c}^{2}+m_{f}^{2}}\right)-3m_{f}^{2}\ln\left(\frac{\xi_{c}^{2}+m_{f}^{2}}{m_{f}^{2}}\right)\right].$ (48)
The Higgs mass is related to $\xi_{c}$, $m_{f}$, and $m_{B}$ in $\widehat{g}$. This $\xi_{c}$ is to be determined by experimental value of $m_{H}$. Once a finite value of $\xi_{c}$ is obtained, divergence that we face in further renormalization is not a quadratic, but a logarithmic divergence, because we need not use the Higgs potential.
The Higgs field is described by the following effective Lagrangian density
$(\partial_{\mu}H)^{2}-m_{H}^{2}H^{2}+\frac{m_{f}}{m_{B}}g\bar{\psi}\psi H.$ (49)
Whereas the masses of fermion and gauge boson have their origin in the non-Fock vacuum, the Higgs boson is a many-body phenomenon in the Fock vacuum.
Compared to the BEH model $L_{0}(x)+L^{\prime}_{h}(x)$, this picture of Higgs boson has following features.
(1) In the experimental tests of Higgs particle, the Higgs coupling to massive gauge bosons is an important channel. In the BEH model, the gauge coupling $|(i\partial_{\mu}+gB_{\mu})(v_{h}+h_{1}+ih_{2})|^{2}$ predicts a direct coupling of the massive gauge boson to the Higgs field $h_{1}$ or Goldstone mode $h_{2}$ as follows
$g^{2}v_{h}^{2}B^{\mu}B_{\mu}\left(1+\frac{h_{1}}{v_{h}}\right)^{2}+g^{2}B^{\mu}B_{\mu}h_{2}^{2}+2gB^{\mu}(h_{1}\partial_{\mu}h_{2}+h_{2}\partial_{\mu}h_{1})+(c.c).$ (50)
In the present model, the Higgs field $H$ is not an elementary field but represents collective excitation, so that gauge coupling of $H$ with the same $g$ does not exist. Instead of $m_{B}^{2}B^{\mu}B_{\mu}(1+h_{1}/v_{h})^{2}$, the effective coupling of $H$ to $B_{\mu}$ arises in the perturbation process induced by $\widehat{g}\bar{\psi}\psi H$ and $g\bar{\psi}\gamma^{\mu}\psi B_{\mu}$, in which such couplings are relayed by $\bar{\psi}$ and $\psi$. The gauge coupling of $h_{1}$ and $h_{2}$ in the BEH mechanism may be a useful phenomenology if $g$ is adjusted properly.
(2) In the GWS model, Higgs and Goldstone couplings to fermions are proportional to fermion’s mass. In the present model, such couplings are made by $(m_{f}/m_{B})g\bar{\psi}\psi H$ in Eq.(49) and $(\sqrt{2}g/m_{B})\bar{\psi}\gamma^{\mu}\psi\partial_{\mu}G$ in Eq.(34), respectively. For the Goldstone’s coupling to the fermion, it arises from the vacuum $|\widetilde{0}\rangle$ itself. The Yukawa interaction $(m_{f}/v_{h})(v_{h}+h_{1}+ih_{2})\bar{\varphi}\varphi$ is a simple phenomenology for the response of physical vacuum to the gauge field.
6 Discussion
The following total Lagrangian density $\widetilde{L}(x)$ on the Fock space is obtained. After rewriting $\varphi$ and $\bar{\varphi}$ with $\psi$ and $\bar{\psi}$, $\widetilde{L}(x)$ is given by
$\widetilde{L}(x)=$ - $\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+\bar{\psi}(i\partial_{\mu}+gB_{\mu})\gamma^{\mu}\psi-m_{f}\bar{\psi}\psi$ (51)
$+$ $\frac{1}{2}m_{B}^{2}B^{\mu}B_{\mu}+\frac{m_{B}}{\sqrt{2}}B_{\mu}\partial^{\mu}G+(\partial_{\mu}G)^{2}+\frac{\sqrt{2}g}{m_{B}}\bar{\psi}\gamma^{\mu}\psi\partial_{\mu}G$
$+$ $(\partial_{\mu}H)^{2}-m_{H}^{2}H^{2}+\frac{m_{f}}{m_{B}}g\bar{\psi}\psi H.$
In this $\widetilde{L}(x)$, the double role of the Higgs potential is dissolved, and each role is played by each physical process. Fermion’s mass and the vacuum condensate are not related. Chiral-symmetry breaking is not induced by the Higgs potential, but occurs in the restoration process of the Fock vacuum. The vacuum condensate $v_{h}$ arises from the statistical gap in the transverse excitation. Whereas the Higgs and Goldstone fields in the BEH model are defined by the Higgs potential, the Higgs boson represents the many-body phenomenon in the real world, and the Goldstone field arises from the phase factor in the vacuum $|\widetilde{0}\rangle$. Parameters $\mu$ and $\lambda$ in the Higgs potential are replaced by parameters that have physical interpretations,
$m_{B}^{2}$ $=$ $g^{2}\left(\frac{2l_{c}^{2}}{d_{m}^{3}}\right)^{2},\qquad(=g^{2}\frac{\mu^{2}}{2\lambda})$
$m_{f}$ $=$ $|U_{0}|.$
$m_{H}^{2}$ $=$ $\frac{\widehat{g}^{2}}{4\pi^{2}}\left[\xi_{c}^{2}+\frac{3}{2}U_{0}^{2}\left(1-\frac{U_{0}^{2}}{\xi_{c}^{2}+U_{0}^{2}}\right)-3U_{0}^{2}\ln\left(1+\frac{\xi_{c}^{2}}{U_{0}^{2}}\right)\right],\qquad(=2\mu^{2})$ (52)
where $\widehat{g}=(m_{f}/m_{B})g=|U_{0}|(d_{m}^{3}/\sqrt{2}l_{c}^{2})$.
The present model has implications for some fundamental problems that lie in the BEH model.
(a) The BEH mechanism, before applied to the electroweak interaction as a part of the GWS model, was nothing but a theoretical possibility. When applied to the electroweak interaction, this mechanism became a realistic model, and the mass acquisition not only for gauge boson but also for fermion became necessary. The Yukawa interaction was a useful phenomenological resolution to it, but the origin of the vacuum condensate $v_{h}$ remains to be proved. A concrete image of $v_{h}$ can be obtained when the distribution of massless fermion and antifermion in the non-Fock vacuum is considered in position space.
(b) Since the present model does not assume the Higgs potential $4\lambda v_{h}h_{1}^{3}-\lambda h_{1}^{4}$, the quadratic divergence does not occur in the perturbation calculation, provided that $\xi_{c}$ is determined using the experimental value of $m_{H}$ in Eq.(52). The divergence we must renormalize is only logarithmic one. The present model proposes a solution to the fine-tuning problem of the quadratic divergence.
(c) According to the lattice model, which strictly preserves local gauge invariance at each stage of argument, the vacuum expectation value (VEV) of the gauge-dependent quantity vanishes when calculated without gauge fixing. Hence, if $h(x)$ follows $h(x)\rightarrow h(x)\exp(i\theta(x))$ under $A_{\mu}(x)\rightarrow A_{\mu}(x)-ie^{-1}\partial_{\mu}\theta(x)$, $\langle h(x)\rangle=0$ is unavoidable. (Elitzur-De Angelis-De Falco-Guerra theorem) [26][27]. This is because the local character of gauge symmetry effectively breaks the connection in the degrees of freedom defined at different space-time points. If the Higgs boson is an elementary particle, the finite $\langle h(x)\rangle$ that depends on the gauge-fixing procedure does not match its fundamental nature. Instead of $v_{h}=\langle h(x)\rangle$, the vacuum is characterized by $\langle\widetilde{0}|\int[\varphi^{\dagger}(0)\varphi(0)]^{2}d^{4}Y|\widetilde{0}\rangle$ in Eq.(25). Because this condensate is gauge invariant, there is no need to worry about the vanishing of its VEV.
(d) In the Higgs potential, $\lambda|v_{h}+h_{1}+ih_{2}|^{4}$ predicts the triple and quartic self-couplings of the Higgs boson. More complex many-body effects than that in Figure 2(b) may correspond to such self couplings. So far, the agreement of the GWS model to experiments is satisfactory. When more precise measurements are performed, however, there is a possibility of deviation, especially for light quarks and leptons of the first and second generations. The above $\widetilde{L}(x)$ predicts some different results from those by the BEH model. The next subject is to extend it to the electroweak interaction.
Appendix A The vacuum $|\widetilde{0}\rangle$ satisfying $\widetilde{a}^{s}(\mbox{$p$})|\widetilde{0}\rangle=\widetilde{b}^{s}(-\mbox{$p$})|\widetilde{0}\rangle=0$
The vacuum satisfying $\widetilde{a}^{s}(\mbox{$p$})|\widetilde{0}\rangle=\widetilde{b}^{s}(-\mbox{$p$})|\widetilde{0}\rangle=0$ is as follows. In the expansion
$e^{-iK}Fe^{iK}=F+[-iK,F]+\frac{1}{2!}\left[-iK,[-iK,F]\right]+\cdots,$ (53)
we regard an operator $a^{s}(\mbox{$p$})$ or $b^{s}(-\mbox{$p$})$ as $F$, and the following operator as $K$,
$i\sum_{p,s}\theta_{\mbox{$p$}}[b^{s\dagger}(-\mbox{$p$})a^{s\dagger}(\mbox{$p$})-a^{s}(\mbox{$p$})b^{s}(-\mbox{$p$})].$ (54)
Hence, Eqs.(5) and (6) can be rewritten in a compact form
$\widetilde{a}^{s}(\mbox{$p$})=e^{-iK}a^{s}(\mbox{$p$})e^{iK},\quad\widetilde{b}^{s}(-\mbox{$p$})=e^{-iK}b^{s}(-\mbox{$p$})e^{iK}.$ (55)
The vacuum $|\widetilde{0}\rangle$ satisfying $\widetilde{a}^{s}(\mbox{$p$})|\widetilde{0}\rangle=\widetilde{b}^{s}(-\mbox{$p$})|\widetilde{0}\rangle=0$ is simply expressed as $|\widetilde{0}\rangle=e^{-iK}|0\rangle$. Hence, we obtain
$|\widetilde{0}\rangle$ $=$ $\exp\left(\sum_{p,s}\theta_{\mbox{$p$}}[b^{s\dagger}(-\mbox{$p$})a^{s\dagger}(\mbox{$p$})-a^{s}(\mbox{$p$})b^{s}(-\mbox{$p$})]\right)|0\rangle.$ (56)
$=$ $\prod_{p,s}\left[\sum_{n}\frac{1}{n!}\theta_{\mbox{$p$}}^{n}[b^{s\dagger}(-\mbox{$p$})a^{s\dagger}(\mbox{$p$})-a^{s}(\mbox{$p$})b^{s}(-\mbox{$p$})]^{n}\right]|0\rangle.$
[…]
$\sum_{n}\frac{\theta^{n}}{n!}(b^{\dagger}a^{\dagger}-ab)^{n}|0\rangle$ $=$ $|0\rangle+\theta b^{\dagger}a^{\dagger}|0\rangle-\frac{\theta^{2}}{2!}abb^{\dagger}a^{\dagger}|0\rangle-\frac{\theta^{3}}{3!}b^{\dagger}a^{\dagger}abb^{\dagger}a^{\dagger}|0\rangle$ (57)
$+$ $\frac{\theta^{4}}{4!}abb^{\dagger}a^{\dagger}abb^{\dagger}a^{\dagger}|0\rangle+\cdots.$
In this expansion, $\cos\theta_{\mbox{$p$}}$ appears in the sum of even-order terms of $\theta$, and $\sin\theta_{\mbox{$p$}}$ appears in the sum of odd-order terms, and then Eq.(7) is yielded.
Appendix B Metric and energy gap
(1) Consider a Bose field $f(x)$ with the following kinetic energy
$H_{ef}=-\int d^{3}xf^{\dagger}(x)\Delta f(x).$ (58)
The square of the infinitesimal line-element is a quadratic function of $dx$ as $dl^{2}=g_{\mu\nu}(x)dx^{\mu}dx^{\nu}$, where $g_{\mu\nu}(x)$ is a metric tensor ($\mu,\nu=x,y,z$). The inner product of the field $f(x)$ is defined as
$\langle f(x)|f(x)\rangle=\int\sqrt{|g(x)|}d^{3}xf^{\dagger}(x)f(x),$ (59)
where $|g(x)|$ is a determinant of $g_{\mu\nu}(x)$. Consider a gradient of $f$ like $A^{\mu}=g_{\mu\nu}\partial f/\partial x^{\nu}$. Since the metric depends on the position, a derivative of a given vector $A^{\mu}$ with respect to $x^{\mu}$ is a covariant derivative as follows
$\frac{DA^{\mu}}{dx^{\mu}}=\frac{dA^{\mu}}{dx^{\mu}}+\Gamma^{\mu}_{\nu\mu}A^{\nu},$ (60)
where $\Gamma^{\mu}_{\nu\mu}$ is a connection coefficient. This $\Gamma^{\mu}_{\nu\mu}$ is expressed by the determinant of the metric tensor $|g(x)|=|g_{\mu\nu}(x)|$ as follows
$\Gamma^{\mu}_{\nu\mu}=\frac{1}{2|g|}\frac{\partial|g|}{\partial x^{\nu}}.$ (61)
With this expression, the covariant derivative of $A^{\mu}$ is given by
$\frac{DA^{\mu}}{dx^{\mu}}=\frac{1}{\sqrt{|g|}}\frac{\partial(\sqrt{|g|}A^{\mu})}{\partial x^{\mu}}.$ (62)
(2) The matrix element of the Laplacian operator is obtained by inserting $D\partial/\partial^{2}x$ between $f^{\dagger}(x)$ and $f(x)$ in Eq.(59). For $D\partial f/\partial^{2}x$, we use Eq.(62) with $A^{\mu}=g_{\mu\nu}\partial f/\partial x^{\nu}$. After integration by parts, we get
$\langle f(x)|-\Delta|f(x)\rangle$ $=$ $-\int\sqrt{|g|}d^{3}xf^{\dagger}\frac{1}{\sqrt{|g|}}\frac{\partial}{\partial x^{\mu}}\left(\sqrt{|g|}g_{\mu\nu}\frac{\partial f}{\partial x^{\nu}}\right)$ (63)
$=$ $\int\sqrt{|g(x)|}g_{\mu\nu}d^{3}x\frac{\partial f^{\dagger}}{\partial x^{\mu}}\frac{\partial f}{\partial x^{\nu}}.$
Our interest is how the metric reflecting permutation symmetry of $f$ looks in the flat space, because we observe it as an effect of Bose statistics in experiments. To rewrite Eq.(63), we introduce a new field $\widehat{f}(x)=|g(x)|^{1/4}f(x)$ as follows
$\int\sqrt{|g(x)|}d^{3}xf^{\dagger}(x)f(x)=\int d^{3}x\widehat{f}^{\dagger}(x)\widehat{f}(x).$ (64)
The gradient in Eq.(58) is rewritten using $\widehat{f}(x)$
$\frac{\partial f}{\partial x}=|g|^{-1/4}\left(\frac{\partial}{\partial x}-\frac{1}{4}\frac{\partial\ln|g|}{\partial x}\right)\widehat{f}(x).$ (65)
Hence, the matrix element of Laplacian in Eq.(63)
$\langle f(x)|-\Delta|f(x)\rangle=\int d^{3}xg_{\mu\nu}\left(\frac{\partial}{\partial x^{\mu}}-\frac{1}{4}\frac{\partial\ln|g|}{\partial x^{\mu}}\right)\widehat{f}^{\dagger}(x)\left(\frac{\partial}{\partial x^{\nu}}-\frac{1}{4}\frac{\partial\ln|g|}{\partial x^{\nu}}\right)\widehat{f}(x),$ (66)
is rewritten using the integration by parts as follows
$\langle f(x)|-\Delta|f(x)\rangle=\int d^{3}xg_{\mu\nu}\frac{\partial\widehat{f}^{\dagger}}{\partial x^{\mu}}\frac{\partial\widehat{f}}{\partial x^{\nu}}+\int W(x)\widehat{f}^{\dagger}(x)\widehat{f}(x)d^{3}x,$ (67)
where
$W(x)=\frac{1}{4}\frac{\partial}{\partial x^{\mu}}\left(g_{\mu\nu}\frac{\partial\ln|g|}{\partial x^{\nu}}\right)+\frac{1}{16}g_{\mu\nu}\left(\frac{\partial\ln|g|}{\partial x^{\mu}}\right)\left(\frac{\partial\ln|g|}{\partial x^{\nu}}\right).$ (68)
This $W(x)$ is a square of the finite energy gap in the excitation spectrum of $\widehat{f}(x)$ [22].